### v1

# GARCH copulas and GARCH-mimicking copulas

Abstract
The bivariate copulas that describe the dependencies and partial dependencies of lagged variables in strictly stationary, first-order GARCH-type processes are investigated. It is shown that the copulas of symmetric GARCH processes are jointly symmetric but non-exchangeable, while the copulas of processes with symmetric innovation distributions and asymmetric leverage effects have weaker h-symmetry; copulas with asymmetric innovation distributions have neither form of symmetry. Since the actual copulas are typically inaccessible, due to the unknown functional forms of the marginal distributions of GARCH processes, methods of mimicking them are proposed. These rely on constructions that combine standard bivariate copulas for positive dependence with two uniformity-preserving transformations known as v-transforms. A variety of new copulas are introduced and the ones providing the best fit to simulated data from GARCH-type processes are identified. A method of constructing tractable simplified d-vines using linear v-transforms is described and shown to coincide with the vt-d-vine model when the two v-transforms are identical.
Keywords: Time series; copulas; vine copulas; GARCH processes; v-transforms.
1 Introduction
The ARCH process (Engle,, 1982), the GARCH process (Bollerslev,, 1986) and a whole lexicon of related GARCH-type models have proved to be remarkably successful models for time series displaying stochastic volatility, such as asset price return series. In this paper we analyse the serial dependence structure of first-order GARCH-type models in terms of the implied bivariate copulas that describe the dependence of pairs of variables at different lags. Our aim is to understand how we could mimic such dependence structures with appropriately chosen bivariate copulas arranged in d-vines; we do this with a view to developing models that might offer additional flexibility in practice. The approach we take recognizes that uniformity-preserving transformations, known as v-transforms (McNeil,, 2021), are embedded in the implied copulas of classical GARCH models and suggests that they could be used as explicit modelling elements to construct vines for applications.
The essential mechanism of a GARCH-type model is simple and intuitive; the conditional distribution from which a new observation is drawn has a standard deviation that is a function of the size of previous returns, so that large market moves have a tendency to beget further large market moves and create periods of volatility. The ARCH and GARCH prototypes of Engle, (1982) and Bollerslev, (1986) have inspired a vast class of related GARCH-type models, including the GJR-GARCH model, which incorporates leverage effects to reflect different contributions to volatility by returns of differing signs (Glosten et al.,, 1993), power-GARCH models (Ding et al.,, 1993) and exponential GARCH models (Nelson,, 1991), to name a but a few of the more important extensions. A number of comprehensive surveys of the extended GARCH family have been published including Bollerslev et al., (1992) Bollerslev et al., (1994) and Shephard, (1996).
Notwithstanding the appealing basic mechanism for explaining volatility, GARCH-type models are also something of a black box. The stationary distribution is not generally known in closed form, although it is known that many models gives rise to distributions with Pareto tails (Mikosch and Stărică,, 2000). In particular, the serial dependence structure is not well understood. If $(X_{t})_{t\in\mathbb{Z}}$ follows a stationary GARCH-type model then the distribution of $(X_{1},\ldots,X_{n})^{\top}$ for $n\geqslant 2$ must have a copula, but this copula also has no simple, accessible closed form.
It may be asked why we would want to mimic GARCH-type models with vine copulas, given that GARCH models have proved themselves in practice over four decades. While many asset return series are well modelled by GARCH, we do encounter series where a power-lawed marginal distribution is sub-optimal for the data; see, for example, Bladt and McNeil, 2022b . For such datasets it would be convenient to be able to construct models that combined bespoke choices of marginal distribution, or non-parametric margins, with a copula that could mimic the kind of serial dependence that is implicit in the GARCH family. If we understand the types of copula that are implicit in the GARCH family, it may also be possible to improve upon these copulas by adding further flexibility.
Vine copulas also have a large literature, including Joe, (1996, 1997), Bedford and Cooke, (2001, 2002), Kurowicka and Cooke, (2006), Aas et al., (2009) and Smith et al., (2010). The most suitable vine structure for a time series is the d-vine, which is able to describe strict stationarity of a random vector under additional translation-invariance restrictions on the vine structure (Nagler et al.,, 2022). Markov copula models (Darsow et al.,, 1992; Chen and Fan,, 2006; Beare,, 2010; Domma et al.,, 2009) are simple examples of stationary first-order d-vines and a number of authors have written on their higher-order extensions (Ibragimov,, 2009; Beare and Seo,, 2015; Brechmann and Czado,, 2015; Loaiza-Maya et al.,, 2018; Nagler et al.,, 2022). The last-named paper introduces the terminology s-vine for stationary multivariate vines, which includes the stationary univariate d-vine as a special case.
The joint density of the distribution of $(X_{1},\ldots,X_{n})^{\top}$ in a strictly stationary process, such as a stationary model of GARCH type, can be decomposed as a d-vine. In other words we may write
$f_{X_{1},\ldots,X_{n}}(x_{1},\ldots,x_{n})=\prod_{i=1}^{n}f_{X}(x_{i})\prod_{k=1}^{n-1}\prod_{j=k+1}^{n}c_{k}(F_{X_{1}|X_{2},\ldots,X_{k}}(x_{j-k}\mid x_{j-k+1},\ldots,x_{j-1}),\\ F_{X_{k+1}|X_{2},\ldots,X_{k}}(x_{j}\mid x_{j-k+1},\ldots,x_{j-1})\mid x_{j-k+1},\ldots,x_{j-1})$ (1)
where $f_{X}$ denotes the marginal density, $F_{X_{i}|X_{2},\ldots,X_{k}}$ denotes the conditional distribution function of $X_{i}$ given $X_{2},\ldots,X_{k}$ and $c_{k}(u,v\mid x_{2},\ldots,x_{k})$ is the density of the copula of the conditional distribution of $(X_{1},X_{k+1})^{\top}$ given $X_{2}=x_{2},\ldots,X_{k}=x_{k}$. In the case $k=1$ the conditional distribution is simply the marginal distribution $F_{X}$ and $c_{1}$ is simply the density of the copula of the unconditional distribution of $(X_{1},X_{2})^{\top}$. When the process is first-order Markov, all of the terms involving $c_{k}(u,v\mid\cdot)$ for $k>1$ disappear, but this will only apply to the ARCH(1)-type models among the models we analyse.
Although the densities of GARCH-type models may be written as d-vines, they may not in general be written as so-called simplified d-vines, that is models in which the conditional copula densities $c_{k}$ do not depend explicitly on the values $x_{j-k+1},\ldots,x_{j-1}$ in equation (1) for $k\geqslant 2$. When d-vines are used in practical modelling applications they are almost always simplified d-vines. However, as pointed out by a number of authors (Haff et al.,, 2010; Stöber et al.,, 2013; Spanhel and Kurz,, 2019; Mroz et al.,, 2021), these may not be able to capture all the features present in an arbitrary joint density, such as that of a GARCH-type model. In the course of our analysis we return to this question.
We confine our attention to first-order GARCH-type models. In practice models of order (1,1) often work very well and adding higher-order terms yields only minor improvements. We also look in some detail at ARCH(1)-type models for two reasons: they are first-order Markov models and therefore particularly easy to analyse; they give rise to the essential copula forms that are encountered in the more general class, as will be seen in this paper.
Loaiza-Maya et al., (2018) have previously investigated the copula of the classical ARCH(1) process. Since this model is first-order Markov, its serial dependence may be fully described by a first-order d-vine with a single bivariate copula, namely the copula of the distribution of two successive values $(X_{1},X_{2})^{\top}$. Using a simulation approach, they identify the cross-shaped nature of the copula density and suggest that mixtures of positive dependence and negative dependence copulas (such as a t copula and a 90-degree rotated t copula) can model the dependence structure effectively. Zhao et al., (2022) argue that t copulas in d-vines can approximate the dependence structure of GARCH(1,1) and GJR-GARCH(1,1) models. We will find some evidence in support of t copulas but also argue that they miss some important asymmetries.
We make a number of original contributions. First, we analyse the structure of the bivariate copula densities $c_{k}$ in (1) for classical GARCH-type models; we show that symmetric models give rise to jointly symmetric copulas while models with symmetric innovation distributions and asymmetric volatility have copulas with weaker symmetry, which we call h-symmetry. We confirm the theoretical insights by using numerical approximations to provide accurate contour plots of the copula densities. These reveal hitherto overlooked features, such as the non-exchangeability of the copulas. We also examine the effects of leverage, asymmetric innovations and autoregressive conditional mean terms on the copulas and on their symmetries.
Second, we show how uniformity-preserving transformations known as v-transforms (McNeil,, 2021) play an implicit role in the copulas of classic symmetric models and in models with certain forms of asymmetry. We investigate methods of constructing copulas with v-transforms to mimic the behaviour of the copulas implied by GARCH-type models. Using simulated data we identify the copulas based on v-transforms that provide the best fit to simulated data and show that these typically do better than mixture copulas.
Finally, we draw conclusions with respect to the construction of d-vine copulas to mimic the serial-dependence structure of GARCH-type models. We also show how d-vine models with appropriately chosen copulas can be related to the vt-ARMA and vt-d-vine models proposed by McNeil, (2021) and Bladt and McNeil, 2022b .
2 Preliminary results on symmetry
2.1 Definitions
Symmetry plays an important role in this paper and we give a definition to make our use of various symmetry concepts clear; parts (iii) and (iv) are from Nelsen, (2006).
Definition 1.
Let $(Y,Z)^{\top}$ be a generic random vector taking values in $\mathbb{R}^{2}$.
  1. (i)
    If $(Y,Z-b)^{\top}\stackrel{{\scriptstyle\text{\tiny d}}}{{=}}(Y,b-Z)^{\top}$ for some $b\in\mathbb{R}$ we say that $(Y,Z)^{\top}$ is h-symmetric about $b$ (symmetric about a horizontal axis at $b$).
  2. (ii)
    If $(Y-a,Z)^{\top}\stackrel{{\scriptstyle\text{\tiny d}}}{{=}}(a-Y,Z)^{\top}$ for some $a\in\mathbb{R}$ we say that $(Y,Z)^{\top}$ is v-symmetric about $a$ (symmetric about a vertical axis at $a$).
  3. (iii)
    If $(Y-a,Z-b)^{\top}\stackrel{{\scriptstyle\text{\tiny d}}}{{=}}(a-Y,b-Z)^{\top}$ for $a,b\in\mathbb{R}$ we say that $(Y,Z)^{\top}$ is radially symmetry about $(a,b)$.
  4. (iv)
    If $(Y,Z)^{\top}$ has all three of the above forms of symmetry then we say that it is jointly symmetric about $(a,b)$.
Remarks 1.
  1. 1.
    When $(Y,Z)^{\top}$ is h-symmetric about $b$, the marginal distribution of $Z$ is symmetric about $b$; when $(Y,Z)^{\top}$ is v-symmetric about $a$, the marginal distribution of $Y$ is symmetric about $a$. The reverse implications are not true.
  2. 2.
    It may be noted that we can always shift the distribution of $(Y,Z)^{\top}$ in two-dimensional space so that axes or centres of symmetry are at zero or the origin. Where $a$ and $b$ are not relevant to the result, we drop the explicit reference to them and just refer to h-symmetry, v-symmetry, radial symmetry and joint symmetry.
  3. 3.
    H-symmetry and v-symmetry together imply radial symmetry and therefore joint symmetry. However radial symmetry does not necessarily imply h-symmetry and v-symmetry, although it does imply symmetry of both marginal distributions.
The symmetries in Definition 1 have the following implications for the copula of $(Y,Z)^{\top}$.
Lemma 1.
Let $(Y,Z)^{\top}$ be a random vector with continuous marginal distributions and copula $C$.
  1. (i)
    If $(Y,Z)^{\top}$ is h-symmetric then $C(u,v)=u-C(u,1-v)$, for $(u,v)\in[0,1]^{2}$.
  2. (ii)
    If $(Y,Z)^{\top}$ is v-symmetric then $C(u,v)=v-C(1-u,v)$, for $(u,v)\in[0,1]^{2}$,
  3. (iii)
    If $(Y,Z)^{\top}$ is radially symmetric then $C(u,v)=u+v+C(1-u,1-v)-1$ , for $(u,v)\in[0,1]^{2}$.
Proof.
Since the copula is invariant under strictly increasing transformations of the margins, such as changes of location, we can, without loss of generality, assume that the axes and centres of symmetry are at zero and the origin. For case (i) observe that if $(Y,Z)^{\top}\stackrel{{\scriptstyle\text{\tiny d}}}{{=}}(Y,-Z)^{\top}$ then the joint distribution function satisfies
$F_{Y,Z}(y,z)=\mathbb{P}(Y\leqslant y,Z\leqslant z)=\mathbb{P}(Y\leqslant y,-Z\leqslant z)$ $=\mathbb{P}(Y\leqslant y)-P(Y\leqslant y,Z\leqslant-z)$
$=F_{Y}(y)-F_{Y,Z}(y,-z)$ (2)
for all $(y,z)\in\mathbb{R}^{2}$ and application of Sklar’s Theorem (Sklar,, 1959) implies that the (unique) copula $C$ must satisfy
$C(F_{Y}(y),F_{Z}(z))=F_{Y}(y)-C(F_{Y}(y),F_{Z}(-z))=F_{Y}(y)-C(F_{Y}(y),1-F_{Z}(z))$
where $F_{Y}$ and $F_{Z}$ denote the marginal distribution functions and the final equality follows from the symmetry of the marginal distribution of $Z$. Hence we obtain that $C(u,v)=u-C(u,1-v)$ at all points in the unit square. The proof of (ii) is obviously identical to (i) while (iii) is a well-known result found in Nelsen, (2006) (Theorem 2.7.3). ∎
Definition 2.
A copula $C$ satisfying $C(u,v)=u-C(u,1-v)$ for all $(u,v)\in[0,1]^{2}$ will be said to have h-symmetry. Similarly, a copula satisfying $C(u,v)=v-C(1-u,v)$ will be said to have v-symmetry. A copula satisfying $C(u,v)=u+v+C(1-u,1-v)-1$ will be said to have radial symmetry. A copula with all three forms of symmetry will be said to be jointly symmetric.
We note that joint symmetry of the copula is the bivariate special case of a property that Durante and Fuchs, (2019) call reflection invariance.
2.2 Correlation and rank correlation
Nelsen, (2006) notes that jointly symmetric distributions have zero correlation (when finite second moments permit the measure to be defined). Perhaps surprisingly, h-symmetry on its own is a sufficient condition for an absence of both correlation and rank correlation. The same is true of v-symmetry.
Proposition 1.
If a random vector $(Y,Z)^{\top}$ has finite second moments and is either h-symmetric or v-symmetric then the Pearson correlation $\rho(Y,Z)$ is zero.
Proof.
We give the proof in the case of h-symmetry, the method for v-symmetry being identical. If $(Y,Z)^{\top}$ is h-symmetric about $b$ then $\mathbb{E}(Z)=b$ and we can write $\operatorname{cov}(Y,Z)=\mathbb{E}(YZ)-\mathbb{E}(Y)\mathbb{E}(Z)=\mathbb{E}(Y(Z-b))+\mathbb{E}(Yb)-\mathbb{E}(Y)b=\mathbb{E}(Y(Z-b))$. However, h-symmetry implies $\mathbb{E}(Y(Z-b))=\mathbb{E}(Y(b-Z))=-\mathbb{E}(Y,(Z-b))$ implying that $\mathbb{E}(Y(Z-b))=0$ from which the result follows. ∎
Proposition 2.
If a random vector $(Y,Z)^{\top}$ has continuous marginal distributions and is either h-symmetric or v-symmetric, then the Spearman rank correlation $\rho_{S}(Y,Z)$ and the Kendall rank correlation $\tau(Y,Z)$ are both zero.
Proof.
Again it suffices to give the proof in the case of h-symmetry. Let the unique copula of $(Y,Z)^{\top}$ be denoted $C$. For Spearman’s rank correlation we need to calculate $\rho_{S}(Y,Z)=12\int_{0}^{1}\int_{0}^{1}(C(u,v)-uv)\mathrm{d}u\mathrm{d}v$. From the identity $C(u,v)=u-C(u,1-v)$ we infer that
$\int_{0}^{1}\int_{0}^{1}C(u,v)\mathrm{d}u\mathrm{d}v$ $=$ $\int_{0}^{1}\int_{0}^{1}u\mathrm{d}u\mathrm{d}v-\int_{0}^{1}\int_{0}^{1}C(u,1-v)\mathrm{d}u\mathrm{d}v=\frac{1}{2}-\int_{0}^{1}\int_{0}^{1}C(u,y)\mathrm{d}u\mathrm{d}y$
where we have made the change of variable $y=1-v$. We must have $\int_{0}^{1}\int_{0}^{1}C(u,v)\mathrm{d}u\mathrm{d}v=\tfrac{1}{4}$ and then simple calculations give $\rho_{S}(Y,Z)=0$.
In the case of Kendall’s rank correlation we need to compute $\tau(Y,Z)=4\mathbb{E}(C(U,V))-1$ where $(U,V)^{\top}$ is a pair of variables with distribution function $C$. The identity $C(u,v)=u-C(u,1-v)$ implies that $(U,V)^{\top}\stackrel{{\scriptstyle\text{\tiny d}}}{{=}}(U,1-V)^{\top}$; indeed $(U,V-0.5)^{\top}\stackrel{{\scriptstyle\text{\tiny d}}}{{=}}(U,0.5-V)^{\top}$ since $(U,V)^{\top}$ is h-symmetric about 0.5. Hence we infer that
$\mathbb{E}(C(U,V))$ $=$ $\mathbb{E}(U)-\mathbb{E}(C(U,1-V))=\frac{1}{2}-\mathbb{E}(C(U,V))$
which implies that $\mathbb{E}(C(U,V))=\tfrac{1}{4}$ and $\tau(Y,Z)=0$. ∎
Despite the absence of correlation in distributions with these forms of symmetry, we will see in this paper that they can have quite strong forms of dependence.
2.3 Copula densities
We will analyse GARCH-type models in terms of their joint densities. Henceforth we assume that the generic random vector $(Y,Z)^{\top}$ has joint density $f_{Y,Z}$. It is easily inferred from Definition 1, part (i), that h-symmetry is equivalent to $f_{Y,Z}(y,b+z)=f_{Y,Z}(y,b-z)$ for all $(y,z)\in\mathbb{R}^{2}$ and some $b\in\mathbb{R}$. Similarly, v-symmetry and radial symmetry are equivalent to the density satisfying $f_{Y,Z}(a+y,z)=f_{Y,Z}(a-y,z)$ for some $a\in\mathbb{R}$ and $f_{Y,Z}(a+y,b+z)=f_{Y,Z}(a-y,b-z)$ for some $a,b\in\mathbb{R}$, respectively.
The density $c$ of the copula $C$ of $(Y,Z)^{\top}$ is given by
$c(u,v)=\frac{f_{Y,Z}\left(F_{Y}^{-1}(u),F_{Z}^{-1}(v)\right)}{f_{Y}\left(F_{Y}^{-1}(u)\right)f_{Z}\left(F_{Z}^{-1}(v)\right)}$ (3)
where $f_{Y}$ and $f_{Z}$ denote the marginal densities. The identity $C(u,v)=u-C(u,1-v)$ for h-symmetry of the copula $C$ in Definition 2 is equivalent to the identity $c(u,v)=c(u,1-v)$ for densities (symmetry in the line $v=0.5$). Similarly, v-symmetry is equivalent to $c(u,v)=c(1-u,v)$ (symmetry in the line $u=0.5$) and radial symmetry is equivalent to $c(u,v)=c(1-u,1-v)$.
The following result provides the first connection between copula densities and v-transforms. The function $\mathcal{V}(u)=|2u-1|$ appearing in this result is the first and simplest example of a v-transform, as we will discuss in Section 4.
Proposition 3.
Let $(Y,Z)^{\top}$ be a random vector with joint density $f_{Y,Z}$ and copula density $c$. Then, for all $(u,v)\in[0,1]^{2}$,
  1. (i)
    if $(Y,Z)^{\top}$ is h-symmetric about $b$, $c(u,v)=c^{*}(u,|2v-1|)$ where $c^{*}$ is the density of the copula of $(Y,|Z-b|)^{\top}$;
  2. (ii)
    if $(Y,Z)^{\top}$ is v-symmetric about $a$, $c(u,v)=c^{*}(|2u-1|,v)$ where $c^{*}$ is the density of the copula of $(|Y-a|,Z)^{\top}$.
Proof.
It suffices to show part (i). We set $\tilde{Z}=Z-b$ and observe that $(Y,\tilde{Z})^{\top}$ is h-symmetric about 0. We then find that
$F_{Y,|\tilde{Z}|}(y,z)=\mathbb{P}(Y\leqslant y,|\tilde{Z}|\leqslant z)=\mathbb{P}(Y\leqslant y,\tilde{Z}\leqslant z)-\mathbb{P}(Y\leqslant y,\tilde{Z}\leqslant-z)=F_{Y,\tilde{Z}}(y,z)-F_{Y,\tilde{Z}}(y,-z)$
and hence $f_{Y,|\tilde{Z}|}(y,z)=f_{Y,\tilde{Z}}(y,z)+f_{Y,\tilde{Z}}(y,-z)=2f_{Y,\tilde{Z}}(y,z)$ by h-symmetry about zero. Similarly, $f_{|\tilde{Z}|}(z)=2f_{\tilde{Z}}(z)$. From (3) we have
$c^{*}(u,|2v-1|)=\frac{f_{Y,|\tilde{Z}|}\left(F_{Y}^{-1}(u),F_{|\tilde{Z}|}^{-1}(|2v-1|)\right)}{f_{Y}\left(F_{Y}^{-1}(u)\right)f_{|\tilde{Z}|}\left(F_{|\tilde{Z}|}^{-1}(|2v-1|)\right)}=\frac{f_{Y,\tilde{Z}}\left(F_{Y}^{-1}(u),F_{|\tilde{Z}|}^{-1}(|2v-1|)\right)}{f_{Y}\left(F_{Y}^{-1}(u)\right)f_{\tilde{Z}}\left(F_{|\tilde{Z}|}^{-1}(|2v-1|)\right)}.$
Since $F_{|\tilde{Z}|}(z)=2F_{\tilde{Z}}(z)-1$ for $z\geqslant 0$, it follows that $F_{|\tilde{Z}|}^{-1}(q)=F_{\tilde{Z}}^{-1}((1+q)/2)$ for $q\in[0,1]$. Therefore, for $v\leqslant 0.5$, we have
$F_{|\tilde{Z}|}^{-1}(|2v-1|)=F_{|\tilde{Z}|}^{-1}(1-2v)=F_{\tilde{Z}}^{-1}(1-v)=-F_{\tilde{Z}}^{-1}(v)$
while, for $v\geqslant 0.5$, we have $F_{|\tilde{Z}|}^{-1}(|2v-1|)=F_{|\tilde{Z}|}^{-1}(2v-1)=F_{\tilde{Z}}^{-1}(v)$. Using the h-symmetry of $f_{Y,\tilde{Z}}$ and the symmetry of $f_{\tilde{Z}}$ we infer that $c^{*}(u,|2v-1|)=\tilde{c}(u,v)$ where $\tilde{c}$ is the density of the copula of $(Y,\tilde{Z})^{\top}$. The identity $c^{*}(u,|2v-1|)=c(u,v)$ follows from the fact that $(Y,Z)^{\top}$ and $(Y,\tilde{Z})^{\top}$ have the same copula. ∎
Clearly, by putting both parts of Proposition 3 together, we see that the density of the copula of a random vector $(Y,Z)^{\top}$ that is jointly symmetric about $(a,b)$ satisfies $c(u,v)=c^{*}(|2u-1|,|2v-1|)$ where $c^{*}$ is the density of the copula of $(|Y-a|,|Z-b|)^{\top}$.
3 ARCH and GARCH copulas
3.1 ARCH and GARCH processes
We define a broad class of first-order GARCH-type models and establish nomenclature for some special cases.
Definition 3 (First-order GARCH processes).
  1. 1.
    A process of GARCH(1,1) type is a strictly stationary process $(X_{t})_{t\in\mathbb{Z}}$ satisfying equations of the form
    $X_{t}=\sigma_{t}\;\epsilon_{t}$ (4)
    where $(\epsilon_{t})_{t\in\mathbb{Z}}$ is a sequence of iid innovations, drawn from a continuous distribution with mean zero and variance one, and $(\sigma_{t})_{t\in\mathbb{Z}}$ is a sequence of variables satisfying $\sigma_{t}=\sigma(X_{t-1},\sigma_{t-1})$ where $\sigma:\mathbb{R}\times\mathbb{R}^{+}\to\mathbb{R}^{+}$ is a strictly positive-valued parametric volatility function.
  2. 2.
    A process of ARCH(1) type is a process of GARCH(1,1) type where the volatility function $\sigma(x,s)$ depends only on $x$.
  3. 3.
    The classical GARCH(1,1) process is a process of GARCH(1,1) type type in which the volatility function satisfies
    $\sigma(x,s)=\sqrt{\alpha_{0}+\alpha_{1}x^{2}+\beta_{1}s^{2}}$ (5)
    where $\alpha_{0}>0$, $\alpha_{1}\geqslant 0$, $\beta_{1}\geqslant 0$ and $\mathbb{E}(\ln(\alpha_{1}\epsilon_{1}^{2}+\beta_{1}))<0$ for an innovation $\epsilon_{1}$.
  4. 4.
    The classical ARCH(1) process is the classical GARCH(1,1) process with $\beta_{1}=0$.
  5. 5.
    A process of GARCH(1,1) type is said to have symmetric distribution, if the distribution of $\epsilon_{1}$ has a density that is symmetric around zero, and to have symmetric volatility if $\sigma(x,w)$ is an even function of $x$. It is said to be symmetric if it has symmetric distribution and symmetric volatility.
The mathematical properties of the classical GARCH(1,1) process have been widely studied and the conditions on the parameters in part 3 of Definition 3 are known to guarantee a strictly stationary solution to the equations (4(Nelson,, 1990; Bougerol and Picard,, 1992). For a finite-variance solution it is usual to assume the stronger condition that $\alpha_{1}+\beta_{1}<1$. It is known that the marginal or stationary distribution $F_{X}$ has regularly varying (power-law) tails and Mikosch and Stărică, (2000) showed that the tail index $\zeta$ may be calculated as the root of the equation $\mathbb{E}\left(\left(\alpha_{1}\epsilon_{1}^{2}+\beta_{1}\right)^{\zeta/2}\right)=1$. However, the exact functional forms of $F_{X}$ or its density $f_{X}$ are not available.
3.2 The first-order copula of GARCH(1,1)-type processes
Simple calculations show that the joint density of a pair of successive variables $(X_{1},X_{2})^{\top}$ in a process of GARCH(1,1) type is given by
$f_{X_{1},X_{2}}(x,y)=\int_{0}^{\infty}f_{X_{1},X_{2},\sigma_{1}}(x,y,s)\mathrm{d}s$ $=\int_{0}^{\infty}f_{X_{2}\mid X_{1},\sigma_{1}}(y\mid x,s)f_{X_{1}\mid\sigma_{1}}(x\mid s)f_{\sigma}(s)\mathrm{d}s$
$=\int_{0}^{\infty}\frac{1}{\sigma(x,s)}f_{\epsilon}\left(\frac{y}{\sigma(x,s)}\right)\frac{1}{s}f_{\epsilon}\left(\frac{x}{s}\right)f_{\sigma}(s)\mathrm{d}s$ (6)
where $f_{\sigma}$ is the marginal density of the volatility process $(\sigma_{t})_{t\in\mathbb{Z}}$ and $f_{\epsilon}$ is the density of the innovations. Applying (3), the density $c_{1}$ of the copula $C_{1}$ of $(X_{1},X_{2})^{\top}$ is thus
$c_{1}(u,v)=\frac{\int_{0}^{\infty}\frac{1}{s\sigma\left(F_{X}^{-1}(u),s\right)}f_{\epsilon}\left(\frac{F_{X}^{-1}(v)}{\sigma\left(F_{X}^{-1}(u),s\right)}\right)f_{\epsilon}\left(\frac{F_{X}^{-1}(u)}{s}\right)f_{\sigma}(s)\mathrm{d}s}{f_{X}\left(F_{X}^{-1}(u)\right)f_{X}\left(F_{X}^{-1}(v)\right)}$ (7)
where
$f_{X}(x)=\int_{0}^{\infty}f_{X_{1}\mid\sigma_{1}}(x\mid s)f_{\sigma}(s)\mathrm{d}s=\int_{0}^{\infty}\frac{1}{s}f_{\epsilon}\left(\frac{x}{s}\right)f_{\sigma}(s)\mathrm{d}s$ (8)
is the marginal density of the process $(X_{t})_{t\in\mathbb{Z}}$.
In the case of a process of ARCH(1) type (where $\sigma(x,s)$ only depends on $x$) these expressions simplify considerably and become
$f_{X_{1},X_{2}}(x,y)$ $=\frac{1}{\sigma(x)}f_{\epsilon}\left(\frac{y}{\sigma(x)}\right)f_{X}(x),\quad\quad c_{1}(u,v)=\frac{1}{\sigma\left(F_{X}^{-1}(u)\right)}f_{\epsilon}\left(\frac{F_{X}^{-1}(v)}{\sigma\left(F_{X}^{-1}(u)\right)}\right)\frac{1}{f_{X}(F_{X}^{-1}(v))}\,.$ (9)
The key insight from these calculations is the following.
Proposition 4.
  1. (i)
    In a GARCH(1,1)-type process with symmetric distribution, the distribution of $(X_{1},X_{2})^{\top}$ is h-symmetric about zero.
  2. (ii)
    In a symmetric GARCH(1,1)-type process, the distribution of $(X_{1},X_{2})^{\top}$ is jointly symmetric about the origin.
Proof.
We consider the joint density $f_{X_{1},X_{2}}(x,y)$ in (6).
  1. (i)
    It is clear that a symmetric innovation distribution implies that $f_{X_{1},X_{2}}(x,y)=f_{X_{1},X_{2}}(x,-y)$, meaning the random vector $(X_{1},X_{2})^{\top}$ is h-symmetric about zero.
  2. (ii)
    An even volatility function also implies that $f_{X_{1},X_{2}}(x,y)=f_{X_{1},X_{2}}(-x,y)$ and so $(X_{1},X_{2})^{\top}$ is v-symmetric about zero and hence radially and jointly symmetric about the origin.
In case (i) above we can apply Lemma 1 to infer that the copula $C_{1}$ is h-symmetric and we can apply Proposition 3 to infer that its density may be written in the form $c_{1}(u,v)=c_{X_{1},|X_{2}|}(u,|2v-1|)$ where $c_{X_{1},|X_{2}|}$ is the density of the copula of $(X_{1},|X_{2}|)^{\top}$. In case (ii) we infer that the copula $C_{1}$ is jointly symmetric and its density may be written in the form $c_{1}(u,v)=c_{|X_{1}|,|X_{2}|}(|2u-1|,|2v-1|)$ where $c_{|X_{1}|,|X_{2}|}$ is the density of the copula of $(|X_{1}|,|X_{2}|)^{\top}$.
Despite the strong symmetry revealed by Proposition 4, there is one important form of symmetry that symmetric processes of GARCH(1,1) type do not have in general, namely exchangeability. It is not true that $f_{X_{1},X_{2}}(x,y)=f_{X_{1},X_{2}}(y,x)$ for all $(x,y)^{\top}\in\mathbb{R}^{2}$ so the bivariate distributions of the pairs $(X_{1},X_{2})^{\top}$ and $(X_{2},X_{1})^{\top}$ differ. This implies that ARCH and GARCH are not time-reversible models. In general $f_{X_{2}\mid X_{1}}(y\mid x)\neq f_{X_{1}\mid X_{2}}(y\mid x)$ and thus, in an intuitive sense, they are models in which the predictive distribution of the future given the past differs from the “predictive distribution” of the past given the future. From an econometric point of view, this would be considered natural. The non-exchangeability carries over to the copula so that we have in general that $C_{1}(u,v)\neq C_{1}(v,u)$ on $[0,1]^{2}$. Despite this observation, it is notable that the majority of the copulas employed in vine copula models of financial time series are exchangeable ones.
3.3 Computational examples
In this section we use computational methods to approximate and plot the joint densities $f_{X_{1},X_{2}}$ and copula densities $c_{1}$ in a number of models of GARCH(1,1) type. We concentrate on processes of ARCH(1) type in our illustrations since their first-order Markov structure simplifies computation of the unknown marginal density $f_{X}$ and distribution function $F_{X}$, which are key ingredients in (9). Beginning with the equation
$f_{X}(x)=\int_{-\infty}^{\infty}f_{\epsilon}\left(\frac{x}{\sigma(y)}\right)\frac{1}{\sigma(y)}f_{X}(y)\mathrm{d}y,$ (10)
which follows easily from (9), we approximate the integral by a sum of weighted values of $f_{X}(y_{j})$ on a grid of points $\mathcal{Y}=\{y_{1},\ldots,y_{n}\}$; we can then solve a linear system for $f_{X}(y_{j}),y_{j}\in\mathcal{Y}$, which is accurate enough for our purposes.
We also consider the classical GARCH(1,1) model with Gaussian innovations. To analyse this process we use simulation and estimation to find a parametric model for the unknown marginal density $f_{\sigma}$ of the volatility process. We can then integrate out volatility in expressions like (6), (7) and (8) [^1] All details of the model for $f_{\sigma}$ and the discretization method for ARCH(1)-type processes can be found in the code that accompanies this paper..
The parameter values for all the models we consider are summarized in Table 1.
Process $\alpha_{0}$ $\alpha_{1}$ $\beta_{1}$ $\gamma_{1}$
ARCH(1) 0.4 0.6 0 0
ARCH(1) with leverage effect 0.4 0.3 0 0.4
GARCH(1,1) 0.1 0.3 0.6 0
GARCH(1,1) with leverage effect 0.1 0.1 0.6 0.4
Table 1: Parameter values for processes analysed in paper. The volatility function $\sigma(x,s)$ in all models satisfies $\sigma(x,s)^{2}=\alpha_{1}+(\alpha_{1}+\gamma_{1}I_{\{x<0\}})x^{2}+\beta_{1}s^{2}$. Innovation distributions used are either Gaussian, Student t with $\nu=4$ or $\nu=2.5$ degrees of freedom or skewed Student t with $\nu=4$ degrees of freedom and skewness parameter $\lambda=0.8$; all innovation distributions are scaled to have mean zero and variance one.
Figure 1: Contour plots of $f_{X_{1},X_{2}}(x_{1},x_{2})$ for symmetric GARCH(1,1)-type processes; see Table 1 for all parameter values. Top left, top right, bottom left: classical ARCH(1) processes with, respectively, Gaussian, Student t4 and Student t2.5 innovations. Bottom right: classical GARCH(1,1) process with Gaussian innovations.
Figure 2: Contour plots of $c_{1}(u,v)$ for symmetric GARCH(1,1)-type processes; see Table 1 for all parameter values. Top left, top right, bottom left: classical ARCH(1) processes with, respectively, Gaussian, Student t4 and Student t2.5 innovations. Bottom right: classical GARCH(1,1) process with Gaussian innovations.
  1. 1.
    Classical ARCH(1) process with Gaussian innovations (symmetric) and parameter values $\alpha_{0}=0.4$ and $\alpha_{1}=0.6$ as in Table 1; this process has unit variance and tail index $\zeta=3.82$. The joint density $f_{X_{1},X_{2}}$ and copula density $c_{1}$ are shown in the top left panels of Figures 1 and 2 respectively. The non-exchangeability, evidenced by a lack of symmetry in the dashed lines $x_{2}=x_{1}$ and $v=u$, is evident in both pictures. We can also observe the joint symmetry, that is the rotational symmetry of order 2 as well as the h-symmetry and v-symmetry.
    Changing $\alpha_{1}$ changes the degree of dependence in the pictures. In particular, as $\alpha_{1}\to 0$ the copula converges to the independence copula while increasing $\alpha_{1}$ serves to strengthen the cross-shaped character of the copula density. For normal innovations, the theoretical maximum value of $\alpha_{1}$ to obtain a strictly stationary process is approximately 3.56, but for values $\alpha_{1}\geqslant 1$ we obtain infinite-variance processes which are seldom used in practical applications and become increasingly pathological with increasing $\alpha_{1}$ as their tail indexes converge to zero (Mikosch and Stărică,, 2000).
  2. 2.
    Classical ARCH(1) process with Student innovations (symmetric). This is shown in the top right panels of Figures 1 and 2 for a process with the same parameters as in the previous example and degree of freedom $\nu=4$ for the innovation distribution; the tail index of the process is 2.76. Remarkably, the resulting joint density appears to be very close to spherical and the copula appears to be very close to an exchangeable copula for this combination of parameters; the joint symmetry is again apparent. The bottom left panels of Figures 1 and 2 show a process in which $\nu=2.5$ and the tail index is 2.09; these show non-exchangeability more clearly.
    It is notable that, although the tail index of the marginal distribution increases with decreasing degree of freedom (increasing tail index of the innovation density), the dependence actually weakens. We can estimate a distance $D_{1}=\int_{0}^{1}\int_{0}^{1}|c_{1}(u,v)-1|\mathrm{d}u\mathrm{d}v$ between the copula density $c_{1}$ and the density of the independence copula; for the Gaussian, t4 and t2.5 copulas we obtain respectively 0.213, 0.158 and 0.092.
  3. 3.
    Classical GARCH(1,1) process with Gaussian innovations (symmetric). This is shown in the bottom right panels of Figures 1 and 2 for a process with Gaussian innovations and parameters as in Table 1, which has variance one and tail index 4.09.
    It is well known that, for a covariance stationary GARCH(1,1) process ($\alpha_{1}+\beta_{1}<1$), the squared process $(X_{t}^{2})_{t\in\mathbb{Z}}$ behaves like an ARMA(1,1) process with autoregressive parameter $\alpha_{1}+\beta_{1}$ and moving-average parameter $-\beta_{1}$; see, for example, McNeil et al., (2015, Section 4.2.2). Thus the value of $\alpha_{1}+\beta_{1}$ controls the persistence, or the rate at which serial dependencies in the squared (or absolute value) process die away with lag. For a fixed value of $\alpha_{1}+\beta_{1}$ the magnitude of the serial dependency at lag 1 increases with $\alpha_{1}$, so that the ARCH(1) special case actually has the strongest dependence and the copula tends towards independence as $\alpha_{1}\to 0$. The parameters we have chosen are not necessarily typical of real financial data, where the ARCH effect $\alpha_{1}$ might be smaller and the GARCH effect $\beta_{1}$ larger, but have instead been selected to give a picture that has more evident dependence. The main point is to show that the behaviour of the GARCH(1,1) copula is qualitatively similar to that of the ARCH(1) copula.
  4. 4.
    ARCH(1) model with Gaussian innovations and leverage (symmetric distribution). Here we use a volatility function $\sigma(x)$ satisfying $\sigma(x)^{2}=\alpha_{0}+(\alpha_{1}+\gamma_{1}I_{\{x<0\}})x^{2}$ for $\gamma_{1}>0$ so that a negative return produces a larger value of the volatility than a positive return of the same magnitude. From now on, we concentrate on contour plots of the copula only, as shown in the top left panel of Figure 3. It can be seen that leverage removes radial symmetry and v-symmetry, but there is still h-symmetry (symmetry in the line $v=0.5$).
  5. 5.
    Classical ARCH(1) model with skewed Student innovations (asymmetric distribution but symmetric volatility). This is shown in the top right panel of Figure 3 for an innovation distribution that is skewed to the left (towards negative shocks). The skewness removes all the previous symmetries in the pictures.
  6. 6.
    Absolute value copula for classical ARCH(1) model with Gaussian innovations. This is shown in the bottom left of Figure 3 for the same process as in the top left of Figure 2. It can be thought of as a blown-up version of the top right quadrant of the latter and clearly shows non-exchangeability and stronger dependence in the upper tail. Note that this picture also illustrates the density of the volatility copula of ARCH(1), that is the copula of $(\sigma_{1},\sigma_{2})^{\top}$; this follows because $\sigma_{t}$ is a strictly increasing function of $|X_{t-1}|$ for all $t$ and copulas are invariant under strictly increasing transformations of variables. For the same reason, it is also the copula density for the variables $(X_{1}^{2},X_{2}^{2})^{\top}$.
Figure 3: Introducing asymmetry. Top left: ARCH(1) with Gaussian innovations in which the volatility function has leverage and the parameters are $\alpha_{0}=0.4$, $\alpha_{1}=0.3$ and $\gamma_{1}=0.4$. Top right: classical ARCH(1) model with with 4 degrees of freedom and skewness parameter $\lambda=0.8$ where $\alpha_{0}=0.4$ and $\alpha_{1}=0.6$. Bottom left: copula density for copula of the absolute values for classical ARCH(1) copula with Gaussian innovations. Bottom right: AR(1)-ARCH(1) model with Gaussian innovations and parameters $\phi=0.3$, $\alpha_{0}=0.4$ and $\alpha_{1}=0.6$.
3.4 AR(1) model with ARCH(1) errors
We consider briefly a time series model with stochastic volatility as well as a non-zero stochastic conditional mean term. We consider an AR(1) model with ARCH(1) errors and show the implications of the conditional mean term. Specifically, we consider the first-order Markov process $(X_{t})_{t\in\mathbb{Z}}$ satisfying equations
$X_{t}=\phi X_{t-1}+\sigma(X_{t-1})\;\epsilon_{t}\quad\text{with}\quad\sigma(x)=\sqrt{\alpha_{0}+\alpha_{1}x^{2}},$ (11)
where the innovations $(\epsilon_{t})_{t\in\mathbb{Z}}$ have a symmetric distribution. This model was proposed and analysed by Borkovec and Klüppelberg, (2001) [^2] It differs from the usual version of the AR(1)-ARCH(1) model, $X_{t}=\phi X_{t-1}+\sigma(X_{t-1}-\phi X_{t-2})\;\epsilon_{t}$, but this is not first-order Markov and thus less straightforward to treat with our methods. who showed that it is covariance stationary provided $\phi^{2}+\alpha_{1}<1$ and has a symmetric stationary distribution $F_{X}$.
The formulas (9) require only simple modification to account for the non-zero conditional mean term. In particular, the copula becomes
$c_{1}(u,v)=\frac{1}{\sigma\left(F_{X}^{-1}(u)\right)}f_{\epsilon}\left(\frac{F_{X}^{-1}(v)-\phi F_{X}^{-1}(u)}{\sigma\left(F_{X}^{-1}(u)\right)}\right)\frac{1}{f_{X}(F_{X}^{-1}(v))}\,.$
It may be easily verified that the copula density satisfies $c(u,v)=c(1-u,1-v)$ and so the copula is radially symmetric. However, the copula is neither h-symmetric nor v-symmetric, and thus it is not jointly symmetric. The copula density is shown in the bottom right panel of Figure 3 in the case where $\phi=0.3$, $\alpha_{0}=0.4$ and $\alpha_{1}=0.6$. It is evident that more of the probability mass is now on the main diagonal of the copula.
In the remainder of the paper we concentrate on models without a conditional mean term, but this example gives some clues about the effect on the copula of serial dependence in the untransformed process $(X_{t})_{t\in\mathbb{Z}}$ in combination with stochastic volatility, a point we return to in our discussion in Section 7.
3.5 Higher-order conditional copulas of GARCH(1,1)-type processes
For higher-order copulas we introduce the compact notation $\bm{X}_{[j,l]}=(X_{j},\ldots,X_{l})^{\top}$ to denote random vectors consisting of a contiguous block of random variables; we apply the same notation to realisations of the random vectors. Whenever the notation is used with $j>l$, this means that the variables are taken in reverse order to their natural temporal order.
Now we consider the copula $C_{k}$ of the conditional distribution of $(X_{1},X_{k+1})^{\top}$ given $\bm{X}_{[2:k]}=\bm{x}_{[2:k]}$ for $k\geqslant 2$. The density of the conditional copula takes the form
$c_{k}(u,v\mid\bm{x}_{[2:k]})=\frac{f_{X_{1},X_{k+1}\mid\bm{X}_{[2:k]}}\left(F_{X_{1}\mid\bm{X}_{[2:k]}}^{-1}\left(u\mid\bm{x}_{[2:k]}\right),F_{X_{k+1}\mid\bm{X}_{[2:k]}}^{-1}\left(v\mid\bm{x}_{[2:k]}\right)\mid\bm{x}_{[2:k]}\right)}{f_{X_{1}\mid\bm{X}_{[2:k]}}\left(F_{X_{1}\mid\bm{X}_{[2:k]}}^{-1}\left(u\mid\bm{x}_{[2:k]}\right)\mid\bm{x}_{[2:k]}\right)f_{X_{k+1}\mid\bm{X}_{[2:k]}}\left(F_{X_{k+1}\mid\bm{X}_{[2:k]}}^{-1}\left(v\mid\bm{x}_{[2:k]}\right)\mid\bm{x}_{[2:k]}\right)}$
Note that this copula density will simply be the independence copula density for a process of ARCH(1) type, since $X_{1}$ and $X_{3}$ are conditionally independent given $X_{2}$. Consider, therefore a process with an explicit GARCH effect, i.e. $\sigma(x,s)$ depends also on $s$.
To develop intuition we consider the case where $k=2$. The symmetry properties can be investigated by considering the joint conditional density $f_{X_{1},X_{3}\mid X_{2}}(x,y\mid z)$. To this end consider the joint density
$f_{X_{1},X_{2},X_{3}}(x,z,y)$ $=\int_{0}^{\infty}f_{X_{1},X_{2},X_{3},\sigma_{1}}(x,z,y,s)\mathrm{d}s$
$=\int_{0}^{\infty}f_{X_{3}\mid X_{2},X_{1},\sigma_{1}}(y\mid z,x,s)f_{X_{2}\mid X_{1},\sigma_{1}}(z\mid x,s)f_{X_{1}\mid\sigma_{1}}(x\mid s)f_{\sigma}(s)\mathrm{d}s$
$=\int_{0}^{\infty}\frac{1}{\sigma(z,\sigma(x,s))}f_{\epsilon}\left(\frac{y}{\sigma(z,\sigma(x,s))}\right)\frac{1}{\sigma(x,s)}f_{\epsilon}\left(\frac{z}{\sigma(x,s)}\right)\frac{1}{s}f_{\epsilon}\left(\frac{x}{s}\right)f_{\sigma}(s)\mathrm{d}s.$
It is straightforward to see that a process with symmetric innovations will satisfy $f_{X_{1},X_{2},X_{3}}(x,z,-y)=f_{X_{1},X_{3},X_{2}}(x,z,y)$ while a symmetric process will also satisfy $f_{X_{1},X_{2},X_{3}}(-x,z,y)=f_{X_{1},X_{2},X_{3}}(x,z,y)$. These properties will also hold for the conditional density $f_{X_{1},X_{3}\mid X_{2}}(x,y\mid z)=f_{X_{1},X_{2},X_{3}}(x,z,y)/f_{X}(z)$ and thus, by Lemma 1, the various forms of symmetry will be inherited by the conditional copula density $c_{2}(u,v\mid z)$. By applying a similar analysis to the joint density $f_{X_{1},\ldots,X_{k+1}}(x,\ldots,y)$ we can therefore establish the following result.
Proposition 5.
  1. (i)
    In a GARCH(1,1)-type process with symmetric distribution, the conditional copula $C_{k}$ is h-symmetric about zero for $k\geqslant 2$.
  2. (ii)
    In a symmetric GARCH(1,1)-type process the conditional copula is jointly symmetric about the origin for $k\geqslant 2$.
Proof.
Suppose we write $s_{2}(x,s)=\sigma(x,s)$, $s_{3}(z_{1},x,s)=\sigma(z_{1},s_{2}(x,s))$ and then for $j=4,\ldots,k+1$ we define recursive functions $s_{j}(z_{j-2},\ldots,z_{1},x,s)=\sigma(z_{j-2},s_{j-1}(z_{j-3},\ldots,z_{1},x,s))$. This permits us to write the joint density as
$f_{X_{1},\ldots,X_{k+1}}(x,z_{1},\ldots,z_{k-1},y)=\int_{0}^{\infty}f_{X_{k+1}\mid\sigma_{k+1}}(y\mid s_{k+1}(z_{k-1},\ldots,z_{1},x,s))\\ \left(\prod_{j=3}^{k}f_{X_{j}\mid\sigma_{j}}(z_{j-1}\mid s_{j}(z_{j-2},\ldots,z_{1},x,s))\right)f_{X_{2}\mid\sigma_{2}}(z_{1}\mid s_{2}(x,s))f_{X_{1}\mid\sigma_{1}}(x\mid s)f_{\sigma}(s)\mathrm{d}s$
It is clear that this joint density satisfies $f_{X_{1},\ldots,X_{k+1}}(x,\ldots,y)=f_{X_{1},\ldots,X_{k+1}}(x,\ldots,-y)$ when the innovation distribution is symmetric. Moreover it is also clear that in the symmetric model we have $f_{X_{1},\ldots,X_{k+1}}(x,\ldots,y)=f_{X_{1},\ldots,X_{k+1}}(-x,\ldots,y)$ since all of the functions $s_{2},\ldots,s_{k+1}$ are even functions of $x$. These properties will also hold for the conditional density
$f_{X_{1},X_{k+1}\mid\bm{X}_{[2:k]}}(x,y\mid z_{1},\ldots,z_{k-1})=f_{X_{1},\ldots,X_{k+1}}(x,z_{1},\ldots,z_{k-1},y)/f_{X_{2},\ldots,X_{k}}(z_{1},\ldots,z_{k-1})$
and thus we can apply Lemma 1 to infer that the assertions regarding symmetry of the copula hold. ∎
Remark 1.
Since the joint density of $(X_{1},X_{k+1})^{\top}$ can be obtained as
$f_{X_{1},X_{k+1}}(x,y)=\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}f_{X_{1},\ldots,X_{k+1}}(x,z_{1},\ldots,z_{k-1},y)\mathrm{d}z_{1}\cdots\mathrm{d}z_{k-1}$
it follows that the same method of proof can be used to show that the copula of the unconditional distribution of $(X_{1},X_{k+1})^{\top}$ shares the same symmetry properties as the conditional copula.
An interesting question concerns the degree to which $c_{k}(u,v\mid\cdot)$ depends on the realized values of the conditioning variables. We investigate this question for $k=2$ and and write $c_{2}(u,v;w):=c_{2}(u,v\mid F_{X}^{-1}(w))$. The latter function takes the form
$c_{2}(u,v;w)=\frac{f_{X_{1},X_{3}\mid X_{2}}\left(F_{X_{1}\mid X_{2}}^{-1}\left(u\mid F_{X}^{-1}(w)\right),F_{X_{3}\mid X_{2}}^{-1}\left(v\mid F_{X}^{-1}(w)\right)\mid F_{X}^{-1}(w)\right)}{f_{X_{1}\mid X_{2}}\left(F_{X_{1}\mid X_{2}}^{-1}\left(u\mid F_{X}^{-1}(w)\right)\mid F_{X}^{-1}(w)\right)f_{X_{3}\mid X_{2}}\left(F_{X_{3}\mid X_{2}}^{-1}\left(v\mid F_{X}^{-1}(w)\right)\mid F_{X}^{-1}(w)\right)}$
and can be calculated approximately for the classical GARCH(1,1) model building on our analysis in Section 3.3.
In Figure 4 we show contour plots for the cases where $w=0.5$, i.e. the intermediate observation is at the median of the marginal distribution (zero in this case), and $w=0.75$, i.e. the intermediate observation is at a quartile of the marginal distribution. While the two pictures are not identical, they are relatively similar to each other in terms of the shapes and heights of the contour lines. To examine this further we attempt to compute an approximate value for the absolute distance between the conditional copula densities and the independence copula density using the measure $D_{2}(w)=\int_{0}^{1}\int_{0}^{1}|c_{2}(u,v;w)-1|\mathrm{d}u\mathrm{d}v$. We obtain the estimates $D_{2}(0.5)=0.10$ and $D_{2}(0.75)=0.09$, which can be compared with the equivalent measure $D_{1}=0.16$ for the copula density $c_{1}$ in Figure 2.
As might be expected, the dependence weakens from lag 1 to lag 2. Moreover, the value of the conditioning variable does not seem to have a particularly large effect on the conditional copula at lag 2. This might be taken as weak evidence that the simplifying assumption is approximately tenable; obviously we should look at values of $w$ further into the tail, but our approximation methods then become more unreliable.
Figure 4: Contour plots of $c_{2}(u,v;w)$ for the classical GARCH(1,1) process with Gaussian innovations. In the left-hand figure $w=0.5$ and in the right-hand figure $w=0.75$.
4 Copula densities and v-transforms
4.1 V-transforms and their inverses
The function $\mathcal{V}(u)=|2u-1|$ which appeared in Proposition 3 is a symmetric v-shaped function and the simplest example of a v-transform. These are special cases of uniform-measure-preserving transformations (Porubský et al.,, 1988) and are extensively analysed in McNeil, (2021) and Bladt and McNeil, 2022b . In this section we recap the essential elements of the theory and examine its relevance for understanding and mimicking the behaviour of the copula of processes of GARCH(1,1) type.
Definition 4.
A mapping $\mathcal{V}:[0,1]\to[0,1]$ is a v-transform if it takes the form
$\mathcal{V}(u)=\begin{cases}(1-u)-(1-\delta)\Psi\left(\frac{u}{\delta}\right)&u\leqslant\delta,\\ u-\delta\Psi^{-1}\left(\frac{1-u}{1-\delta}\right)&u>\delta,\end{cases}$ (12)
for a parameter $10<\delta<$ 1, known as the fulcrum, and a continuous and strictly increasing distribution function $\Psi$ on $[0,1]$, known as the generator.
Mappings of this form are always v-shaped with $\mathcal{V}(0)=\mathcal{V}(1)=1$ and $\mathcal{V}(\delta)=0$. They preserve uniformity in the sense that if $U\sim U(0,1)$, then $\mathcal{V}(U)\sim U(0,1)$. If the generator $\Psi$ is taken to be standard uniform, we obtain the linear v-transform, which satisfies $\mathcal{V}(u;\delta)=(\delta-u)/\delta$ for $u\leqslant\delta$ and $\mathcal{V}(u;\delta)=(u-\delta)/(1-\delta)$ for $u>\delta$; this includes the special case $\mathcal{V}(u;0.5)=|2u-1|$, which is the symmetric v-transform. It is possible to introduce further parameters. For example, by taking $\Psi(x)=x^{\kappa}$ for $\kappa>0$, we obtain a two-parameter family
$\mathcal{V}(u;\delta,\kappa)=\begin{cases}(1-u)-(1-\delta)\left(\frac{u}{\delta}\right)^{\kappa}&u\leqslant\delta,\\ u-\delta\left(\frac{1-u}{1-\delta}\right)^{1/\kappa}&u>\delta.\end{cases}$ (13)
For further results we need some more notation. For every $y\in(0,1]$ there are two possible values $u$ on either side of the fulcrum $\delta$ such that $\mathcal{V}(u)=y$ and these may be written as $\mathcal{V}^{-1}(y)$ and $\mathcal{V}^{-1}(y)+y$ where $\mathcal{V}^{-1}(y)=\inf\{u:\mathcal{V}(u)=y\}$ is a partial inverse of $\mathcal{V}$. We also introduce the function $\Delta(y)=-1/\mathcal{V}^{\prime}(\mathcal{V}^{-1}(y))$ where $\mathcal{V}^{\prime}$ is the derivative of $\mathcal{V}$. If $U\sim U(0,1)$ then McNeil, (2021) shows that $\Delta(y)=\mathbb{P}(U\leqslant\delta\mid\mathcal{V}(U)=y)$ for $y\in(0,1]$; for the linear v-transform $\Delta(y)=\delta$, but otherwise $\Delta(y)$ depends explicitly on $y$.
We can stochastically invert a v-transform in such a way that we also preserve uniformity of random variables. For $y\in(0,1]$, the stochastic inverse $\mathcal{V}^{\leftarrow}(y)$ of the v-transform $\mathcal{V}$ is defined to take the value $\mathcal{V}^{-1}(y)$ (from the left branch of the transform) with probability $\Delta(y)$ and the value $\mathcal{V}^{-1}(y)+y$ (from the right branch of the transform) with probability $1-\Delta(y)$; for $y=0$ we set $\mathcal{V}^{\leftarrow}(y)=\delta$. If $U\sim U(0,1)$ then it can also be shown that $\mathcal{V}^{\leftarrow}(U)\sim U(0,1)$. Note, however, that while $\mathcal{V}\circ\mathcal{V}^{\leftarrow}(y)=y$ holds for all $y\in[0,1]$, the equation $\mathcal{V}^{\leftarrow}\circ\mathcal{V}(u)=u$ is only true with probability $\Delta(u)$, if $u<\delta$, and with probability $1-\Delta(u)$, if $u>\delta$.
4.2 V-transformation and inverse v-transformation of copulas
The following result, which is an adaptation of a result in McNeil, (2021), is key to constructing new copulas using v-transforms and their inverses.
Theorem 1.
Let $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ be v-transforms with associated partial inverses $\mathcal{V}_{1}^{-1}$ and $\mathcal{V}_{2}^{-1}$ and $\Delta$-functions $\Delta_{1}$ and $\Delta_{2}$. Let $(U,V)^{\top}$ be distributed according to a copula $C_{U,V}$ with density $c_{U,V}$.
  1. (i)
    $(\mathcal{V}_{1}(U),\mathcal{V}_{2}(V))^{\top}$ is distributed according to a copula $C_{\mathcal{V}_{1}(U),\mathcal{V}_{2}(V)}$ with density
    $c_{\mathcal{V}_{1}(U),\mathcal{V}_{2}(V)}(u,v)=\;$ $c_{U,V}(\mathcal{V}_{1}^{-1}(u),\mathcal{V}_{2}^{-1}(v))\;\Delta_{1}(u)\,\Delta_{2}(v)\;+$
    $c_{U,V}(\mathcal{V}_{1}^{-1}(u)+u,\mathcal{V}_{2}^{-1}(v))\;\big{(}1-\Delta_{1}(u)\big{)}\,\Delta_{2}(v)\;+$ (14)
    $c_{U,V}(\mathcal{V}_{1}^{-1}(u),\mathcal{V}_{2}^{-1}(v)+v)\;\Delta_{1}(u)\,\big{(}1-\Delta_{2}(v)\big{)}\;+$
    $c_{U,V}(\mathcal{V}_{1}^{-1}(u)+u,\mathcal{V}_{2}^{-1}(v)+v)\;\big{(}1-\Delta_{1}(u)\big{)}\,\big{(}1-\Delta_{2}(v)\big{)}.$
  2. (ii)
    $(\mathcal{V}_{1}^{\leftarrow}(U),\mathcal{V}_{2}^{\leftarrow}(V))^{\top}$ is distributed according to a copula $C_{\mathcal{V}^{\leftarrow}_{1}(U),\mathcal{V}^{\leftarrow}_{2}(V)}$ with density
    $c_{\mathcal{V}_{1}^{\leftarrow}(U),\mathcal{V}_{2}^{\leftarrow}(V)}(u,v)=c_{U,V}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v)).$ (15)
Proof.
This may be proved by a simple adaptation of the proof of Theorem 3 in McNeil, (2021) which deals with the case of two (or more) identical v-transforms. ∎
We refer to the copula density $c_{\mathcal{V}_{1}(U),\mathcal{V}_{2}(V)}$ in part (i) as a v-transformed version of $c_{U,V}$ while we refer to the copula density $c_{\mathcal{V}_{1}^{\leftarrow}(U),\mathcal{V}_{2}^{\leftarrow}(V)}$ in part (ii) as an inverse-v-transformed version of $c_{U,V}$. Inverse v-transformation of a copula density using (15) followed by v-transformation using (14) returns the original copula density $c_{U,V}$ (since for any v-transform $\mathcal{V}\circ\mathcal{V}^{\leftarrow}(y)=y$ for $y\in[0,1]$). However v-transformation of a copula density using (14) followed by inverse v-transformation using (15) does not, in general, return the original density $c$, except in special cases, such as the following.
Corollary 1.
If $C_{U,V}$ is a jointly symmetric copula and $\mathcal{V}_{1}(u)=\mathcal{V}_{2}(u)=|2u-1|$ for $u\in[0,1]$, then the density of the copula $C_{\mathcal{V}_{1}(U),\mathcal{V}_{2}(V)}$ satisfies
$c_{\mathcal{V}_{1}(U),\mathcal{V}_{2}(V)}(u,v)$ $=$ $c_{U,V}\left(\frac{1-u}{2},\frac{1-v}{2}\right)=c_{U,V}\left(\frac{1-u}{2},\frac{1+v}{2}\right)$ (16)
$=$ $c_{U,V}\left(\frac{1+u}{2},\frac{1-v}{2}\right)=c_{U,V}\left(\frac{1+u}{2},\frac{1+v}{2}\right)$
and
$c_{\mathcal{V}_{1}^{\leftarrow}\circ\mathcal{V}_{1}(U),\mathcal{V}_{2}^{\leftarrow}\circ\mathcal{V}_{2}(V)}(u,v)=c_{\mathcal{V}_{1}(U),\mathcal{V}_{2}(V)}(|2u-1|,|2v-1|)=c_{U,V}(u,v).$ (17)
Proof.
By part (i) of Theorem 1 and equation (14), the copula density satisfies
$c_{\mathcal{V}_{1}(U),\mathcal{V}_{2}(V)}(u,v)$ $=\frac{1}{4}\left(c_{U,V}\left(\frac{1-u}{2},\frac{1-v}{2}\right)+c_{U,V}\left(\frac{1-u}{2},\frac{1+v}{2}\right)\right.$
$\left.+c_{U,V}\left(\frac{1+u}{2},\frac{1-v}{2}\right)+c_{U,V}\left(\frac{1+u}{2},\frac{1+v}{2}\right)\right)$
and (16) follows easily from the joint symmetry property of $C_{U,V}$. The identity (17) is then easily verified by considering the values of $|2u-1|$ and $|2v-1|$ in the four quadrants of the unit square defined by the lines $u=0.5$ and $v=0.5$ and using each of the four equations in (16) in turn. ∎
Remark 2.
We can easily derive a corresponding result for the simpler case where the copula $C_{U,V}$ has h-symmetry and $V$ is v-transformed with the symmetric v-transform. The copula density of the random vector $(U,|2V-1|))^{\top}$ can be shown to satisfy
$c_{U,|2V-1|}(u,v)=c_{U,V}\left(u,\frac{1-v}{2}\right)=c_{U,V}\left(u,\frac{1+v}{2}\right)$ (18)
and it is easily verified that $c_{U,|2V-1|}(u,|2v-1|)=c_{U,V}(u,v)$.
Remark 3.
To see that the identity (17) does not always hold when $\mathcal{V}_{1}(u)=\mathcal{V}_{2}(u)=|2u-1|$, it suffices to start with any copula $C_{U,V}$ that is not jointly symmetric, since application of the inverse v-transform in the final step always results in a jointly symmetric copula. For example, when $C_{U,V}=C^{\text{Ga}}_{\rho}$, the Gaussian copula with parameter $\rho>0$, v-transformation followed by inverse v-transformation results in the jointly symmetric mixture $0.5C^{\text{Ga}}_{\rho}+0.5C^{\text{Ga}}_{-\rho}$.
Corollary 1 and Remark 2 allow us to state the following simple result, which is relevant to our study of the symmetries in the densities of copulas of GARCH(1,1) type.
Proposition 6.
Let $C$ be an absolutely continuous copula with density $c$.
  1. (i)
    $C$ is h-symmetric if and only if $c(u,v)=c^{*}(u,|2v-1|)$ for some copula density $c^{*}$.
  2. (ii)
    $C$ is jointly symmetric if and only if $c(u,v)=c^{*}(|2u-1|,|2v-1|)$ for some copula density $c^{*}$.
Proof.
It is trivial to verify that copulas with the structures $c^{*}(u,|2v-1|)$ and $c^{*}(|2u-1|,|2v-1|)$ are h-symmetric and jointly symmetric, respectively. To show the reverse implication in part (i), we choose $c^{*}$ to be the copula density of $(U,|2V-1|)^{\top}$ where $(U,V)^{\top}$ is a random vector with copula density $c$. To show the reverse implication in part (ii), we choose $c^{*}$ to be the copula density of $(|2U-1|,|2V-1|)^{\top}$ where $(U,V)^{\top}$ is a random vector with copula density $c$. ∎
4.3 Implicit v-transforms in general bivariate densities
In this section we use part (i) of Theorem 1 to find a condition for our generic random vector $(Y,Z)^{\top}$ on $\mathbb{R}^{2}$ which allows us to write its copula density $c$ in the general form $c(u,v)=c^{*}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v))$ for v-transforms $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ and a copula density $c^{*}$. This result requires the concept of a v-shaped transformation on $\mathbb{R}$, which should be distinguished from a v-transform.
Definition 5.
A function $T$ is a v-shaped function on $\mathbb{R}$ if it can be written in the form
$T(x)=\begin{cases}T_{L}(\mu_{T}-x)&\quad x\leqslant\mu_{T}\\ T_{R}(x-\mu_{T})&\quad x>\mu_{T}\end{cases}$
where $T_{L}$ and $T_{R}$ are strictly increasing, continuous and differentiable functions on $\mathbb{R}^{+}=[0,\infty)$ such that $T_{L}(0)=T_{R}(0)$ and $\mu_{T}\in\mathbb{R}$.
Theorem 2.
Let $(Y,Z)^{\top}$ be a random vector in $\mathbb{R}^{2}$ with joint density $f_{Y,Z}$, marginals densities $f_{Y}$ and $f_{Z}$ and copula density $c$. If the density ratio $f_{Y,Z}(y,z)/(f_{Y}(y)f_{Z}(z))$ can be written in the form
$\frac{f_{Y,Z}(y,z)}{f_{Y}(y)f_{Z}(z)}=g\left(T_{1}(y),T_{2}(z)\right)$
for two v-shaped transformations $T_{1}$ and $T_{2}$ on $\mathbb{R}$ and some function $g:\mathbb{R}\times\mathbb{R}\to\mathbb{R}^{+}$, then we may write $c(u,v)=c^{*}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v))$ on $[0,1]^{2}$ where $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ are v-transforms given by $\mathcal{V}_{1}(u)=F_{T_{1}(Y)}\circ T_{1}\circ F_{Y}^{-1}(u)$ and $\mathcal{V}_{2}(v)=F_{T_{2}(Z)}\circ T_{2}\circ F_{Z}^{-1}(v)$ and where $c^{*}$ is the density of the copula of $(T_{1}(Y),T_{2}(Z))^{\top}$.
Proof.
The density $c$ is the density of the df of the random variables $(U,V)^{\top}$ where $U=F_{Y}(Y)$ and $V=F_{Z}(Z)$. In McNeil, (2021) it is shown that $F_{T_{1}(Y)}(T_{1}(Y))=\mathcal{V}_{1}(U)$ for the v-transform given by $\mathcal{V}_{1}=F_{T_{1}(Y)}\circ T_{1}\circ F_{Y}^{-1}$; similarly $F_{T_{2}(Z)}(T_{2}(Z))=\mathcal{V}_{2}(V)$ for $\mathcal{V}_{2}=F_{T_{2}(Z)}\circ T_{2}\circ F_{Z}^{-1}$. Hence we may write
$c(u,v)=g\left(T_{1}(F_{Y}^{-1}(u)),T_{2}(F_{Z}^{-1}(v))\right)=g\left(F_{T_{1}(Y)}^{-1}(\mathcal{V}_{1}(u)),F_{T_{2}(Z)}^{-1}(\mathcal{V}_{2}(v))\right).$ (19)
Let $c^{*}$ be the density of the df of $(\mathcal{V}_{1}(U),\mathcal{V}_{2}(V))^{\top}$ and thus the density of the copula of $(T_{1}(Y),T_{2}(Z))^{\prime}$. We can compute $c^{*}$ from (19) using (14). Since $\mathcal{V}_{1}(\mathcal{V}_{1}^{-1}(u))=\mathcal{V}_{1}(\mathcal{V}_{1}^{-1}(u)+u)=u$ and $\mathcal{V}_{2}(\mathcal{V}_{2}^{-1}(v))=\mathcal{V}_{2}(\mathcal{V}_{2}^{-1}(v)+v)=v$ we infer that $c^{*}(u,v)=g\left(F_{T_{1}(Y)}^{-1}(u),F_{T_{2}(Z)}^{-1}(v)\right)$ and hence it follows that $c(u,v)=c^{*}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v))$. ∎
We give two examples to illustrate the use of Theorem 2.
Example 1 (Symmetric GARCH-type process).
Theorem 2 gives an alternative proof of the fact that the copula density $c_{1}$ in a symmetric GARCH-type process can be written in the form $c^{*}(|2u-1|,|2v-1|)$ where $c^{*}$ is the density of the copula of $(|X_{1}|,|X_{2}|)^{\top}$. The joint density in (6) may be written as
$f_{X_{1},X_{2}}(x,y)=\int_{0}^{\infty}\frac{1}{w\sigma\left(|x|,s\right)}f_{\epsilon}\left(\frac{|y|}{\sigma\left(|x|,s\right)}\right)f_{\epsilon}\left(\frac{|x|}{s}\right)f_{\sigma}(s)\mathrm{d}s$
in terms of the absolute value function, which is a v-shaped transformation on $\mathbb{R}$; moreover the marginal distribution satisfies $f_{X}(x)=f_{X}(|x|)$.
Example 2 (ARCH(1) process with symmetric innovations and leverage).
For the process of ARCH(1) type with leverage and symmetric distribution in the upper left of Figure 3 we observe that the volatility function $\sigma(x)$ is itself a v-shaped transformation with $\mu_{\sigma}=0$, $\sigma_{L}(x)=\sqrt{\alpha_{0}+(\alpha_{1}+\gamma_{1})x^{2}}$ and $\sigma_{R}(x)=\sqrt{\alpha_{0}+\alpha_{1}x^{2}}$. In this case, it follows from the joint density in (9) that we may write
$\frac{f_{X_{1},X_{2}}(x,y)}{f_{X}(x)f_{X}(y)}=\frac{f_{\epsilon}\left(\frac{|y|}{\sigma(x)}\right)}{\sigma(x)f_{X}(|y|)}=g(\sigma(x),|y|)$
and conclude that the copula density satisfies $c_{1}(u,v)=c_{1}^{*}(\mathcal{V}_{1}(u),|2v-1|)$ where $c_{1}^{*}$ is the copula density of $(\sigma(X_{1}),|X_{2}|)^{\top}$ or, in other words, $(\sigma_{t},|X_{t}|)^{\top}$ for any $t$. To derive $\mathcal{V}_{1}$ we must find the composition of functions $\mathcal{V}_{1}(u)=F_{\sigma}\circ\sigma\circ F_{X}^{-1}(u)$. By using the fact that the distribution function of $\sigma_{t}$ is
$F_{\sigma}(s)=F_{X}\left(\sqrt{\frac{s^{2}-\alpha_{0}}{\alpha_{1}}}\right)-F_{X}\left(-\sqrt{\frac{s^{2}-\alpha_{0}}{\alpha_{1}+\gamma_{1}}}\right),$
we obtain after some calculation that
$\mathcal{V}_{1}(u)=\begin{cases}F_{X}\left(-\left(1+\frac{\gamma_{1}}{\alpha_{1}}\right)^{\frac{1}{2}}F_{X}^{-1}(u)\right)-u,&u\leqslant 0.5,\\ u-F_{X}\left(-\left(1+\frac{\gamma_{1}}{\alpha_{1}}\right)^{-\frac{1}{2}}F_{X}^{-1}(u)\right),&u>0.5,\end{cases}$
which is an asymmetric function for $\gamma_{1}\neq 0$.
For the classical ARCH(1) process with asymmetric innovation density in the upper right panel of Figure 3 it does not seem to be generally possible to write the copula density in the form $c_{1}(u,v)=c_{1}^{*}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v))$. Although the density ratio can be written in the form
$\frac{f_{X_{1},X_{2}}(x,y)}{f_{X}(x)f_{X}(y)}=\frac{f_{\epsilon}\left(\frac{y}{\sigma(x)}\right)}{\sigma(x)f_{X}(y)}=h(\sigma(x),y),$
it does not appear to be possible to find a second v-shaped transformation $T_{2}$ such that $h(\sigma(x),y)=g(\sigma(x),T_{2}(y))$. This conclusion extends to the more general models of first-order GARCH type with asymmetric innovation distributions.
Despite the negative conclusion, copula densities of the form $c(u,v)=c^{*}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v))$ are still promising candidates for approximating or mimicking the implicit copulas of GARCH(1,1)-type models and this is the subject of the next section.
5 GARCH-mimicking copulas
5.1 Mimicking jointly symmetric GARCH copulas
There are a number of ways of finding jointly symmetric copulas that might be capable of mimicking the jointly symmetric copula of a symmetric GARCH process. We could consider:
  1. 1.
    the copula of a spherically symmetric distribution;
  2. 2.
    an equal (50/50) mixture of a radially symmetric copula and its 90-degree rotation;
  3. 3.
    an equal (25/25/25/25) mixture of any copula and its rotations through 90, 180 and 270 degrees;
  4. 4.
    a copula constructed from another copula $C^{*}$ by inverse v-transformation with the symmetric v-transform, i.e. a copula with density of the form $c^{*}(|2u-1|,|2v-1|)$.
In view of Proposition 6, the densities of the copulas arising from the first 3 approaches can also be written in the form $c(u,v)=c^{*}(|2u-1|,|2v-1|)$ for some implicit copula $C^{*}$ with density $c^{*}$, so the 4th approach should be considered as the overarching approach. Proposition 3, part (ii), and the lower-left panel of Figure 3 give clues about the kind of copula density $c^{*}$ we should be targeting. In particular, we note that there should be strong dependence in the upper right tail and some asymmetry.
An example of (1) is the copula of a spherical t distribution, i.e. a bivariate t distribution located at the origin with identical marginal distributions and correlation equal to zero. We call this copula the spherical t copula and denote it $C^{t}_{\nu}$. Since a spherical t-distributed random vector $(Y,Z)^{\top}$ is jointly symmetric, so is its copula. As the degree-of-freedom parameter tends to zero, this copula becomes increasingly cross-shaped and converges to an equal mixture of the comonotonicity and countermonotonicity copulas (McNeil et al.,, 2022).
It is of interest to identify the copula corresponding to v-transformation of $C^{t}_{\nu}$, in other words, the copula of $(|Y|,|Z|)^{\top}$. We call this copula the absolute spherical t copula (ast) and denote it $C^{|t|}_{\nu}$. By Corollary 1 and (16) its density must satisfy
$c^{|t|}_{\nu}(u,v)=c^{t}_{\nu}\left(\frac{1-u}{2},\frac{1-v}{2}\right)=c^{t}_{\nu}\left(\frac{1-u}{2},\frac{1+v}{2}\right)=c^{t}_{\nu}\left(\frac{1+u}{2},\frac{1-v}{2}\right)=c^{t}_{\nu}\left(\frac{1+u}{2},\frac{1+v}{2}\right).$
Integrating any one of these equalities and observing the boundary conditions for a copula gives us an expression for $C^{|t|}_{\nu}$. For example, integrating the final equality gives
$C^{|t|}_{\nu}=4C_{\nu}^{t}\left(\frac{1+u}{2},\frac{1+v}{2}\right)-u-v-1.$ (20)
This one-parameter copula, which appears to be new, interpolates between independence (as $\nu\to\infty$) and comonotonicity (as $\nu\to 0$). For small degrees of freedom, the density $c^{|t|}_{\nu}$ does bear some resemblance to the lower-left panel of Figure 3, but cannot reproduce the non-exchangeability.
An example of (2) would be an equal mixture of the t copula $C^{t}_{\nu,\rho}$ and the t copula $C^{t}_{\nu,-\rho}$ for some $\rho\geqslant 0$; obviously this contains the spherical t copula $C^{t}_{\nu}$ as the special case where $\rho=0$. Examples of (3) could be constructed using the Gumbel, Clayton or Joe copulas, to take advantage of the strong tail dependence of these copulas in either the upper-right or lower-left corners; this includes the equally-weighted special case of the so-called convex Gumbel copula (Junker and May,, 2005) used by Loaiza-Maya et al., (2018) for financial return data. Mixtures of types (2) and (3) all imply v-transformed copulas $C^{*}$, whose forms may be easily derived using the method we have used for the spherical t copula. If we follow the explicit construction (4), we need to choose a copula $C^{*}$ that has stronger dependence in the upper tail than the lower tail. This suggests considering the Gumbel, Joe or 180-degree-rotated (survival) Clayton copulas.
So far none of these constructions address the issue of non-exchangeability which was clearly evident in certain examples in Figure 2. One possible construction for non-exchangeable copulas is the one proposed by Khoudraji, (1995) and later generalized by Liebscher, (2008). In this construction two copulas $C_{A}$ and $C_{B}$ are combined to give a new copula
$C_{K}(u,v)=C_{A}(u^{1-a_{1}},v^{1-a_{2}})C_{B}(u^{a_{1}},v^{a_{2}})$ (21)
by choosing constants $a_{i}\in[0,1]$ for $i=1,2$ and the resulting copula $C_{K}$ is non-exchangeable when $a_{1}\neq a_{2}$. In our use of this construction we set $C_{B}$ to be the independence copula. We apply the Khoudraji extension to copulas before either mixing them with their rotations or applying inverse v-transformation.
We note, however, that the class of copulas in (21) does have one problematic feature for maximum-likelihood inference; even when $C_{B}$ is the independence copula, in order to evaluate the copula density of $c_{K}$ of $C_{K}$, we need to be able to evaluate, not only the copula density $c_{A}$, but also the copula $C_{A}$ and its partial derivatives. This is a particular challenge for copulas related to the t copula [^3] For example, existing R software for evaluating the distribution function of the t copula imposes some limitations on $\nu$. The algorithms in the R package copula (Hofert et al.,, 2023), which are based on methods by Gentz and Bretz (Gentz and Bretz,, 2009), are limited to integer $\nu$ and this is insufficiently fine-grained for applications of $C^{|t|}_{\nu}$. The C++ library underlying rvinecopulib (Nagler and Vatter,, 2022) imposes the restriction that $\nu\geqslant 2$..
5.2 Simulation study of jointly symmetric case
We explore the ideas of the previous section via a Monte Carlo simulation study. We generate data from the classical symmetric ARCH(1) and GARCH(1,1) processes using the parameter values and innovation distributions summarized in Table 1. We then fit copulas with the property of joint symmetry to the generated realizations of the pairs $(X_{t},X_{t+1})^{\top}$.
In this and all subsequent simulations, the analyses are based on simulated samples of 50000 points. Copulas are estimated by the method of pseudo maximum likelihood (PML) (Genest and Rivest,, 1993); in other words, marginal distributions are estimated empirically, the data are mapped to the unit square using the empirical distribution function and copula parameters are then determined by maximum likelihood. Models are compared according to their AIC values.
The first copula fitted in Table 2 is $C^{t}_{\nu}$, the copula of the spherical t distribution, i.e. the t copula with the correlation parameter fixed at $\rho=0$; the only free parameter is the degree-of-freedom parameter $\nu$. We also label this copula iv-ast, to emphasize that it can be viewed as the inverse v-transformed version of the absolute spherical t copula $C^{|t|}_{\nu}$. The second model is the generalisation of $C^{t}_{\nu}$ consisting of an equal mixture of the t copula $C^{t}_{\nu,\rho}$ and its 90-degree rotation $C^{t}_{\nu,-\rho}$ for $\rho\geqslant 0$.
According to AIC, the $C_{\nu}^{t}$ is the best model in the table in two of six cases, which are the cases where the innovations are Student t4. The mixture of two t copulas is an inferior model to $C_{\nu}^{t}$ in 5 of 6 cases, the only exception being GARCH with normal innovations. The maximum likelihood method for the t mixture returns an estimate $\rho\approx 0$ in those 5 cases, which accounts for the fact that the difference in AIC values is approximately 2.
The next panel of 3 models contains equal mixtures of the Clayton, Gumbel and Joe copulas and their 90, 180 and 270 degree rotations. The Gumbel case is the equally-weighted special case of the convex Gumbel copula of Junker and May, (2005), as also used by Loaiza-Maya et al., (2018). The next 3 copulas in the first panel of Table 2 are inverse-v-transformed versions of the upper tail-dependent Gumbel, Joe and Clayton survival copulas (the latter described as Clayton-180), i.e. copulas with densities of the form $c^{*}(|2u-1|,|2v-1|)$ where $c^{*}$ is the density of either the Clayton survival copula, the Gumbel copula or the Joe copulas.
Comparing all of the 1-parameter models, the $C_{\nu}^{t}$ (iv-ast) copula is always the best model; second place is usually occupied by the iv-Joe copula, except for ARCH with Gaussian innovations where the iv-Clayton180 copula is favoured. Interestingly, while the iv-Gumbel copula is the worst of the inverse-v-transformed copulas, the mixture of Gumbels is often the best of the mixtures.
The remaining models in Table 2 are non-exchangeable. The fourth panel of the table contains mixture copulas, except that the Khoudraji extension in (21) is applied to each copula before the mixture of the 4 different rotations is constructed. Thus these models introduce two further parameters $a_{1}$ and $a_{2}$. The final panel contains inverse-v-transformed versions of the Khoudraji extensions of the Clayton survival, Gumbel and Joe copulas. It would be appealing to add the Khoudraji construction based on the iv-ast copula, but this model is computationally more demanding and not currently a viable candidate model for practical applications.
The best of the 3-parameter non-exchangeable models are always in the inverse-v-transformed class. The mixture copulas are only really competitive for t2.5 distributed innovations. The best-fitting copulas for Gaussian innovations (iv-K-Clayton-190 for ARCH and iv-K-Joe for GARCH) give estimates of $a_{2}$ which are very close to zero, while the best-fitting copula for t2.5 innovations (iv-K-Joe) give estimates of $a_{1}$ which are very close to zero. This makes sense in relation to Figure 2 since the non-exchangeability in the copulas seems to be caused by a “twist” in different directions.
It is notable that the best model of all is iv-ast in the two cases where the innovations are t4. In the case of ARCH, this is in accord with the near-exchangeability evident in Figures 1 and 2. The estimated degree of freedom parameters are 2.56 for ARCH and 2.67 for GARCH.
Model ARCH(1) GARCH(1,1)
Copula $p$ | $(\epsilon_{t})$ Gauss t4 t2.5 Gauss t4 t2.5
$C^{t}_{\nu}$ (iv-ast) 1 482 0 115 56 0 77
mix-t 2 484 2 117 54 2 79
mix-Clayton 1 1630 426 178 444 284 127
mix-Gumbel 1 867 150 171 181 113 138
mix-Joe 1 1936 593 265 619 424 214
iv-Clayton-180 1 483 146 281 127 98 198
iv-Gumbel 1 1121 444 368 375 300 272
iv-Joe 1 604 50 134 97 58 113
mix-K-Clayton 3 1124 430 91 338 281 52
mix-K-Gumbel 3 728 154 55 152 117 46
mix-K-Joe 3 1634 597 167 585 428 133
iv-K-Clayton-180 3 0 147 145 42 91 101
iv-K-Gumbel 3 552 448 210 283 296 141
iv-K-Joe 3 47 54 0 0 56 0
Table 2: Comparison of copula fits when jointly symmetric copulas are fitted to simulated data from symmetric ARCH(1) and GARCH(1,1) processes by the PML method. Each column shows the differences between AIC and the minimum value of AIC for that column; thus a 0 indicates the model with lowest AIC.
5.3 Simulation study of h-symmetric case with leverage
Model ARCH(1) GARCH(1,1)
Copula $p$ | $(\epsilon_{t})$ Gauss t4 t2.5 Gauss t4 t2.5
iv-ast 1 980 224 222 683 262 173
iv-Clayton-180 1 1009 358 363 806 384 299
iv-Joe 1 1114 267 237 727 319 200
iva-ast 2 258 0 131 23 0 68
iva-Clayton-180 2 364 152 280 209 139 202
iva-Joe 2 374 48 152 54 60 100
iv-K-Clayton-180 3 655 361 239 715 386 238
iv-K-Joe 3 672 271 113 619 323 132
iva-K-Clayton-180 4 22 153 131 151 136 112
iva-K-Joe 4 0 52 0 0 61 0
Table 3: Comparison of AIC values when various h-symmetric copulas are fitted to simulated data from ARCH(1) and GARCH(1,1) processes with symmetric innovations and leverage by the PML method. Each column shows the differences between AIC and the minimum value of AIC for that column; thus a 0 indicates the model with lowest AIC.
In this section we consider the ARCH(1) and GARCH(1,1) models with symmetric innovations and leverage. According to Proposition 4, the first-order copula $C_{1}$ is h-symmetric but not jointly symmetric. According to Example 2, the ARCH(1) process has a copula density of the form $c(\mathcal{V}_{1}(u),|2v-1|)$ for an asymmetric v-transform $\mathcal{V}_{1}$ with fulcrum at $\delta=0.5$. For the GARCH(1,1) model we don’t have the same theoretical result but we will try using copulas of the same form. The parameterisation of the ARCH(1) model with leverage is the same as in Figure 3 while the parameters for the GARCH(1,1) are given in Table 1.
In Table 3 we include a selection of the copulas from Table 2, omitting the mixture copulas, since the inverse-v-transformed copulas typically do better, and also omitting the iv-Gumbel copula, since iv-Clayton-180 and iv-Joe generally do better. We add copulas with the prefix ‘iva’ (referring to asymmetric inverse v-transform) which are models of the form $c(u,v)=c^{*}(\mathcal{V}_{1}(u),|2v-1|)$ where $\mathcal{V}_{1}$ is a member of the family $\mathcal{V}(u;\delta,\kappa)$ in (13) with $\delta=0.5$, so that only one extra unknown parameter $\kappa$ is introduced.
The first panel contains jointly symmetric and exchangeable copulas; the second panel contains h-symmetric copulas based on an exchangeable copula for $c^{*}$ but incorporating an asymmetric v-transform for the first variable; the third panel contains jointly symmetric non-exchangeable copulas using the Khoudraji construction; the final panel contains h-symmetric copulas based on a non-exchangeable $c^{*}$ and incorporating an asymmetric v-transform for the first variable.
For the two models with t4 innovations the best fitting copula is the absolute spherical t copula incorporating an asymmetric v-transform. For the other models the best fitting copula is a Khoudraji-transformed version of the Joe copula incorporating an asymmetric v-transform. Results are very similar for the ARCH(1) and GARCH(1,1) process.
5.4 Simulation study of case of asymmetric innovations
Model ARCH(1) GARCH(1,1)
Copula $p$ leverage leverage
iv-ast 1 168 679 49 436
iv-Clayton-180 1 376 881 208 673
iv-Joe 1 220 728 92 487
ivl-ivl-ast 3 116 201 23 92
ivl-ivl-Clayton-180 3 347 558 189 403
ivl-ivl-Joe 3 164 231 64 140
ivl-ivl-K-Clayton-180 5 336 561 191 403
ivl-ivl-K-Joe 5 168 235 68 144
iv2-iv2-ast 5 0 0 0 0
iv2-iv2-Clayton-180 5 228 293 159 277
iv2-iv2-Joe 5 47 24 45 50
iv2-iv2-K-Clayton-180 7 221 297 160 279
iv2-iv2-K-Joe 7 50 28 49 54
Table 4: Comparison of AIC values when various copulas are fitted to simulated data from ARCH(1) and GARCH(1,1) processes with asymmetric innovations by the PML method. Each column shows the differences between AIC and the minimum value of AIC for that column; thus a 0 indicates the model with lowest AIC.
In the final experiment we consider models with skewed t innovations with parameters $\nu=4$ and $\lambda=0.8$. We simulate ARCH(1) and GARCH(1,1) processes, both with and without leverage effects; parameter values are as in Table 1. We have already observed that asymmetric innovations remove all of the previous symmetries we have discussed and so our approach in this section is pure experimentation without the theoretical insights we exploited in the previous experiments.
The first set of 3 models in Table 4 consists of jointly symmetric and exchangeable copulas, included as a benchmark. The iv-ast copula is the best of these but substantial improvements can be made by introducing asymmetry.
The second set of 3 models is based on copula densities of the form $c(u,v)=c^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))$ where $\mathcal{V}(\cdot;\delta)$ is the linear v-transform with fulcrum at $\delta$ (asymmetric when $\delta\neq 0.5$) and $c^{*}$ is either the ast, Clayton survival or Joe copula [^4] It should be noted that full maximum likelihood estimation with respect to all parameters including the parameters of the v-transform presents no problems for these models since the ast, Clayton survival and Joe copula densities do not have asymptotes at $(0,0)$. If this were not the case we can run into a threshold estimation issue, as discussed in Bladt and McNeil, 2022b .. These obviously lower the AIC values considerably at the expense of 2 extra parameters. The third set of 2 models applies the Khoudraji transformation to $c^{*}$ in the case of Clayton and Joe, but this only offers an improvement in one case (Clayton-180 and ARCH(1) process).
The fourth set of models is based on copula densities $c(u,v)=c^{*}(\mathcal{V}(u;\delta_{1},\kappa_{1}),\mathcal{V}(v;\delta_{2},\kappa_{2}))$ where $\mathcal{V}(\cdot;\delta,\kappa)$ is the 2-parameter v-transform in (13); such v-transforms are both non-linear and asymmetric. These models give the best results of all, with the model where $c^{*}$ is the ast copula taking first place for all 4 datasets, followed by the model based on Joe. Addition of the Khoudraji transformation in the final set of models again offers no discernible improvement to the models based on Joe. It seems that allowing asymmetric v-transform largely removes the need to consider the Khoudraji transformation, even when the v-transforms are linear.
6 Mimicking GARCH(1,1) processes with simplified d-vines
6.1 Practical issues in the construction of simplified d-vines
In the final sections of this paper we provide some guidance to the construction of simplified d-vine copula models to mimic the behaviour of GARCH-type processes and offer potential alternative models for financial time series. The previous sections have given some insight into the kinds of pair copulas we should consider. We now turn to some of the issues that arise when we want to embed these in simplified vine structures.
Equation (1) in the Introduction implies that the density $c_{(n)}$ of the copula of the variables $(X_{1},\ldots,X_{n})^{\top}$ in a strictly stationary time series model, such as GARCH, can be written in the form
$c_{(n)}(u_{1},\ldots,u_{n})=\\ \prod_{k=1}^{n-1}\prod_{j=k+1}^{n}c_{k}(B_{k-1}(u_{j-k};\bm{u}_{[j-k+1:j-1]}),R_{k-1}(u_{j};\bm{u}_{[j-1:j-k+1]})\mid F_{X}^{-1}(u_{j-k+1}),\ldots,F_{X}^{-1}(u_{j-1}))$ (22)
where, for $k\in\mathbb{N}$, we define the functions $R_{k}:(0,1)\times(0,1)^{k}\to(0,1)$ and $B_{k}:(0,1)\times(0,1)^{k}$
$R_{k}(x;\bm{u})$ $=$ $F_{X_{k+1}|\bm{X}_{[1:k]}}(F_{X}^{-1}(x)\mid F_{X}^{-1}(u_{k}),\ldots,F_{X}^{-1}(u_{1}))$
$B_{k}(x;\bm{u})$ $=$ $F_{X_{1}|\bm{X}_{[2:k+1]}}(F_{X}^{-1}(x)\mid F_{X}^{-1}(u_{1}),\ldots,F_{X}^{-1}(u_{k}))$
which we refer to as forward and backward Rosenblatt functions, respectively. The notation is based on the one used in Bladt and McNeil, 2022a with minor changes. Note that, in (22), we adopt the convention that $R_{0}(x;\cdot)=B_{0}(x;\cdot)=x$ for all $x\in(0,1)$.
In practice we want to mimic the copula (22) with a simplified d-vine, i.e. a model where the copula has joint density of the simpler form
$c_{(n)}(\bm{u})=\prod_{k=1}^{n-1}\prod_{j=k+1}^{n}c_{k}\Big{(}B_{k-1}(u_{j-k};\bm{u}_{[j-k+1:j-1]}),R_{k-1}(u_{j};\bm{u}_{[j-1:j-k+1]})\Big{)}.$ (23)
In this case the Rosenblatt functions are determined in a recursive, interlacing fashion by $R_{1}(x;u)=h_{1,1}(u,x)$, $B_{1}(x;v)=h_{1,2}(x,v)$ and
[…]
where $h_{k,1}(u,v)=\frac{\partial}{\partial u}C_{k}(u,v)$ and $h_{k,2}(u,v)=\frac{\partial}{\partial v}C_{k}(u,v)$ are the so-called $h$-functions of the copula $C_{k}$ with density $c_{k}$.
It seems unlikely that we can perfectly mimic a GARCH-type process with a simplified d-vine, since the true copula (22) cannot be shown to conform to the simplifying assumption. However, the weak dependence of $c_{2}(u,v;w)$ on the conditioning variable investigated in Section 3.5 gives some optimism that we could at least approximate a GARCH-type model with a simplified d-vine.
Other than ARCH(1)-type processes, which are Markov processes of order 1, GARCH(1,1)-type processes have infinite memory in the sense that $X_{t}$ depends on all previous values $X_{t-k}$, for $k\in\mathbb{N}$, albeit the serial dependence on past observations dies away relatively quickly; for example, the autocorrelation function of the squared values of the classical GARCH(1,1) process decays like an ARMA(1,1) process. To mimic a GARCH(1,1)-type process, it seems natural to consider d-vine processes with infinite-order partial copula dependence as in Bladt and McNeil, 2022a . These are models which are built from a sequence of pair copulas $(C_{k})_{k\in\mathbb{N}}$ which is parsimoniously parameterized using the Kendall partial autocorrelation function (kpacf) or, in other words, the sequence of Kendall rank correlations of $(C_{k})_{k\in\mathbb{N}}$.
GARCH(1,1) models are very parsimoniously parameterized. For example, a model that incorporates both leverage and skewed Student t innovations would have only 6 parameters ($\alpha_{0}$, $\alpha_{1}$, $\beta_{1}$, $\gamma_{1}$, $\nu$, $\lambda$). In our d-vine analogues we also need to control the number of parameters and the method of parameterization via the kpacf developed in Bladt and McNeil, 2022a can be adapted to our context by applying it to an inverse-v-transformed version of the copula series $(C^{*}_{k})_{k\in\mathbb{N}}$. To be more precise, each pair copula $C_{k}$ used in the sequence defining the d-vine would be constructed to have density $c_{k}(u,v)=c_{k}^{*}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v))$ where the v-transforms $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ would be two different, probably asymmetric functions with a fixed form for all $k$; in essence they would take care of the effects caused by leverage and asymmetric innovations in a similar manner to $\gamma_{1}$ and $\lambda$ in the aforementioned GARCH(1,1) model. The copula sequence $(C^{*}_{k})_{k\in\mathbb{N}}$ would contain Joe, Clayton survival or ast copulas with Kendall correlations $\tau_{k}$ that coincided with the terms of the kpacf of a classical Gaussian ARMA(1,1) process, or some other low-order linear process. The combination of ARMA(1,1) and two linear v-transforms would give precisely 4 parameters to describe the d-vine copula process, although two or more parameters would also be required to model the marginal distribution of $(X_{t})_{t\in\mathbb{Z}}$.
We have listed the ast copula $C^{|t|}_{\nu}$ as a candidate pair copula for the sequence $(C^{*}_{k})_{k\in\mathbb{N}}$, but it is not clear whether there is a simple parametric formula to link Kendall’s tau to the degree of freedom parameter $\nu$. Simple calculations from (20) show that
$\tau(\nu)=64\int_{0}^{0.5}\int_{0}^{0.5}C_{\nu}^{t}(u,v)c_{\nu}^{t}(u,v)\mathrm{d}u,\mathrm{d}v-1$ (25)
where $C_{\nu}^{t}$ and $c_{\nu}^{t}$ are the spherical t copula and its density, and that $\tau(\nu)$ ranges in the interval $(0,1)$ as $\nu$ ranges in the interval $(0,\infty)$. An analytical expression for $\tau(\nu)$ would make the ast copula a very attractive choice for these models but, in the absence of such an expression, the Joe copula may be the easier option.
Finally we consider the choice of v-transforms for $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$. To estimate a d-vine model by maximum likelihood we have to be able to evaluate a likelihood based on (23) which means evaluating the Rosenblatt forward and backward functions using the recursive expressions (24). This in turn entails evaluating multiple $h$-functions. Moreover, to generate realizations from a copula with joint density (23), we have to be able to invert the Rosenblatt forward functions which entails the evaluation of inverse $h$-functions. To this end it is convenient to have closed form expressions for the copulas $C_{k}$ and their $h$-functions and inverse $h$-functions. For copula densities $c_{k}$ obtained from the inverse-v-transformation construction in (15) this is unfortunately not straightforward, except in the case where we have copula densities of the form $c(u,v)=c^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))$ constructed from two linear v-transforms $\mathcal{V}_{\delta_{1}}$ and $\mathcal{V}_{\delta_{2}}$ and a copula density $c^{*}$ which itself has a tractable joint distribution function and $h$-functions. We investigate this in the next section.
6.2 Tractable processes using linear v-transforms
The following results shows that all the formulas required to work with linear v-transforms are easily obtained.
Proposition 7.
The copula $C$ with density given by $c(u,v)=c^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))$ is
$C(u,v)=\delta_{1}^{I_{\{u\leqslant\delta_{1}\}}}(\delta_{1}-1)^{I_{\{u>\delta_{1}\}}}\delta_{2}^{I_{\{v\leqslant\delta_{2}\}}}(\delta_{2}-1)^{I_{\{v>\delta_{2}\}}}C^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))+\delta_{1}v+\delta_{2}u-\delta_{1}\delta_{2}$ (26)
where $C^{*}$ is the copula with density $c^{*}$. Writing $h_{1}^{*}$ and $h_{2}^{*}$ for the partial derivatives of $C^{*}$, the corresponding partial derivatives $h_{1}$ and $h_{2}$ of $C$ are
$h_{1}(u,v)$ $=-\delta_{2}^{I_{\{v\leqslant\delta_{2}\}}}(\delta_{2}-1)^{I_{\{v>\delta_{2}\}}}h_{1}^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))+\delta_{2}$
$h_{2}(u,v)$ $=-\delta_{1}^{I_{\{u\leqslant\delta_{1}\}}}(\delta_{1}-1)^{I_{\{u>\delta_{1}\}}}h_{2}^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))+\delta_{1}$
and satisfy the equations
$\mathcal{V}\left(h_{1}(u,v);\delta_{2}\right)$ $=h_{1}^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))$ (27)
$\mathcal{V}\left(h_{2}(u,v);\delta_{1}\right)$ $=h_{2}^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2})).$
Moreover the inverses $h_{1}^{-1}(u,v)=\{x:h_{1}(u,x)=v\}$ and $h_{2}^{-1}(u,v)=\{y:h_{2}(y,v)=u\}$ are
$h_{1}^{-1}(u,v)$ $=-\delta_{2}^{I_{\{v\leqslant\delta_{2}\}}}(\delta_{2}-1)^{I_{\{v>\delta_{2}\}}}h_{1}^{*-1}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))+\delta_{2}$
$h_{2}^{-1}(u,v)$ $=-\delta_{1}^{I_{\{u\leqslant\delta_{1}\}}}(\delta_{1}-1)^{I_{\{u>\delta_{1}\}}}h_{2}^{*-1}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))+\delta_{1}$
where $h_{1}^{*-1}$ and $h_{2}^{*-1}$ are the inverses of $h_{1}^{*}$ and $h_{2}^{*}$.
Proof.
First note that $\frac{\partial^{2}}{\partial u\partial v}C(u,v)=c^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))$. We just need to check that the expression (26) defines a copula, which entails checking that $C(0,v)=0$ and $C(1,v)=v$ for any $v$; by symmetry it would follow that $C(u,0)=0$ and $C(u,1)=u$ for any $u$. Now
$C(0,v)=\delta_{1}\delta_{2}^{I_{\{v\leqslant\delta_{2}\}}}(\delta_{2}-1)^{I_{\{v>\delta_{2}\}}}\mathcal{V}(v;\delta_{2})+\delta_{1}v-\delta_{1}\delta_{2}$
and this yields 0 for both $v\leqslant\delta_{2}$ and $v>\delta_{2}$; moreover
$C(1,v)=(\delta_{1}-1)\delta_{2}^{I_{\{v\leqslant\delta_{2}\}}}(\delta_{2}-1)^{I_{\{v>\delta_{2}\}}}\mathcal{V}(v;\delta_{2})+\delta_{1}v+\delta_{2}(1-\delta_{1})$
and this yields $v$ for both $v\leqslant\delta_{2}$ and $v>\delta_{2}$. The $h$-functions are obtained by differentiation of (26) and the identities (27) are easy to verify. Noting that $h_{1}(u,\delta_{2})=\delta_{2}$ and $h_{2}(u,\delta_{1})=\delta_{1}$, it is also straightforward to derive the inverses of the $h$-functions. ∎
We end by considering the special case of the model with linear v-transforms in which the two v-transforms are identical. We find that the resulting d-vine copula also belongs to the class of vt-d-vines in Bladt and McNeil, 2022b ; these models have been used with real financial data and found to be tractable and effective. The statement of the result and proof use the notation $\mathcal{V}(\bm{u};\delta)$ for the element-by-element v-transformation of a vector $\bm{u}$ with the same linear v-transform.
Theorem 3.
Suppose we construct a d-vine copula (23) using a sequence of copulas $(C_{k})_{k\in\mathbb{N}}$ whose densities are all of the form $c_{k}(u,v)=c_{k}^{*}(\mathcal{V}(u;\delta),\mathcal{V}(v;\delta))$ for a sequence of copulas $(C_{k}^{*})_{k\in\mathbb{N}}$ with densities $(c_{k}^{*})_{k\in\mathbb{N}}$ and a single linear v-transform $\mathcal{V}(\cdot,\delta)$. Then
$c_{(n)}(\bm{u})=\prod_{k=1}^{n-1}\prod_{j=k+1}^{n}c_{k}^{*}\Big{(}B^{*}_{k-1}(\mathcal{V}(u_{j-k};\delta);\mathcal{V}(\bm{u}_{[j-k+1:j-1]};\delta)),R^{*}_{k-1}(\mathcal{V}(u_{j};\delta);\mathcal{V}(\bm{u}_{[j-1:j-k+1]};\delta))\Big{)}$ (28)
where $(R_{k}^{*})_{k\in\mathbb{N}}$ and $(B_{k}^{*})_{k\in\mathbb{N}}$ are the forward and backward Rosenblatt functions corresponding to the copula sequence $(C_{k}^{*})_{k\in\mathbb{N}}$.
Proof.
The result will follow if we can show that the equalities
$\mathcal{V}\left(B_{k-1}(u;\bm{v});\delta\right)=B^{*}_{k-1}(\mathcal{V}(u;\delta);\mathcal{V}(\bm{v};\delta))\quad\text{and}\quad\mathcal{V}\left(R_{k-1}(u;\bm{v});\delta\right)=R^{*}_{k-1}(\mathcal{V}(u;\delta);\mathcal{V}(\bm{v};\delta))$ (29)
hold for any $k\in\mathbb{N}$, $u\in(0,1)$ and $\bm{v}\in(0,1)^{k-1}$. We will argue by induction. The equalities (29) are obviously true for $k=1$ since, by our convention, $B_{0}(u;\cdot)=B_{0}^{*}(u;\cdot)=u$ and $R_{0}(u;\cdot)=R_{0}^{*}(u;\cdot)=u$ for all $u$. Assume the equalities (29) are true when $k=l\in\mathbb{N}$ and consider $k=l+1$. For the first equality we can show, by first using formula (27) in Proposition 7 and then applying the inductive assumption, that
$\mathcal{V}\left(B_{l}(u;\bm{v});\delta\right)$ $=\mathcal{V}\left(h_{l,2}\left(B_{l-1}(u;\bm{v}_{[1,k-1]}),R_{l-1}(v_{k};\bm{v}_{[k-1:1]})\right);\delta\right)$
$=h_{l,2}^{*}\left(\mathcal{V}\left(B_{l-1}(u;\bm{v}_{[1,k-1]});\delta\right),\mathcal{V}\left(R_{l-1}(v_{k};\bm{v}_{[k-1:1]});\delta\right)\right)$
$=h_{l,2}^{*}\left(B^{*}_{l-1}(\mathcal{V}(u;\delta);\mathcal{V}(\bm{v}_{[1,k-1]};\delta)),R^{*}_{l-1}(\mathcal{V}(v_{k};\delta);\mathcal{V}(\bm{v}_{[k-1:1]};\delta))\right)=B^{*}_{l}(\mathcal{V}(u;\delta);\mathcal{V}(\bm{v};\delta))$
and the same argument can be applied to the second equality. ∎
Equation (28) describes a vt-d-vine copula process. In such a model the copula density can be evaluated by first applying the v-transform to the entire vector $(u_{1},\ldots,u_{n})$ and then passing the resulting values to the density of a d-vine copula based on the sequence $(C_{k}^{*})_{k\in\mathbb{N}}$. McNeil, (2021) shows that if we apply such a copula process to a series $(X_{t})_{t\in\mathbb{Z}}$ we effectively use the copula sequence $(C_{k}^{*})_{k\in\mathbb{N}}$ to model the serial dependence structure of an element-wise transformation of the time series $(T(X_{t}))_{t\in\mathbb{Z}}$. The case where the copulas are all Gaussian corresponds to the vt-ARMA model in McNeil, (2021), although this tends to be an inferior model to the ones based on non-Gaussian d-vines in Bladt and McNeil, 2022b . The implied canonical form of $T$ for a linear v-transform $\mathcal{V}(\cdot;\delta)$ and a continuous marginal distribution function $F_{X}$ is the v-shaped function on $\mathbb{R}$ given by
$T(x)=\begin{cases}F_{X}^{-1}\left(F_{X}(x)+\mathcal{V}(F_{X}(x);\delta)\right)-\mu&x\leqslant\mu,\\ x-\mu&x>\mu,\end{cases}$ (30)
where $\mu=F_{X}^{-1}(\delta)$ [^5] The function $T$ in (30) is not the unique transformation that gives a time series whose dependence structure may be described by the joint density (28). There is, in fact, an equivalence class of transformations that all lead to the same copula density when paired with $F_{X}$. For example, in the case where $F_{X}$ is symmetric and $\delta=0.5$ we could also take $T(x)=x^{2}$. We pin down the canonical form by fixing the gradient to be 1 for $x>\mu$. See McNeil, (2021) for more details.. For example, in the case where $F_{X}$ is symmetric around the origin and $\delta=0.5$, $\mu=0$ and $T(x)=|x|$. In general $T$ can be thought of as a function that measures the magnitude of a return relative to $\mu$ in a way that can also account for asymmetric effects for returns that are greater or less than $\mu$.
7 Conclusion
This paper has asked two questions. First, what are the bivariate copulas that underlie the serial dependencies and serial conditional dependencies of pairs of variables $(X_{t},X_{t+k})$ in first-order GARCH(1,1) models. Second, can these copulas be mimicked by simpler constructions with a view to developing simplified d-vines that can behave in a similar fashion to GARCH-type models, but offer more flexible solutions for datasets where GARCH works less well.
With regard to the first question we find interesting symmetries in the dependence structures of the models with symmetric innovation distributions, even when leverage effects are introduced, but these symmetries disappear in the presence of asymmetric innovations, which are often necessary in practice to model real-world datasets. What can generally always be said is that the underlying bivariate copulas are cross-shaped, in the sense that there is a tendency for points to accumulate in all fours corners of the unit square, and that this tendency increases at shorter lags as the size of the ARCH effect increases and persists at longer lags as the aggregate size of the ARCH and GARCH effects increases. Heavier-tailed innovation distributions lead to weaker dependence.
However, the true copulas and conditional copulas of GARCH-type processes remain inaccessible for practical use, mainly because the stationary distribution of these models is not available in closed form. This leads us to consider the second question of whether GARCH copulas can be mimicked with simpler copulas. Motivated by the recognition that v-transforms are implicitly embedded in the copulas of symmetric GARCH models, we propose a construction based on v-transforms and positive-dependence copulas with strong dependence in the upper joint tail; these constructions are naturally cross-shaped. Of the well-known copulas, Joe’s copula in combination with asymmetric v-transforms seems the most promising for mimicking GARCH. However, we also identify a new copula - the absolute spherical t copula - which seems to offer an even more promising basis for the construction.
Could these copulas be used to construct simplified d-vines to mimic GARCH processes? To this question we give a more qualified answer. While the Markovian first-order ARCH-type models can definitely be mimicked by a d-vine(1) models, as soon as a GARCH effect is introduced, it is not clear whether the simplifying assumption will restrict our ability to perfectly mimic GARCH behaviour; a first investigation of this question in Section 5 produces some weak evidence that the dependence of the conditional copulas on the conditioning variables may not be particularly strong, but more investigation is required. Moreover, the computational requirements of d-vines, as discussed in Section 6, mean that we may have difficulty in integrating the most promising GARCH-mimicking pair copulas (such as iv2-iv2-ast) into d-vine structures and we may have to compromise and use copulas based on linear v-transforms. Further examination of these issues is required.
Could the GARCH-mimicking copulas be used to construct d-vines to model real financial return data? This is a different question to the question of mimicking GARCH processes. Bladt and McNeil, 2022b  have shown that the class of vt-d-vine processes with linear v-transforms is competitive for modelling many financial return series and, in particular, can give better results for data which depart from the power-lawed marginal behaviour modelled by a GARCH process. The d-vine processes constructed from GARCH-mimicking pair copulas, as described in this paper, can offer additional flexibility because of the use of two v-transforms in their construction, even if we do restrict both to be linear. There is a need for further empirical research to evaluate the potential of such models.
Another issue requiring further research is the construction of d-vines to mimic ARMA-GARCH processes. Real-world financial return series sometimes show serial correlation in the raw data as well as in the squared or absolute value series and this is usually modelled by combining ARMA and GARCH effects. In this paper we have concentrated on pure GARCH-type models and given only a single example in Section 3.4 of an AR-ARCH model. The illustration in the lower-right panel of Figure 3 suggests mixtures of positive dependence copulas and GARCH-mimicking copulas could be successful in addressing combined ARMA and GARCH effects. As a pragmatic alternative, the d-vine models discussed in this paper could be applied to the residuals from an ARMA model instead.

### v2

@@
-GARCH copulas and GARCH-mimicking copulas
+GARCH copulas, v-transforms and D-vines for stochastic volatility
 Abstract
@@
 It is shown that the copulas of symmetric GARCH processes are jointly symmetric but non-exchangeable, while the copulas of processes with symmetric innovation distributions and asymmetric leverage effects have weaker h-symmetry; copulas with asymmetric innovation distributions have neither form of symmetry.
-Since the actual copulas are typically inaccessible, due to the unknown functional forms of the marginal distributions of GARCH processes, methods of mimicking them are proposed.
-These rely on constructions that combine standard bivariate copulas for positive dependence with two uniformity-preserving transformations known as v-transforms.
-A variety of new copulas are introduced and the ones providing the best fit to simulated data from GARCH-type processes are identified.
-A method of constructing tractable simplified d-vines using linear v-transforms is described and shown to coincide with the vt-d-vine model when the two v-transforms are identical.
+Since the true bivariate copulas are typically inaccessible, due to the unknown functional forms of the marginal distributions of GARCH processes, a new class of approximating copulas is proposed.
+These rely on copula density constructions that combine standard bivariate copula densities for positive dependence with two uniformity-preserving transformations known as v-transforms.
+The construction is shown to be particularly effective when applied to the density of the copula of the absolute values of a spherical t distribution.
+Tractable simplified D-vines incorporating the new pair copulas are developed for applications to time series showing stochastic volatility.
+The resulting models are shown to provide better fits to simulated data from GARCH processes, and to a dataset of financial exchange-rate returns, than have previously been obtained using vine copulas.
 Keywords: Time series; copulas; vine copulas; GARCH processes; v-transforms.
 1 Introduction
-The ARCH process (Engle,, 1982), the GARCH process (Bollerslev,, 1986) and a whole lexicon of related GARCH-type models have proved to be remarkably successful models for time series displaying stochastic volatility, such as asset price return series.
-In this paper we analyse the serial dependence structure of first-order GARCH-type models in terms of the implied bivariate copulas that describe the dependence of pairs of variables at different lags.
-Our aim is to understand how we could mimic such dependence structures with appropriately chosen bivariate copulas arranged in d-vines; we do this with a view to developing models that might offer additional flexibility in practice.
-The approach we take recognizes that uniformity-preserving transformations, known as v-transforms (McNeil,, 2021), are embedded in the implied copulas of classical GARCH models and suggests that they could be used as explicit modelling elements to construct vines for applications.
-The essential mechanism of a GARCH-type model is simple and intuitive; the conditional distribution from which a new observation is drawn has a standard deviation that is a function of the size of previous returns, so that large market moves have a tendency to beget further large market moves and create periods of volatility.
-The ARCH and GARCH prototypes of Engle, (1982) and Bollerslev, (1986) have inspired a vast class of related GARCH-type models, including the GJR-GARCH model, which incorporates leverage effects to reflect different contributions to volatility by returns of differing signs (Glosten et al.,, 1993), power-GARCH models (Ding et al.,, 1993) and exponential GARCH models (Nelson,, 1991), to name a but a few of the more important extensions.
-A number of comprehensive surveys of the extended GARCH family have been published including Bollerslev et al., (1992) Bollerslev et al., (1994) and Shephard, (1996).
-Notwithstanding the appealing basic mechanism for explaining volatility, GARCH-type models are also something of a black box.
-The stationary distribution is not generally known in closed form, although it is known that many models gives rise to distributions with Pareto tails (Mikosch and Stărică,, 2000).
+This paper addresses the construction of copula-based models for time series displaying stochastic volatility, such as financial asset-price return series.
+The philosophy of the copula-modelling approach is that, by separating the issues of modelling marginal behaviour and dependence structure, we may obtain composite models that provide superior fit and predictive performance in practice.
+A number of authors (Loaiza-Maya et al.,, 2018; Bladt and McNeil, 2022b, ; Zhao et al.,, 2022) have suggested that simplified D-vines, a class of models based on pair or bivariate copulas, offer a promising framework for modelling volatile return series.
+Finding suitable pair copulas to describe volatile return series is not straightforward, since most existing bivariate copulas are for modelling monotonic dependence relationships.
+The serial dependencies in return series tend to be monotonic in absolute values, but non-monotonic in the raw returns.
+In this paper we attempt to learn from the success of GARCH-type models in practical applications by studying their implied pair copulas.
+Primarily, this is a paper about the dependence structure of GARCH processes rather than empirical data, although we provide an example to show that the insights we gain can lead to highly effective D-vine models for real data.
+The ARCH (Engle,, 1982) and GARCH (Bollerslev,, 1986) prototypes have inspired a vast class of related GARCH-type models, including the GJR-GARCH model, which incorporates leverage effects to reflect different contributions to volatility by returns of differing signs (Glosten et al.,, 1993), power-GARCH models (Ding et al.,, 1993) and exponential GARCH models (Nelson,, 1991), to name but a few of the more important extensions.
+Despite their success in modelling financial data, GARCH-type models are something of a black box.
+The stationary distribution is not generally known in closed form, although many models give rise to distributions with Pareto tails (Mikosch and Stărică,, 2000).
 In particular, the serial dependence structure is not well understood.
-If $(X_{t})_{t\in\mathbb{Z}}$ follows a stationary GARCH-type model then the distribution of $(X_{1},\ldots,X_{n})^{\top}$ for $n\geqslant 2$ must have a copula, but this copula also has no simple, accessible closed form.
-It may be asked why we would want to mimic GARCH-type models with vine copulas, given that GARCH models have proved themselves in practice over four decades.
-While many asset return series are well modelled by GARCH, we do encounter series where a power-lawed marginal distribution is sub-optimal for the data; see, for example, Bladt and McNeil, 2022b .
-For such datasets it would be convenient to be able to construct models that combined bespoke choices of marginal distribution, or non-parametric margins, with a copula that could mimic the kind of serial dependence that is implicit in the GARCH family.
-If we understand the types of copula that are implicit in the GARCH family, it may also be possible to improve upon these copulas by adding further flexibility.
-Vine copulas also have a large literature, including Joe, (1996, 1997), Bedford and Cooke, (2001, 2002), Kurowicka and Cooke, (2006), Aas et al., (2009) and Smith et al., (2010).
+If $(X_{t})_{t\in\mathbb{Z}}$ follows a stationary GARCH-type model then the distribution of $(X_{1},\ldots,X_{n})^{\top}$ for $n\geqslant 2$ must have a copula, but this copula has no simple, accessible closed form.
+Vine copulas also have a large literature, including Joe, (1996), Bedford and Cooke, (2002), Kurowicka and Cooke, (2006), Aas et al., (2009) and Smith et al., (2010).
 The most suitable vine structure for a time series is the D-vine, which is able to describe strict stationarity of a random vector under additional translation-invariance restrictions on the vine structure (Nagler et al.,, 2022).
-Markov copula models (Darsow et al.,, 1992; Chen and Fan,, 2006; Beare,, 2010; Domma et al.,, 2009) are simple examples of stationary first-order d-vines and a number of authors have written on their higher-order extensions (Ibragimov,, 2009; Beare and Seo,, 2015; Brechmann and Czado,, 2015; Loaiza-Maya et al.,, 2018; Nagler et al.,, 2022).
-The last-named paper introduces the terminology s-vine for stationary multivariate vines, which includes the stationary univariate d-vine as a special case.
-The joint density of the distribution of $(X_{1},\ldots,X_{n})^{\top}$ in a strictly stationary process, such as a stationary model of GARCH type, can be decomposed as a d-vine.
-In other words we may write
-$f_{X_{1},\ldots,X_{n}}(x_{1},\ldots,x_{n})=\prod_{i=1}^{n}f_{X}(x_{i})\prod_{k=1}^{n-1}\prod_{j=k+1}^{n}c_{k}(F_{X_{1}|X_{2},\ldots,X_{k}}(x_{j-k}\mid x_{j-k+1},\ldots,x_{j-1}),\\ F_{X_{k+1}|X_{2},\ldots,X_{k}}(x_{j}\mid x_{j-k+1},\ldots,x_{j-1})\mid x_{j-k+1},\ldots,x_{j-1})$ (1)
-where $f_{X}$ denotes the marginal density, $F_{X_{i}|X_{2},\ldots,X_{k}}$ denotes the conditional distribution function of $X_{i}$ given $X_{2},\ldots,X_{k}$ and $c_{k}(u,v\mid x_{2},\ldots,x_{k})$ is the density of the copula of the conditional distribution of $(X_{1},X_{k+1})^{\top}$ given $X_{2}=x_{2},\ldots,X_{k}=x_{k}$.
+Markov copula models (Darsow et al.,, 1992; Chen and Fan,, 2006; Beare,, 2010; Domma et al.,, 2009) are examples of stationary first-order D-vines and a number of authors have written on their higher-order and multivariate extensions (Ibragimov,, 2009; Beare and Seo,, 2015; Brechmann and Czado,, 2015; Loaiza-Maya et al.,, 2018; Nagler et al.,, 2022; Zhao et al.,, 2022).
+Our goals in the present paper are threefold.
+First, we want to analyse the structure of the implied pair copulas in first-order GARCH-type models and to investigate how different model features affect the symmetries of these pair copulas.
+Second, since the true GARCH pair copulas are inaccessible for practical purposes, we want to find tractable copulas that can approximate them.
+Finally, with the help of the analyses in the paper, we aim to propose a class of D-vines that behave like GARCH-type models and give a first glimpse of its suitability for real data applications.
+Our approach is similar to the one followed in the insightful paper of Loaiza-Maya et al., (2018), in which the authors use simulated data to reveal the cross-shaped dependence structure of the copula that describes the serial dependence structure of a classical symmetric ARCH(1) process [^1] Loaiza-Maya et al., (2018) and Smith and Maneesoonthorn, (2018) also analyse the first-order copulas of discrete-time stochastic volatility models, an alternative model class for financial asset-price returns we do not consider; they find similar cross-shaped behaviour..
+They use this analysis to propose D-vines of low Markov order (1 or 5) in which the pair copulas are mixtures of positive dependence and negative dependence copulas; they show that these models give better forecasts of the conditional quantiles (value-at-risk) of exchange-rate returns than classical symmetric models of first-order GARCH type.
+Another paper related to our own is by Zhao et al., (2022), who argue that t copulas in D-vines can approximate the dependence structure of GARCH(1,1) and GJR-GARCH(1,1) models.
+Our paper also extends ideas in Bladt and McNeil, 2022b by proposing a class of models that partly overlaps with the vt-D-vines in that paper but offers additional flexibility.
+To explain our contributions, we note that the joint density of the distribution of $(X_{1},\ldots,X_{n})^{\top}$ in a strictly stationary process $(X_{t})_{t\in\mathbb{Z}}$, such as a stationary model of GARCH type, can be decomposed as a D-vine of the form
+$f_{X_{1},\ldots,X_{n}}(x_{1},\ldots,x_{n})=\prod_{i=1}^{n}f_{X}(x_{i})\\ \prod_{k=1}^{n-1}\prod_{j=k+1}^{n}c_{k}(F_{X_{1}|\bm{X}_{[2:k]}}(x_{j-k}\mid\bm{x}_{[j-k+1:j-1]}),F_{X_{k+1}|\bm{X}_{[2:k]}}(x_{j}\mid\bm{x}_{[j-k+1:j-1]})\mid\bm{x}_{[j-k+1:j-1]})$ (1)
+where we use the compact notation $\bm{x}_{[l:m]}:=(x_{l},\ldots,x_{m})^{\top}$, $f_{X}$ denotes the marginal density, $F_{X_{i}|\bm{X}_{[2:k]}}$ is the conditional distribution function of $X_{i}$ given $X_{2},\ldots,X_{k}$ and $c_{k}(u,v\mid\bm{x}_{[j-k+1:j-1]})$ is the density of the copula of the conditional distribution of $(X_{1},X_{k+1})^{\top}$ given $X_{2}=x_{j-k+1},\ldots,X_{k}=x_{j-1}$.
 In the case $k=1$ the conditional distribution is simply the marginal distribution $F_{X}$ and $c_{1}$ is simply the density of the copula of the unconditional distribution of $(X_{1},X_{2})^{\top}$.
-When the process is first-order Markov, all of the terms involving $c_{k}(u,v\mid\cdot)$ for $k>1$ disappear, but this will only apply to the ARCH(1)-type models among the models we analyse.
-Although the densities of GARCH-type models may be written as d-vines, they may not in general be written as so-called simplified d-vines, that is models in which the conditional copula densities $c_{k}$ do not depend explicitly on the values $x_{j-k+1},\ldots,x_{j-1}$ in equation (1) for $k\geqslant 2$.
-When d-vines are used in practical modelling applications they are almost always simplified d-vines.
-However, as pointed out by a number of authors (Haff et al.,, 2010; Stöber et al.,, 2013; Spanhel and Kurz,, 2019; Mroz et al.,, 2021), these may not be able to capture all the features present in an arbitrary joint density, such as that of a GARCH-type model.
-In the course of our analysis we return to this question.
-We confine our attention to first-order GARCH-type models.
-In practice models of order (1,1) often work very well and adding higher-order terms yields only minor improvements.
-We also look in some detail at ARCH(1)-type models for two reasons: they are first-order Markov models and therefore particularly easy to analyse; they give rise to the essential copula forms that are encountered in the more general class, as will be seen in this paper.
-Loaiza-Maya et al., (2018) have previously investigated the copula of the classical ARCH(1) process.
-Since this model is first-order Markov, its serial dependence may be fully described by a first-order d-vine with a single bivariate copula, namely the copula of the distribution of two successive values $(X_{1},X_{2})^{\top}$.
-Using a simulation approach, they identify the cross-shaped nature of the copula density and suggest that mixtures of positive dependence and negative dependence copulas (such as a t copula and a 90-degree rotated t copula) can model the dependence structure effectively.
-Zhao et al., (2022) argue that t copulas in d-vines can approximate the dependence structure of GARCH(1,1) and GJR-GARCH(1,1) models.
-We will find some evidence in support of t copulas but also argue that they miss some important asymmetries.
-We make a number of original contributions.
-First, we analyse the structure of the bivariate copula densities $c_{k}$ in (1) for classical GARCH-type models; we show that symmetric models give rise to jointly symmetric copulas while models with symmetric innovation distributions and asymmetric volatility have copulas with weaker symmetry, which we call h-symmetry.
-We confirm the theoretical insights by using numerical approximations to provide accurate contour plots of the copula densities.
-These reveal hitherto overlooked features, such as the non-exchangeability of the copulas.
-We also examine the effects of leverage, asymmetric innovations and autoregressive conditional mean terms on the copulas and on their symmetries.
-Second, we show how uniformity-preserving transformations known as v-transforms (McNeil,, 2021) play an implicit role in the copulas of classic symmetric models and in models with certain forms of asymmetry.
-We investigate methods of constructing copulas with v-transforms to mimic the behaviour of the copulas implied by GARCH-type models.
-Using simulated data we identify the copulas based on v-transforms that provide the best fit to simulated data and show that these typically do better than mixture copulas.
-Finally, we draw conclusions with respect to the construction of d-vine copulas to mimic the serial-dependence structure of GARCH-type models.
-We also show how d-vine models with appropriately chosen copulas can be related to the vt-ARMA and vt-d-vine models proposed by McNeil, (2021) and Bladt and McNeil, 2022b .
-2 Preliminary results on symmetry
-2.1 Definitions
-Symmetry plays an important role in this paper and we give a definition to make our use of various symmetry concepts clear; parts (iii) and (iv) are from Nelsen, (2006).
-Definition 1.
-Let $(Y,Z)^{\top}$ be a generic random vector taking values in $\mathbb{R}^{2}$.
-  1. (i)
-    If $(Y,Z-b)^{\top}\stackrel{{\scriptstyle\text{\tiny d}}}{{=}}(Y,b-Z)^{\top}$ for some $b\in\mathbb{R}$ we say that $(Y,Z)^{\top}$ is h-symmetric about $b$ (symmetric about a horizontal axis at $b$).
-  2. (ii)
-    If $(Y-a,Z)^{\top}\stackrel{{\scriptstyle\text{\tiny d}}}{{=}}(a-Y,Z)^{\top}$ for some $a\in\mathbb{R}$ we say that $(Y,Z)^{\top}$ is v-symmetric about $a$ (symmetric about a vertical axis at $a$).
-  3. (iii)
-    If $(Y-a,Z-b)^{\top}\stackrel{{\scriptstyle\text{\tiny d}}}{{=}}(a-Y,b-Z)^{\top}$ for $a,b\in\mathbb{R}$ we say that $(Y,Z)^{\top}$ is radially symmetry about $(a,b)$.
-  4. (iv)
-    If $(Y,Z)^{\top}$ has all three of the above forms of symmetry then we say that it is jointly symmetric about $(a,b)$.
-Remarks 1.
-  1. 1.
-    When $(Y,Z)^{\top}$ is h-symmetric about $b$, the marginal distribution of $Z$ is symmetric about $b$; when $(Y,Z)^{\top}$ is v-symmetric about $a$, the marginal distribution of $Y$ is symmetric about $a$.
-The reverse implications are not true.
-  2. 2.
-    It may be noted that we can always shift the distribution of $(Y,Z)^{\top}$ in two-dimensional space so that axes or centres of symmetry are at zero or the origin.
-Where $a$ and $b$ are not relevant to the result, we drop the explicit reference to them and just refer to h-symmetry, v-symmetry, radial symmetry and joint symmetry.
-  3. 3.
-    H-symmetry and v-symmetry together imply radial symmetry and therefore joint symmetry.
-However radial symmetry does not necessarily imply h-symmetry and v-symmetry, although it does imply symmetry of both marginal distributions.
-The symmetries in Definition 1 have the following implications for the copula of $(Y,Z)^{\top}$.
-Lemma 1.
-Let $(Y,Z)^{\top}$ be a random vector with continuous marginal distributions and copula $C$.
-  1. (i)
-    If $(Y,Z)^{\top}$ is h-symmetric then $C(u,v)=u-C(u,1-v)$, for $(u,v)\in[0,1]^{2}$.
-  2. (ii)
-    If $(Y,Z)^{\top}$ is v-symmetric then $C(u,v)=v-C(1-u,v)$, for $(u,v)\in[0,1]^{2}$,
-  3. (iii)
-    If $(Y,Z)^{\top}$ is radially symmetric then $C(u,v)=u+v+C(1-u,1-v)-1$ , for $(u,v)\in[0,1]^{2}$.
-Proof.
-Since the copula is invariant under strictly increasing transformations of the margins, such as changes of location, we can, without loss of generality, assume that the axes and centres of symmetry are at zero and the origin.
-For case (i) observe that if $(Y,Z)^{\top}\stackrel{{\scriptstyle\text{\tiny d}}}{{=}}(Y,-Z)^{\top}$ then the joint distribution function satisfies
-$F_{Y,Z}(y,z)=\mathbb{P}(Y\leqslant y,Z\leqslant z)=\mathbb{P}(Y\leqslant y,-Z\leqslant z)$ $=\mathbb{P}(Y\leqslant y)-P(Y\leqslant y,Z\leqslant-z)$
-$=F_{Y}(y)-F_{Y,Z}(y,-z)$ (2)
-for all $(y,z)\in\mathbb{R}^{2}$ and application of Sklar’s Theorem (Sklar,, 1959) implies that the (unique) copula $C$ must satisfy
-$C(F_{Y}(y),F_{Z}(z))=F_{Y}(y)-C(F_{Y}(y),F_{Z}(-z))=F_{Y}(y)-C(F_{Y}(y),1-F_{Z}(z))$
-where $F_{Y}$ and $F_{Z}$ denote the marginal distribution functions and the final equality follows from the symmetry of the marginal distribution of $Z$.
-Hence we obtain that $C(u,v)=u-C(u,1-v)$ at all points in the unit square.
-The proof of (ii) is obviously identical to (i) while (iii) is a well-known result found in Nelsen, (2006) (Theorem 2.7.3).
-∎
-Definition 2.
-A copula $C$ satisfying $C(u,v)=u-C(u,1-v)$ for all $(u,v)\in[0,1]^{2}$ will be said to have h-symmetry.
-Similarly, a copula satisfying $C(u,v)=v-C(1-u,v)$ will be said to have v-symmetry.
-A copula satisfying $C(u,v)=u+v+C(1-u,1-v)-1$ will be said to have radial symmetry.
-A copula with all three forms of symmetry will be said to be jointly symmetric.
-We note that joint symmetry of the copula is the bivariate special case of a property that Durante and Fuchs, (2019) call reflection invariance.
-2.2 Correlation and rank correlation
-Nelsen, (2006) notes that jointly symmetric distributions have zero correlation (when finite second moments permit the measure to be defined).
-Perhaps surprisingly, h-symmetry on its own is a sufficient condition for an absence of both correlation and rank correlation.
-The same is true of v-symmetry.
-Proposition 1.
-If a random vector $(Y,Z)^{\top}$ has finite second moments and is either h-symmetric or v-symmetric then the Pearson correlation $\rho(Y,Z)$ is zero.
-Proof.
-We give the proof in the case of h-symmetry, the method for v-symmetry being identical.
-If $(Y,Z)^{\top}$ is h-symmetric about $b$ then $\mathbb{E}(Z)=b$ and we can write $\operatorname{cov}(Y,Z)=\mathbb{E}(YZ)-\mathbb{E}(Y)\mathbb{E}(Z)=\mathbb{E}(Y(Z-b))+\mathbb{E}(Yb)-\mathbb{E}(Y)b=\mathbb{E}(Y(Z-b))$.
-However, h-symmetry implies $\mathbb{E}(Y(Z-b))=\mathbb{E}(Y(b-Z))=-\mathbb{E}(Y,(Z-b))$ implying that $\mathbb{E}(Y(Z-b))=0$ from which the result follows.
-∎
-Proposition 2.
-If a random vector $(Y,Z)^{\top}$ has continuous marginal distributions and is either h-symmetric or v-symmetric, then the Spearman rank correlation $\rho_{S}(Y,Z)$ and the Kendall rank correlation $\tau(Y,Z)$ are both zero.
-Proof.
-Again it suffices to give the proof in the case of h-symmetry.
-Let the unique copula of $(Y,Z)^{\top}$ be denoted $C$.
-For Spearman’s rank correlation we need to calculate $\rho_{S}(Y,Z)=12\int_{0}^{1}\int_{0}^{1}(C(u,v)-uv)\mathrm{d}u\mathrm{d}v$.
-From the identity $C(u,v)=u-C(u,1-v)$ we infer that
-$\int_{0}^{1}\int_{0}^{1}C(u,v)\mathrm{d}u\mathrm{d}v$ $=$ $\int_{0}^{1}\int_{0}^{1}u\mathrm{d}u\mathrm{d}v-\int_{0}^{1}\int_{0}^{1}C(u,1-v)\mathrm{d}u\mathrm{d}v=\frac{1}{2}-\int_{0}^{1}\int_{0}^{1}C(u,y)\mathrm{d}u\mathrm{d}y$
-where we have made the change of variable $y=1-v$. We must have $\int_{0}^{1}\int_{0}^{1}C(u,v)\mathrm{d}u\mathrm{d}v=\tfrac{1}{4}$ and then simple calculations give $\rho_{S}(Y,Z)=0$.
-In the case of Kendall’s rank correlation we need to compute $\tau(Y,Z)=4\mathbb{E}(C(U,V))-1$ where $(U,V)^{\top}$ is a pair of variables with distribution function $C$.
-The identity $C(u,v)=u-C(u,1-v)$ implies that $(U,V)^{\top}\stackrel{{\scriptstyle\text{\tiny d}}}{{=}}(U,1-V)^{\top}$; indeed $(U,V-0.5)^{\top}\stackrel{{\scriptstyle\text{\tiny d}}}{{=}}(U,0.5-V)^{\top}$ since $(U,V)^{\top}$ is h-symmetric about 0.5.
-Hence we infer that
-$\mathbb{E}(C(U,V))$ $=$ $\mathbb{E}(U)-\mathbb{E}(C(U,1-V))=\frac{1}{2}-\mathbb{E}(C(U,V))$
-which implies that $\mathbb{E}(C(U,V))=\tfrac{1}{4}$ and $\tau(Y,Z)=0$. ∎
-Despite the absence of correlation in distributions with these forms of symmetry, we will see in this paper that they can have quite strong forms of dependence.
-2.3 Copula densities
-We will analyse GARCH-type models in terms of their joint densities.
-Henceforth we assume that the generic random vector $(Y,Z)^{\top}$ has joint density $f_{Y,Z}$.
-It is easily inferred from Definition 1, part (i), that h-symmetry is equivalent to $f_{Y,Z}(y,b+z)=f_{Y,Z}(y,b-z)$ for all $(y,z)\in\mathbb{R}^{2}$ and some $b\in\mathbb{R}$.
-Similarly, v-symmetry and radial symmetry are equivalent to the density satisfying $f_{Y,Z}(a+y,z)=f_{Y,Z}(a-y,z)$ for some $a\in\mathbb{R}$ and $f_{Y,Z}(a+y,b+z)=f_{Y,Z}(a-y,b-z)$ for some $a,b\in\mathbb{R}$, respectively.
-The density $c$ of the copula $C$ of $(Y,Z)^{\top}$ is given by
-$c(u,v)=\frac{f_{Y,Z}\left(F_{Y}^{-1}(u),F_{Z}^{-1}(v)\right)}{f_{Y}\left(F_{Y}^{-1}(u)\right)f_{Z}\left(F_{Z}^{-1}(v)\right)}$ (3)
-where $f_{Y}$ and $f_{Z}$ denote the marginal densities.
-The identity $C(u,v)=u-C(u,1-v)$ for h-symmetry of the copula $C$ in Definition 2 is equivalent to the identity $c(u,v)=c(u,1-v)$ for densities (symmetry in the line $v=0.5$).
-Similarly, v-symmetry is equivalent to $c(u,v)=c(1-u,v)$ (symmetry in the line $u=0.5$) and radial symmetry is equivalent to $c(u,v)=c(1-u,1-v)$.
-The following result provides the first connection between copula densities and v-transforms.
-The function $\mathcal{V}(u)=|2u-1|$ appearing in this result is the first and simplest example of a v-transform, as we will discuss in Section 4.
-Proposition 3.
-Let $(Y,Z)^{\top}$ be a random vector with joint density $f_{Y,Z}$ and copula density $c$. Then, for all $(u,v)\in[0,1]^{2}$,
-  1. (i)
-    if $(Y,Z)^{\top}$ is h-symmetric about $b$, $c(u,v)=c^{*}(u,|2v-1|)$ where $c^{*}$ is the density of the copula of $(Y,|Z-b|)^{\top}$;
-  2. (ii)
-    if $(Y,Z)^{\top}$ is v-symmetric about $a$, $c(u,v)=c^{*}(|2u-1|,v)$ where $c^{*}$ is the density of the copula of $(|Y-a|,Z)^{\top}$.
-Proof.
-It suffices to show part (i). We set $\tilde{Z}=Z-b$ and observe that $(Y,\tilde{Z})^{\top}$ is h-symmetric about 0. We then find that
-$F_{Y,|\tilde{Z}|}(y,z)=\mathbb{P}(Y\leqslant y,|\tilde{Z}|\leqslant z)=\mathbb{P}(Y\leqslant y,\tilde{Z}\leqslant z)-\mathbb{P}(Y\leqslant y,\tilde{Z}\leqslant-z)=F_{Y,\tilde{Z}}(y,z)-F_{Y,\tilde{Z}}(y,-z)$
-and hence $f_{Y,|\tilde{Z}|}(y,z)=f_{Y,\tilde{Z}}(y,z)+f_{Y,\tilde{Z}}(y,-z)=2f_{Y,\tilde{Z}}(y,z)$ by h-symmetry about zero. Similarly, $f_{|\tilde{Z}|}(z)=2f_{\tilde{Z}}(z)$. From (3) we have
-$c^{*}(u,|2v-1|)=\frac{f_{Y,|\tilde{Z}|}\left(F_{Y}^{-1}(u),F_{|\tilde{Z}|}^{-1}(|2v-1|)\right)}{f_{Y}\left(F_{Y}^{-1}(u)\right)f_{|\tilde{Z}|}\left(F_{|\tilde{Z}|}^{-1}(|2v-1|)\right)}=\frac{f_{Y,\tilde{Z}}\left(F_{Y}^{-1}(u),F_{|\tilde{Z}|}^{-1}(|2v-1|)\right)}{f_{Y}\left(F_{Y}^{-1}(u)\right)f_{\tilde{Z}}\left(F_{|\tilde{Z}|}^{-1}(|2v-1|)\right)}.$
-Since $F_{|\tilde{Z}|}(z)=2F_{\tilde{Z}}(z)-1$ for $z\geqslant 0$, it follows that $F_{|\tilde{Z}|}^{-1}(q)=F_{\tilde{Z}}^{-1}((1+q)/2)$ for $q\in[0,1]$. Therefore, for $v\leqslant 0.5$, we have
-$F_{|\tilde{Z}|}^{-1}(|2v-1|)=F_{|\tilde{Z}|}^{-1}(1-2v)=F_{\tilde{Z}}^{-1}(1-v)=-F_{\tilde{Z}}^{-1}(v)$
-while, for $v\geqslant 0.5$, we have $F_{|\tilde{Z}|}^{-1}(|2v-1|)=F_{|\tilde{Z}|}^{-1}(2v-1)=F_{\tilde{Z}}^{-1}(v)$.
-Using the h-symmetry of $f_{Y,\tilde{Z}}$ and the symmetry of $f_{\tilde{Z}}$ we infer that $c^{*}(u,|2v-1|)=\tilde{c}(u,v)$ where $\tilde{c}$ is the density of the copula of $(Y,\tilde{Z})^{\top}$.
-The identity $c^{*}(u,|2v-1|)=c(u,v)$ follows from the fact that $(Y,Z)^{\top}$ and $(Y,\tilde{Z})^{\top}$ have the same copula.
-∎
-Clearly, by putting both parts of Proposition 3 together, we see that the density of the copula of a random vector $(Y,Z)^{\top}$ that is jointly symmetric about $(a,b)$ satisfies $c(u,v)=c^{*}(|2u-1|,|2v-1|)$ where $c^{*}$ is the density of the copula of $(|Y-a|,|Z-b|)^{\top}$.
-3 ARCH and GARCH copulas
+When the process is first-order Markov, such as a model of ARCH(1) type, all of the terms involving $c_{k}(u,v\mid\cdot)$ for $k>1$ disappear.
+Although the densities of GARCH-type models may be written as D-vines, they may not in general be written as so-called simplified D-vines, that is models in which the conditional copula densities $c_{k}$ do not depend explicitly on the conditioning values $\bm{x}_{[j-k+1:j-1]}$ in equation (1) for $k\geqslant 2$.
+When D-vines are used in practical modelling applications they are always simplified D-vines.
+However, as pointed out by a number of authors (Haff et al.,, 2010; Stöber et al.,, 2013; Spanhel and Kurz,, 2019; Mroz et al.,, 2021), these may not be able to capture all the features present in a given joint density, such as that of a GARCH-type model.
+Our first contributions relate to the analysis of the GARCH pair copula densities $c_{k}$ in (1) for first-order GARCH-type models and are described in Section 2.
+We provide theory to show that (i) symmetric models give rise to jointly symmetric copulas, (ii) models with symmetric innovation distributions and leverage have copulas with weaker symmetry, which we call h-symmetry, and (iii) models with asymmetric innovations lose all symmetry.
+We confirm the theoretical insights by using numerical approximations to provide accurate contour plots of the copula densities which reveal hitherto overlooked features, such as the non-exchangeability of copulas even in the most symmetric models.
+The theory complements and extends Loaiza-Maya et al., (2018), who consider the copula density $c_{1}$ in the ARCH(1) model and use a simulation approach and adaptive kernel density estimation to elicit its cross-shaped form.
+The second body of material, found in Section 3 relates to the search for copulas that can approximate the GARCH pair copulas.
+We show how uniformity-preserving transformations known as v-transforms (McNeil,, 2021) play an implicit role in the copulas of classic symmetric models and in models with certain forms of asymmetry.
+We investigate methods of constructing copulas to approximate GARCH pair copulas using a technique we call stochastic inversion of v-transforms and we identify the copulas that provide the best fit to simulated data from GARCH-type models.
+We find that inverse-v-transformed versions of positive dependence copulas with upper tail dependence, such as survival Clayton, Joe, or a new copula we refer to as absolute spherical t, typically fit better than mixture copulas.
+Finally, in Section 4, we propose a class of simplified D-vines that behave like GARCH models and that could be used to model real data.
+To support the use of simplified D-vines, we provide some evidence in Section 2.5 that the dependence of higher order conditional copulas on conditioning variables may be weak.
+Our simplified D-vines combine the inverse-v-transformed pair copulas developed in Section 3 with a parsimonious method of parameterization via the partial autocorrelation function of ARMA processes as proposed by Bladt et al., (2025).
+The resulting models generalize the vt-D-vine models of Bladt and McNeil, 2022b and provide a better fit to simulated GARCH data than the simplified D-vines of order 1 and 2 based on t copulas used in Zhao et al., (2022).
+A first application to a real-world dataset of daily exchange-rate returns provided and analysed by Loaiza-Maya et al., (2018) suggests that the much high Markov order of our models, and the copula constructions they contain, can lead to better-fitting models than have previously been obtained using D-vines.
+2 An analysis of GARCH pair copulas
 2.1 ARCH and GARCH processes
@@
   2. 2.
-    A process of ARCH(1) type is a process of GARCH(1,1) type where the volatility function $\sigma(x,s)$ depends only on $x$.
+    A process of ARCH(1) type is a process of GARCH(1,1) type where $\sigma(x,s)$ depends only on $x$.
   3. 3.
-    The classical GARCH(1,1) process is a process of GARCH(1,1) type type in which the volatility function satisfies
+    The classical GARCH(1,1) process is a process of GARCH(1,1) type in which $\sigma(x,s)$ satisfies
     $\sigma(x,s)=\sqrt{\alpha_{0}+\alpha_{1}x^{2}+\beta_{1}s^{2}}$ (3)
@@
   5. 5.
-    A process of GARCH(1,1) type is said to have symmetric distribution, if the distribution of $\epsilon_{1}$ has a density that is symmetric around zero, and to have symmetric volatility if $\sigma(x,w)$ is an even function of $x$.
+    A process of GARCH(1,1) type is said to have symmetric distribution, if the distribution of $\epsilon_{1}$ has a density that is symmetric around zero, and to have symmetric volatility if $\sigma(x,s)$ is an even function of $x$.
 It is said to be symmetric if it has symmetric distribution and symmetric volatility.
@@
 For a finite-variance solution it is usual to assume the stronger condition that $\alpha_{1}+\beta_{1}<1$.
-It is known that the marginal or stationary distribution $F_{X}$ has regularly varying (power-law) tails and Mikosch and Stărică, (2000) showed that the tail index $\zeta$ may be calculated as the root of the equation $\mathbb{E}\left(\left(\alpha_{1}\epsilon_{1}^{2}+\beta_{1}\right)^{\zeta/2}\right)=1$.
-However, the exact functional forms of $F_{X}$ or its density $f_{X}$ are not available.
+It is known that the marginal or stationary distribution $F_{X}$ has regularly varying (power-law) tails with a tail index $\zeta$ that may be calculated as the root of the equation $\mathbb{E}\left(\left(\alpha_{1}\epsilon_{1}^{2}+\beta_{1}\right)^{\zeta/2}\right)=1$ (Mikosch and Stărică,, 2000).
+However, the exact functional forms of $F_{X}$ and its density $f_{X}$ are not available.
+2.2 Symmetry concepts for bivariate distributions
+In our analysis of the pair copulas of GARCH processes, we first consider models with symmetric innovation distributions.
+To describe the resulting copulas, it helps to recall the definitions of radial and joint symmetry in Nelsen, (2006) and to identify two further concepts we call h- and v-symmetry (symmetry about horizontal and vertical axes).
+Definition 2.
+Let $(Y,Z)^{\top}$ be a random vector taking values in a set $A\subseteq\mathbb{R}^{2}$.
+  1. (i)
+    If $(Y,Z-b)^{\top}\stackrel{{\scriptstyle\text{\tiny d}}}{{=}}(Y,b-Z)^{\top}$ for some $b\in\mathbb{R}$ we say that $(Y,Z)^{\top}$ is h-symmetric about $b$.
+  2. (ii)
+    If $(Y-a,Z)^{\top}\stackrel{{\scriptstyle\text{\tiny d}}}{{=}}(a-Y,Z)^{\top}$ for some $a\in\mathbb{R}$ we say that $(Y,Z)^{\top}$ is v-symmetric about $a$.
+  3. (iii)
+    If $(Y-a,Z-b)^{\top}\stackrel{{\scriptstyle\text{\tiny d}}}{{=}}(a-Y,b-Z)^{\top}$ for $a,b\in\mathbb{R}$ we say that $(Y,Z)^{\top}$ is radially symmetric about $(a,b)$.
+  4. (iv)
+    If (i), (ii) and (iii) all hold, we say that $(Y,Z)^{\top}$ is jointly symmetric about $(a,b)$.
+Remarks 1.
+  1. 1.
+    We also describe the joint distribution function (df) $F_{YZ}$ of $(Y,Z)^{\top}$ as having these symmetries.
+  2. 2.
+    If (i) holds, the marginal distribution of $Z$ is symmetric about $b$; if (ii) holds, the marginal distribution of $Y$ is symmetric about $a$. The reverse implications are not true.
+  3. 3.
+    H-symmetry and v-symmetry together imply radial and joint symmetry.
+However radial symmetry does not necessarily imply h-symmetry and v-symmetry, although it does imply symmetry of both marginal distributions.
+  4. 4.
+    Nelsen, (2006) notes that jointly symmetric distributions have zero correlation (when finite second moments permit the measure to be defined).
+In Section S1 of the Online Supplement, we show that h-symmetry (or v-symmetry) on its own is sufficient for an absence of both correlation and rank correlation.
+Despite the absence of correlation, distributions with these forms of symmetry can have quite strong forms of dependence.
+If the distribution functions $F_{Y}$ and $F_{Z}$ of $Y$ and $Z$ are continuous, the random vector $(U,V)^{\top}=(F_{Y}(Y),F_{Z}(Z))^{\top}$ has distribution function $C$, known as the copula of $(Y,Z)^{\top}$, and the symmetries have the following implications for $C$.
+Lemma 1.
+Let $(Y,Z)^{\top}$ be a random vector with continuous marginal distributions and copula $C$.
+  1. (i)
+    If $(Y,Z)^{\top}$ is h-symmetric then C(u,v)=uC(u,1v)C(u,v)=u-C(u,1-v), for $(u,v)\in[0,1]^{2}$.
+  2. (ii)
+    If $(Y,Z)^{\top}$ is v-symmetric then C(u,v)=vC(1u,v)C(u,v)=v-C(1-u,v), for $(u,v)\in[0,1]^{2}$,
+  3. (iii)
+    If $(Y,Z)^{\top}$ is radially symmetric then C(u,v)=u+v+C(1u,1v)1C(u,v)=u+v+C(1-u,1-v)-1 , for $(u,v)\in[0,1]^{2}$.
+Proof.
+The proof of (iii) is given by Nelsen, (2006) (Theorem 2.7.3) and (i) and (ii) follow by a similar method. ∎
+This result implies that the random vector $(U,V)^{\top}=(F_{Y}(Y),F_{Z}(Z))^{\top}$, or its df $C$, inherits the properties of h-symmetry, v-symmetry, radial and joint symmetry from $(Y,Z)^{\top}$; the corresponding lines and centre of symmetry are at 0.5 and $(0.5,0.5)$.
+The next result provides the first connection between symmetries of copula functions and v-transforms.
+The function $\mathcal{V}(u)=|2u-1|$ appearing in this result is the simplest example of a v-transform.
+Proposition 1.
+Let $C$ be the distribution function of $(U,V)^{\top}$. Then, for all $(u,v)\in[0,1]^{2}$,
+  1. (i)
+    if $C$ is h-symmetric, $C(u,v)=(u+(-1)^{I_{\{v<0.5\}}}C^{(h)}(u,|2v-1|))/2$ where $C^{(h)}$ is the df of $(U,|2V-1|)^{\top}$;
+  2. (ii)
+    if $C$ is v-symmetric, $C(u,v)=(v+(-1)^{I_{\{u<0.5\}}}C^{(v)}(|2u-1|,v))/2$ where $C^{(v)}$ is the df of $(|2U-1|,V)^{\top}$;
+  3. (iii)
+    if $C$ is jointly symmetric, $C(u,v)=(2u+2v-1+(-1)^{I_{\{(u-0.5)(v-0.5)<0\}}}C^{*}(|2u-1|,|2v-1|))/4$ where $C^{*}$ is the df of $(|2U-1|,|2V-1|)^{\top}$.
+Proof.
+We prove (i) and (iii) since (ii) follows immediately from (i).
+We use the equalities $C^{(h)}(u,v)=\mathbb{P}(U\leqslant u,|2V-1|\leqslant v)=\mathbb{P}(U\leqslant u,|V-0.5|\leqslant v/2)=C(u,(1+v)/2)-C(u,(1-v)/2)$ and part (i) of Lemma 1 to infer that $C^{(h)}(u,|2v-1|)=C(u,1-v)-C(u,v)=u-2C(u,v)$ if $v<0.5$ and $C^{(h)}(u,|2v-1|)=C(u,v)-C(u,1-v)=2C(u,v)-u$ if $v\geqslant 0.5$.
+Hence part (i) follows.
+To prove (iii) note that, if $C$ is jointly symmetric, then $C^{(h)}$ is v-symmetric since $C^{(h)}(u,v)+C^{(h)}(1-u,v)=C(u,(1+v)/2)-C(u,(1-v)/2)+C(1-u,(1+v)/2)-C(1-u,(1-v)/2)=(1+v)/2-(1-v)/2=v$.
+Hence $C^{(h)}(u,|2v-1|)=(|2v-1|+(-1)^{I_{\{u<0.5\}}}C^{*}(|2u-1|,|2v-1|))/2$.
+In combination with part (i) this yields part (iii).
+If $C$ is the copula of the h-symmetric random vector $(Y,Z)^{\top}$ in part (i) of Lemma 1, $C^{(h)}$ in part (i) of Proposition 1 is the copula of $(Y,|Z-b|)^{\top}$.
+This follows by observing that if $V=F_{Z}(Z)$ is the probability-integral transform of $Z$, then the probability-integral transform of |Zb||Z-b| is $F_{|Z-b|}(|Z-b|)=|2V-1|$.
+Similarly, if $C$ is the copula of the jointly symmetric random vector $(Y,Z)^{\top}$, then $C^{*}$ in part (iii) of Proposition 1 is the copula of $(|Y-a|,|Z-b|)^{\top}$.
+In this paper we apply the symmetry concepts to random vectors $(Y,Z)^{\top}$ with a joint density $f_{Y,Z}(y,z)>0$ on $\mathbb{R}^{2}$, implying that the copula $C$ has density $c$ given by
+$c(u,v)=f_{Y,Z}(F_{Y}^{-1}(u),F_{Z}^{-1}(v))/\left(f_{Y}(F_{Y}^{-1}(u))f_{Z}(F_{Z}^{-1}(v))\right),$
+where $f_{Y}$ and $f_{Z}$ denote the marginal densities.
+It is easily inferred from Definition 2 that h-symmetry of $(Y,Z)^{\top}$ is equivalent to $f_{Y,Z}(y,b+z)=f_{Y,Z}(y,b-z)$ for all $(y,z)\in\mathbb{R}^{2}$ and some $b\in\mathbb{R}$, while v-symmetry and radial symmetry are equivalent to $f_{Y,Z}(a+y,z)=f_{Y,Z}(a-y,z)$ for some $a\in\mathbb{R}$ and $f_{Y,Z}(a+y,b+z)=f_{Y,Z}(a-y,b-z)$ for some $a,b\in\mathbb{R}$.
+Similarly, h-symmetry of the copula $C$ is equivalent to the identity c(u,v)=c(u,1v)c(u,v)=c(u,1-v), v-symmetry is equivalent to c(u,v)=c(1u,v)c(u,v)=c(1-u,v) and radial symmetry is equivalent to c(u,v)=c(1u,1v)c(u,v)=c(1-u,1-v).
+When densities exist, the following Corollary of Proposition 1 is immediate.
+Corollary 1.
+Let $C$ be the distribution function of $(U,V)^{\top}$ and assume $C$ has joint density $c$. For all $(u,v)\in[0,1]^{2}$,
+  1. (i)
+    if $C$ is h-symmetric, $c(u,v)=c^{(h)}(u,|2v-1|)$ where $c^{(h)}$ is the density of $(U,|2V-1|)^{\top}$;
+  2. (ii)
+    if $C$ is jointly symmetric, $c(u,v)=c^{*}(|2u-1|,|2v-1|)$ where $c^{*}$ is the density of $(|2U-1|,|2V-1|)^{\top}$.
+Simple calculations show that part (ii) also implies the following series of equations which will be used in the sequel:
+$c^{*}(u,v)=c\left(\frac{1-u}{2},\frac{1-v}{2}\right)=c\left(\frac{1+u}{2},\frac{1-v}{2}\right)=c\left(\frac{1-u}{2},\frac{1+v}{2}\right)=c\left(\frac{1+u}{2},\frac{1+v}{2}\right).$ (4)
 2.3 The first-order copula of GARCH(1,1)-type processes
+We can now apply these symmetry ideas to GARCH-type processes.
 Simple calculations show that the joint density of a pair of successive variables $(X_{1},X_{2})^{\top}$ in a process of GARCH(1,1) type is given by
@@
 where $f_{\sigma}$ is the marginal density of the volatility process $(\sigma_{t})_{t\in\mathbb{Z}}$ and $f_{\epsilon}$ is the density of the innovations.
-Applying (3), the density $c_{1}$ of the copula $C_{1}$ of $(X_{1},X_{2})^{\top}$ is thus
+The density $c_{1}$ of the copula $C_{1}$ of $(X_{1},X_{2})^{\top}$ is thus
 $c_{1}(u,v)=\frac{\int_{0}^{\infty}\frac{1}{s\sigma\left(F_{X}^{-1}(u),s\right)}f_{\epsilon}\left(\frac{F_{X}^{-1}(v)}{\sigma\left(F_{X}^{-1}(u),s\right)}\right)f_{\epsilon}\left(\frac{F_{X}^{-1}(u)}{s}\right)f_{\sigma}(s)\mathrm{d}s}{f_{X}\left(F_{X}^{-1}(u)\right)f_{X}\left(F_{X}^{-1}(v)\right)}$ (6)
@@
 $f_{X_{1},X_{2}}(x,y)$ $=\frac{1}{\sigma(x)}f_{\epsilon}\left(\frac{y}{\sigma(x)}\right)f_{X}(x),\quad\quad c_{1}(u,v)=\frac{1}{\sigma\left(F_{X}^{-1}(u)\right)}f_{\epsilon}\left(\frac{F_{X}^{-1}(v)}{\sigma\left(F_{X}^{-1}(u)\right)}\right)\frac{1}{f_{X}(F_{X}^{-1}(v))}\,.$ (8)
-The key insight from these calculations is the following.
+The key insight from these equations is the following.
 Proposition 2.
@@
   2. (ii)
-    An even volatility function also implies that $f_{X_{1},X_{2}}(x,y)=f_{X_{1},X_{2}}(-x,y)$ and so $(X_{1},X_{2})^{\top}$ is v-symmetric about zero and hence radially and jointly symmetric about the origin.
-In case (i) above we can apply Lemma 1 to infer that the copula $C_{1}$ is h-symmetric and we can apply Proposition 3 to infer that its density may be written in the form $c_{1}(u,v)=c_{X_{1},|X_{2}|}(u,|2v-1|)$ where $c_{X_{1},|X_{2}|}$ is the density of the copula of $(X_{1},|X_{2}|)^{\top}$.
-In case (ii) we infer that the copula $C_{1}$ is jointly symmetric and its density may be written in the form $c_{1}(u,v)=c_{|X_{1}|,|X_{2}|}(|2u-1|,|2v-1|)$ where $c_{|X_{1}|,|X_{2}|}$ is the density of the copula of $(|X_{1}|,|X_{2}|)^{\top}$.
-Despite the strong symmetry revealed by Proposition 4, there is one important form of symmetry that symmetric processes of GARCH(1,1) type do not have in general, namely exchangeability.
-It is not true that $f_{X_{1},X_{2}}(x,y)=f_{X_{1},X_{2}}(y,x)$ for all $(x,y)^{\top}\in\mathbb{R}^{2}$ so the bivariate distributions of the pairs $(X_{1},X_{2})^{\top}$ and $(X_{2},X_{1})^{\top}$ differ.
-This implies that ARCH and GARCH are not time-reversible models.
-In general $f_{X_{2}\mid X_{1}}(y\mid x)\neq f_{X_{1}\mid X_{2}}(y\mid x)$ and thus, in an intuitive sense, they are models in which the predictive distribution of the future given the past differs from the “predictive distribution” of the past given the future.
+    If the volatility function $\sigma(x,s)$ is an even function of $x$, this also implies that $f_{X_{1},X_{2}}(x,y)=f_{X_{1},X_{2}}(-x,y)$ and so $(X_{1},X_{2})^{\top}$ is v-symmetric about zero and hence radially and jointly symmetric about the origin.
+In case (i) above we can apply Lemma 1 to infer that the copula $C_{1}$ is h-symmetric and we can apply Proposition 1 and Corollary 1 to infer that its density may be written in the form $c_{1}(u,v)=c_{1}^{(h)}(u,|2v-1|)$ where $c_{1}^{(h)}$ is the density of the copula of $(X_{1},|X_{2}|)^{\top}$.
+In case (ii) we infer that the copula $C_{1}$ is jointly symmetric and its density may be written in the form $c_{1}(u,v)=c_{1}^{*}(|2u-1|,|2v-1|)$ where $c_{1}^{*}$ is the density of the copula of $(|X_{1}|,|X_{2}|)^{\top}$.
+In the latter case it follows from (4) that $c_{1}^{*}(u,v)=c_{1}((u+1)/2,(v+1)/2)$, which means that the copula density $c_{1}^{*}$ may be viewed as a blown-up version of the upper quadrant of the copula density $c_{1}$, i.e.
+the part of the density on $0.5,1]^{2}$.
+Despite the strong symmetry revealed by Proposition 2, the bivariate distribution of $(X_{1},X_{2})^{\top}$ in a symmetric process of GARCH(1,1) type is not in general exchangeable.
+This implies that ARCH and GARCH are not time-reversible models; in general $f_{X_{2}\mid X_{1}}(y\mid x)\neq f_{X_{1}\mid X_{2}}(y\mid x)$ and thus, in an intuitive sense, they are models in which the predictive distribution of the future given the past differs from the “predictive distribution” of the past given the future.
 From an econometric point of view, this would be considered natural.
-The non-exchangeability carries over to the copula so that we have in general that $C_{1}(u,v)\neq C_{1}(v,u)$ on $[0,1]^{2}$.
-Despite this observation, it is notable that the majority of the copulas employed in vine copula models of financial time series are exchangeable ones.
+The non-exchangeability carries over to the copula so that we have in general that $C_{1}(u,v)\neq C_{1}(v,u)$ on $0,1]^{2}$.
 2.4 Computational examples
-In this section we use computational methods to approximate and plot the joint densities $f_{X_{1},X_{2}}$ and copula densities $c_{1}$ in a number of models of GARCH(1,1) type.
+In this section we use computational methods to approximate and plot the copula density $c_{1}$ in a number of models of GARCH(1,1) type; in Section S2 of the Online Supplement we give some accompanying pictures of the joint density $f_{X_{1},X_{2}}$ for selected models.
 We concentrate on processes of ARCH(1) type in our illustrations since their first-order Markov structure simplifies computation of the unknown marginal density $f_{X}$ and distribution function $F_{X}$, which are key ingredients in (8).
@@
 To analyse this process we use simulation and estimation to find a parametric model for the unknown marginal density $f_{\sigma}$ of the volatility process.
-We can then integrate out volatility in expressions like (6), (7) and (8) [^1] All details of the model for $f_{\sigma}$ and the discretization method for ARCH(1)-type processes can be found in the code that accompanies this paper..
+We can then integrate out volatility in expressions like (5), (6) and (7) [^2] All details of the model for $f_{\sigma}$ and the discretization method for ARCH(1)-type processes can be found in the code that accompanies this paper at https://github.com/ajmcneil/papers..
 The parameter values for all the models we consider are summarized in Table 1.
+In the examples that follow we concentrate on illustrating the copulas densities $c_{1}$; corresponding illustrations of the joint densities $f_{X_{1},X_{2}}$ can be found in Section S2 of the Supplementary Material.
 Process $\alpha_{0}$ $\alpha_{1}$ $\beta_{1}$ $\gamma_{1}$
@@
 The volatility function $\sigma(x,s)$ in all models satisfies $\sigma(x,s)^{2}=\alpha_{1}+(\alpha_{1}+\gamma_{1}I_{\{x<0\}})x^{2}+\beta_{1}s^{2}$.
-Innovation distributions used are either Gaussian, Student t with $\nu=4$ or $\nu=2.5$ degrees of freedom or skewed Student t with $\nu=4$ degrees of freedom and skewness parameter $\lambda=0.8$; all innovation distributions are scaled to have mean zero and variance one.
-Figure 1: Contour plots of $f_{X_{1},X_{2}}(x_{1},x_{2})$ for symmetric GARCH(1,1)-type processes; see Table 1 for all parameter values.
-Top left, top right, bottom left: classical ARCH(1) processes with, respectively, Gaussian, Student t4 and Student t2.5 innovations.
-Bottom right: classical GARCH(1,1) process with Gaussian innovations.
+Innovation distributions used are either Gaussian, Student t with $\xi=4$ or $\xi=2.5$ degrees of freedom or skewed Student t with $\xi=4$ degrees of freedom and skewness parameter $\lambda=0.8$; all innovation distributions are scaled to have mean zero and variance one.
 Figure 1: Contour plots of $c_{1}(u,v)$ for symmetric GARCH(1,1)-type processes; see Table 1 for all parameter values.
@@
 Bottom right: classical GARCH(1,1) process with Gaussian innovations.
-  1. 1.
-    Classical ARCH(1) process with Gaussian innovations (symmetric) and parameter values $\alpha_{0}=0.4$ and $\alpha_{1}=0.6$ as in Table 1; this process has unit variance and tail index $\zeta=3.82$.
-The joint density $f_{X_{1},X_{2}}$ and copula density $c_{1}$ are shown in the top left panels of Figures 1 and 2 respectively.
-The non-exchangeability, evidenced by a lack of symmetry in the dashed lines $x_{2}=x_{1}$ and $v=u$, is evident in both pictures.
-We can also observe the joint symmetry, that is the rotational symmetry of order 2 as well as the h-symmetry and v-symmetry.
-    Changing $\alpha_{1}$ changes the degree of dependence in the pictures.
-In particular, as $\alpha_{1}\to 0$ the copula converges to the independence copula while increasing $\alpha_{1}$ serves to strengthen the cross-shaped character of the copula density.
-For normal innovations, the theoretical maximum value of $\alpha_{1}$ to obtain a strictly stationary process is approximately 3.56, but for values $\alpha_{1}\geqslant 1$ we obtain infinite-variance processes which are seldom used in practical applications and become increasingly pathological with increasing $\alpha_{1}$ as their tail indexes converge to zero (Mikosch and Stărică,, 2000).
-  2. 2.
-    Classical ARCH(1) process with Student innovations (symmetric).
-This is shown in the top right panels of Figures 1 and 2 for a process with the same parameters as in the previous example and degree of freedom $\nu=4$ for the innovation distribution; the tail index of the process is 2.76.
-Remarkably, the resulting joint density appears to be very close to spherical and the copula appears to be very close to an exchangeable copula for this combination of parameters; the joint symmetry is again apparent.
-The bottom left panels of Figures 1 and 2 show a process in which $\nu=2.5$ and the tail index is 2.09; these show non-exchangeability more clearly.
-    It is notable that, although the tail index of the marginal distribution increases with decreasing degree of freedom (increasing tail index of the innovation density), the dependence actually weakens.
+2.4.1 Classical ARCH(1) process with Gaussian innovations (symmetric)
+For the parameter values in Table 1, the contours of the copula density $c_{1}$ are shown in the top left panel of Figure 1 The non-exchangeability, evidenced by a lack of symmetry in the dashed line $v=u$, is evident.
+We can also observe the joint symmetry, that is the h-symmetry and v-symmetry also implying rotational symmetry of order 2.
+Changing $\alpha_{1}$ would change the degree of dependence; in particular, increasing $\alpha_{1}$ would strengthen the cross-shaped character of the copula density.
+In Section S2 of the Supplementary Material we also include a contour plot of the absolute value copula for this model, that is the copula of $(|X_{1}|,|X_{2}|)^{\top}$.
+The picture for the absolute value copula is confirmed to be a blown-up version of the top right quadrant of the contour plot of the copula density $c_{1}$.
+2.4.2 Classical ARCH(1) process with Student innovations (symmetric)
+This is shown in the top right panel of Figure 1 for a process with the same parameters as in the previous example and degree of freedom $\xi=4$ for the innovation distribution.
+Remarkably, the copula appears to be very close to an exchangeable copula for this combination of parameters; the joint symmetry is again apparent.
+The bottom left panel of Figure 1 (see also Figure S1) show a process in which $\xi=2.5$ which shows non-exchangeability more clearly.
+It is notable that, although the heaviness of the tail of the marginal distribution increases with decreasing degree of freedom, the dependence actually weakens.
 We can estimate a distance $D_{1}=\int_{0}^{1}\int_{0}^{1}|c_{1}(u,v)-1|\mathrm{d}u\mathrm{d}v$ between the copula density $c_{1}$ and the density of the independence copula; for the Gaussian, t4 and t2.5 copulas we obtain respectively 0.213, 0.158 and 0.092.
-  3. 3.
-    Classical GARCH(1,1) process with Gaussian innovations (symmetric).
-This is shown in the bottom right panels of Figures 1 and 2 for a process with Gaussian innovations and parameters as in Table 1, which has variance one and tail index 4.09.
-    It is well known that, for a covariance stationary GARCH(1,1) process ($\alpha_{1}+\beta_{1}<1$), the squared process $(X_{t}^{2})_{t\in\mathbb{Z}}$ behaves like an ARMA(1,1) process with autoregressive parameter $\alpha_{1}+\beta_{1}$ and moving-average parameter $-\beta_{1}$; see, for example, McNeil et al., (2015, Section 4.2.2).
+2.4.3 Classical GARCH(1,1) process with Gaussian innovations (symmetric)
+This is shown in the bottom right panel of Figure 1 for a process with Gaussian innovations.
+It is well known that, for a covariance stationary GARCH(1,1) process ($\alpha_{1}+\beta_{1}<1$), the squared process $(X_{t}^{2})_{t\in\mathbb{Z}}$ behaves like an ARMA(1,1) process with autoregressive parameter $\alpha_{1}+\beta_{1}$ and moving-average parameter $\beta_{1}$; see, for example, McNeil et al., (2015, Section 4.2.2).
 Thus the value of $\alpha_{1}+\beta_{1}$ controls the persistence, or the rate at which serial dependencies in the squared (or absolute value) process die away with lag.
@@
 The main point is to show that the behaviour of the GARCH(1,1) copula is qualitatively similar to that of the ARCH(1) copula.
-  4. 4.
-    ARCH(1) model with Gaussian innovations and leverage (symmetric distribution).
-Here we use a volatility function $\sigma(x)$ satisfying $\sigma(x)^{2}=\alpha_{0}+(\alpha_{1}+\gamma_{1}I_{\{x<0\}})x^{2}$ for $\gamma_{1}>0$ so that a negative return produces a larger value of the volatility than a positive return of the same magnitude.
-From now on, we concentrate on contour plots of the copula only, as shown in the top left panel of Figure 3.
+2.4.4 ARCH(1) model with Gaussian innovations and leverage (symmetric distribution)
+In the left panel of Figure 2 we show the copula $c_{1}$ for a model with symmetric innovations and volatility function $\sigma(x)$ satisfying $\sigma(x)^{2}=\alpha_{0}+(\alpha_{1}+\gamma_{1}I_{\{x<0\}})x^{2}$ for $\gamma_{1}>0$, so that a negative return produces a larger value of the volatility than a positive return of the same magnitude.
 It can be seen that leverage removes radial symmetry and v-symmetry, but there is still h-symmetry (symmetry in the line $v=0.5$).
-  5. 5.
-    Classical ARCH(1) model with skewed Student innovations (asymmetric distribution but symmetric volatility).
-This is shown in the top right panel of Figure 3 for an innovation distribution that is skewed to the left (towards negative shocks).
+2.4.5 Classical ARCH(1) model with skewed Student innovations (symmetric volatility)
+The copula density $c_{1}$ is shown in the right panel of Figure 2 for a model with an innovation distribution that is skewed to the left (towards negative shocks).
 The skewness removes all the previous symmetries in the pictures.
-  6. 6.
-    Absolute value copula for classical ARCH(1) model with Gaussian innovations.
-This is shown in the bottom left of Figure 3 for the same process as in the top left of Figure 2.
-It can be thought of as a blown-up version of the top right quadrant of the latter and clearly shows non-exchangeability and stronger dependence in the upper tail.
-Note that this picture also illustrates the density of the volatility copula of ARCH(1), that is the copula of $(\sigma_{1},\sigma_{2})^{\top}$; this follows because $\sigma_{t}$ is a strictly increasing function of $|X_{t-1}|$ for all $t$ and copulas are invariant under strictly increasing transformations of variables.
-For the same reason, it is also the copula density for the variables $(X_{1}^{2},X_{2}^{2})^{\top}$.
 Figure 2: Introducing asymmetry.
-Top left: ARCH(1) with Gaussian innovations in which the volatility function has leverage and the parameters are $\alpha_{0}=0.4$, $\alpha_{1}=0.3$ and $\gamma_{1}=0.4$.
-Top right: classical ARCH(1) model with with 4 degrees of freedom and skewness parameter $\lambda=0.8$ where $\alpha_{0}=0.4$ and $\alpha_{1}=0.6$.
-Bottom left: copula density for copula of the absolute values for classical ARCH(1) copula with Gaussian innovations.
-Bottom right: AR(1)-ARCH(1) model with Gaussian innovations and parameters $\phi=0.3$, $\alpha_{0}=0.4$ and $\alpha_{1}=0.6$.
-3.4 AR(1) model with ARCH(1) errors
-We consider briefly a time series model with stochastic volatility as well as a non-zero stochastic conditional mean term.
-We consider an AR(1) model with ARCH(1) errors and show the implications of the conditional mean term.
-Specifically, we consider the first-order Markov process $(X_{t})_{t\in\mathbb{Z}}$ satisfying equations
-$X_{t}=\phi X_{t-1}+\sigma(X_{t-1})\;\epsilon_{t}\quad\text{with}\quad\sigma(x)=\sqrt{\alpha_{0}+\alpha_{1}x^{2}},$ (11)
-where the innovations $(\epsilon_{t})_{t\in\mathbb{Z}}$ have a symmetric distribution.
-This model was proposed and analysed by Borkovec and Klüppelberg, (2001) [^2] It differs from the usual version of the AR(1)-ARCH(1) model, $X_{t}=\phi X_{t-1}+\sigma(X_{t-1}-\phi X_{t-2})\;\epsilon_{t}$, but this is not first-order Markov and thus less straightforward to treat with our methods.
-who showed that it is covariance stationary provided $\phi^{2}+\alpha_{1}<1$ and has a symmetric stationary distribution $F_{X}$.
-The formulas (9) require only simple modification to account for the non-zero conditional mean term. In particular, the copula becomes
-$c_{1}(u,v)=\frac{1}{\sigma\left(F_{X}^{-1}(u)\right)}f_{\epsilon}\left(\frac{F_{X}^{-1}(v)-\phi F_{X}^{-1}(u)}{\sigma\left(F_{X}^{-1}(u)\right)}\right)\frac{1}{f_{X}(F_{X}^{-1}(v))}\,.$
-It may be easily verified that the copula density satisfies $c(u,v)=c(1-u,1-v)$ and so the copula is radially symmetric.
-However, the copula is neither h-symmetric nor v-symmetric, and thus it is not jointly symmetric.
-The copula density is shown in the bottom right panel of Figure 3 in the case where $\phi=0.3$, $\alpha_{0}=0.4$ and $\alpha_{1}=0.6$.
-It is evident that more of the probability mass is now on the main diagonal of the copula.
-In the remainder of the paper we concentrate on models without a conditional mean term, but this example gives some clues about the effect on the copula of serial dependence in the untransformed process $(X_{t})_{t\in\mathbb{Z}}$ in combination with stochastic volatility, a point we return to in our discussion in Section 7.
-3.5 Higher-order conditional copulas of GARCH(1,1)-type processes
-For higher-order copulas we introduce the compact notation $\bm{X}_{[j,l]}=(X_{j},\ldots,X_{l})^{\top}$ to denote random vectors consisting of a contiguous block of random variables; we apply the same notation to realisations of the random vectors.
-Whenever the notation is used with $j>l$, this means that the variables are taken in reverse order to their natural temporal order.
-Now we consider the copula $C_{k}$ of the conditional distribution of $(X_{1},X_{k+1})^{\top}$ given $\bm{X}_{[2:k]}=\bm{x}_{[2:k]}$ for $k\geqslant 2$.
-The density of the conditional copula takes the form
-$c_{k}(u,v\mid\bm{x}_{[2:k]})=\frac{f_{X_{1},X_{k+1}\mid\bm{X}_{[2:k]}}\left(F_{X_{1}\mid\bm{X}_{[2:k]}}^{-1}\left(u\mid\bm{x}_{[2:k]}\right),F_{X_{k+1}\mid\bm{X}_{[2:k]}}^{-1}\left(v\mid\bm{x}_{[2:k]}\right)\mid\bm{x}_{[2:k]}\right)}{f_{X_{1}\mid\bm{X}_{[2:k]}}\left(F_{X_{1}\mid\bm{X}_{[2:k]}}^{-1}\left(u\mid\bm{x}_{[2:k]}\right)\mid\bm{x}_{[2:k]}\right)f_{X_{k+1}\mid\bm{X}_{[2:k]}}\left(F_{X_{k+1}\mid\bm{X}_{[2:k]}}^{-1}\left(v\mid\bm{x}_{[2:k]}\right)\mid\bm{x}_{[2:k]}\right)}$
-Note that this copula density will simply be the independence copula density for a process of ARCH(1) type, since $X_{1}$ and $X_{3}$ are conditionally independent given $X_{2}$.
-Consider, therefore a process with an explicit GARCH effect, i.e.
-$\sigma(x,s)$ depends also on $s$.
-To develop intuition we consider the case where $k=2$.
-The symmetry properties can be investigated by considering the joint conditional density $f_{X_{1},X_{3}\mid X_{2}}(x,y\mid z)$.
-To this end consider the joint density
-$f_{X_{1},X_{2},X_{3}}(x,z,y)$ $=\int_{0}^{\infty}f_{X_{1},X_{2},X_{3},\sigma_{1}}(x,z,y,s)\mathrm{d}s$
-$=\int_{0}^{\infty}f_{X_{3}\mid X_{2},X_{1},\sigma_{1}}(y\mid z,x,s)f_{X_{2}\mid X_{1},\sigma_{1}}(z\mid x,s)f_{X_{1}\mid\sigma_{1}}(x\mid s)f_{\sigma}(s)\mathrm{d}s$
-$=\int_{0}^{\infty}\frac{1}{\sigma(z,\sigma(x,s))}f_{\epsilon}\left(\frac{y}{\sigma(z,\sigma(x,s))}\right)\frac{1}{\sigma(x,s)}f_{\epsilon}\left(\frac{z}{\sigma(x,s)}\right)\frac{1}{s}f_{\epsilon}\left(\frac{x}{s}\right)f_{\sigma}(s)\mathrm{d}s.$
-It is straightforward to see that a process with symmetric innovations will satisfy $f_{X_{1},X_{2},X_{3}}(x,z,-y)=f_{X_{1},X_{3},X_{2}}(x,z,y)$ while a symmetric process will also satisfy $f_{X_{1},X_{2},X_{3}}(-x,z,y)=f_{X_{1},X_{2},X_{3}}(x,z,y)$.
-These properties will also hold for the conditional density $f_{X_{1},X_{3}\mid X_{2}}(x,y\mid z)=f_{X_{1},X_{2},X_{3}}(x,z,y)/f_{X}(z)$ and thus, by Lemma 1, the various forms of symmetry will be inherited by the conditional copula density $c_{2}(u,v\mid z)$.
-By applying a similar analysis to the joint density $f_{X_{1},\ldots,X_{k+1}}(x,\ldots,y)$ we can therefore establish the following result.
+Left: ARCH(1) with Gaussian innovations in which the volatility function has leverage and the parameters are $\alpha_{0}=0.4$, $\alpha_{1}=0.3$ and $\gamma_{1}=0.4$.
+Right: classical ARCH(1) model with 4 degrees of freedom and skewness parameter $\lambda=0.8$ where $\alpha_{0}=0.4$ and $\alpha_{1}=0.6$.
+2.5 Higher-order copulas of GARCH(1,1)-type processes
+In this section we show that the same symmetries evident in the copula density $c_{1}$ can be found in the density $c_{k}$ of the copula of the conditional distribution of $(X_{1},X_{k+1})^{\top}$ given $\bm{X}_{[2:k]}=\bm{x}_{[2:k]}$ for $k\geqslant 2$, which takes the form
+$c_{k}(u,v\mid\bm{x}_{[2:k]})=\frac{f_{X_{1},X_{k+1}\mid\bm{X}_{[2:k]}}\left(F_{X_{1}\mid\bm{X}_{[2:k]}}^{-1}\left(u\mid\bm{x}_{[2:k]}\right),F_{X_{k+1}\mid\bm{X}_{[2:k]}}^{-1}\left(v\mid\bm{x}_{[2:k]}\right)\mid\bm{x}_{[2:k]}\right)}{f_{X_{1}\mid\bm{X}_{[2:k]}}\left(F_{X_{1}\mid\bm{X}_{[2:k]}}^{-1}\left(u\mid\bm{x}_{[2:k]}\right)\mid\bm{x}_{[2:k]}\right)f_{X_{k+1}\mid\bm{X}_{[2:k]}}\left(F_{X_{k+1}\mid\bm{X}_{[2:k]}}^{-1}\left(v\mid\bm{x}_{[2:k]}\right)\mid\bm{x}_{[2:k]}\right)}$ (10)
+It should be distinguished from the copula density $c_{(k)}$ of the copula $C_{(k)}$ of the marginal distribution of $(X_{1},X_{k+1})^{\top}$ (unless $k=1$).
+In particular, it is the densities $c_{k}$ rather than $c_{(k)}$ that appear in the D-vine representation (1) of the GARCH-type process.
+The former depend on the conditioning variables $\bm{x}_{[2:k]}$ although we would like to be able to assume that this dependence is weak so that the D-vine representation can be approximated by a simplified D-vine.
+To understand the main argument in this section it suffices to consider the case $k=2$.
+We consider a process with an explicit GARCH effect since, for a process of ARCH(1) type, $X_{1}$ and $X_{3}$ are conditionally independent given $X_{2}=z$ and $c_{2}(u,v\mid z)\equiv 1$, the density of the independence copula.
+The joint conditional density $f_{X_{1},X_{3}\mid X_{2}}(x,y\mid z)$ and the joint density $f_{X_{1},X_{3}}(x,y)$ can both be calculated from the joint density $f_{X_{1},X_{2},X_{3}}(x,z,y)$ by observing that $f_{X_{1},X_{3}\mid X_{2}}(x,y\mid z)=f_{X_{1},X_{2},X_{3}}(x,z,y)/f_{X}(z)$ and $f_{X_{1},X_{3}}(x,y)=\int_{-\infty}^{\infty}f_{X_{1},X_{2},X_{3}}(x,z,y)\,\mathrm{d}z$.
+Since
+$f_{X_{1},X_{2},X_{3}}(x,z,y)$ $=\int_{0}^{\infty}f_{X_{3}\mid X_{2},X_{1},\sigma_{1}}(y\mid z,x,s)f_{X_{2}\mid X_{1},\sigma_{1}}(z\mid x,s)f_{X_{1}\mid\sigma_{1}}(x\mid s)f_{\sigma}(s)\mathrm{d}s$
+$=\int_{0}^{\infty}\frac{1}{\sigma(z,\sigma(x,s))}f_{\epsilon}\left(\frac{y}{\sigma(z,\sigma(x,s))}\right)\frac{1}{\sigma(x,s)}f_{\epsilon}\left(\frac{z}{\sigma(x,s)}\right)\frac{1}{s}f_{\epsilon}\left(\frac{x}{s}\right)f_{\sigma}(s)\mathrm{d}s$
+we can infer that $f_{X_{1},X_{2},X_{3}}(x,z,-y)=f_{X_{1},X_{3},X_{2}}(x,z,y)$ for a process with symmetric innovations and $f_{X_{1},X_{2},X_{3}}(-x,z,y)=f_{X_{1},X_{2},X_{3}}(x,z,y)$ for a symmetric process.
+It follows immediately that the conditional and unconditional densities satisfy $f_{X_{1},X_{3}\mid X_{2}}(x,-y\mid z)=f_{X_{1},X_{3}\mid X_{2}}(x,y\mid z)$ and $f_{X_{1},X_{3}}(x,-y)=f_{X_{1},X_{3}}(x,y)$ for a process with symmetric innovations and we have h-symmetry in both cases.
+Moreover, they satisfy $f_{X_{1},X_{3}\mid X_{2}}(-x,y\mid z)=f_{X_{1},X_{3}\mid X_{2}}(x,y\mid z)$ and $f_{X_{1},X_{3}}(-x,y)=f_{X_{1},X_{3}}(x,y)$ for a symmetric process and we have joint symmetry in both cases.
+By Lemma 1, the various forms of symmetry are inherited by the conditional copula $C_{2}$ of $(X_{1},X_{3})^{\top}$ given $X_{2}$ and the marginal copula $C_{(2)}$ of $(X_{1},X_{3})^{\top}$.
+We now give the general result for $k\geqslant 2$.
+Since the logic of its proof is identical to the argument given above, albeit with more cumbersome notation, the general proof is given in the Supplementary Material.
 Proposition 3.
   1. (i)
-    In a GARCH(1,1)-type process with symmetric distribution, the conditional copula $C_{k}$ is h-symmetric about zero for $k\geqslant 2$.
+    In a GARCH(1,1)-type process with symmetric distribution, the conditional copula $C_{k}$ and the marginal copula $C_{(k)}$ are h-symmetric about zero for $k\geqslant 2$.
   2. (ii)
-    In a symmetric GARCH(1,1)-type process the conditional copula is jointly symmetric about the origin for $k\geqslant 2$.
-Proof.
-Suppose we write $s_{2}(x,s)=\sigma(x,s)$, $s_{3}(z_{1},x,s)=\sigma(z_{1},s_{2}(x,s))$ and then for $j=4,\ldots,k+1$ we define recursive functions $s_{j}(z_{j-2},\ldots,z_{1},x,s)=\sigma(z_{j-2},s_{j-1}(z_{j-3},\ldots,z_{1},x,s))$.
-This permits us to write the joint density as
-$f_{X_{1},\ldots,X_{k+1}}(x,z_{1},\ldots,z_{k-1},y)=\int_{0}^{\infty}f_{X_{k+1}\mid\sigma_{k+1}}(y\mid s_{k+1}(z_{k-1},\ldots,z_{1},x,s))\\ \left(\prod_{j=3}^{k}f_{X_{j}\mid\sigma_{j}}(z_{j-1}\mid s_{j}(z_{j-2},\ldots,z_{1},x,s))\right)f_{X_{2}\mid\sigma_{2}}(z_{1}\mid s_{2}(x,s))f_{X_{1}\mid\sigma_{1}}(x\mid s)f_{\sigma}(s)\mathrm{d}s$
-It is clear that this joint density satisfies $f_{X_{1},\ldots,X_{k+1}}(x,\ldots,y)=f_{X_{1},\ldots,X_{k+1}}(x,\ldots,-y)$ when the innovation distribution is symmetric.
-Moreover it is also clear that in the symmetric model we have $f_{X_{1},\ldots,X_{k+1}}(x,\ldots,y)=f_{X_{1},\ldots,X_{k+1}}(-x,\ldots,y)$ since all of the functions $s_{2},\ldots,s_{k+1}$ are even functions of $x$.
-These properties will also hold for the conditional density
-$f_{X_{1},X_{k+1}\mid\bm{X}_{[2:k]}}(x,y\mid z_{1},\ldots,z_{k-1})=f_{X_{1},\ldots,X_{k+1}}(x,z_{1},\ldots,z_{k-1},y)/f_{X_{2},\ldots,X_{k}}(z_{1},\ldots,z_{k-1})$
-and thus we can apply Lemma 1 to infer that the assertions regarding symmetry of the copula hold. ∎
-Remark 1.
-Since the joint density of $(X_{1},X_{k+1})^{\top}$ can be obtained as
-$f_{X_{1},X_{k+1}}(x,y)=\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}f_{X_{1},\ldots,X_{k+1}}(x,z_{1},\ldots,z_{k-1},y)\mathrm{d}z_{1}\cdots\mathrm{d}z_{k-1}$
-it follows that the same method of proof can be used to show that the copula of the unconditional distribution of $(X_{1},X_{k+1})^{\top}$ shares the same symmetry properties as the conditional copula.
-An interesting question concerns the degree to which $c_{k}(u,v\mid\cdot)$ depends on the realized values of the conditioning variables.
-We investigate this question for $k=2$ and and write $c_{2}(u,v;w):=c_{2}(u,v\mid F_{X}^{-1}(w))$.
-The latter function takes the form
-$c_{2}(u,v;w)=\frac{f_{X_{1},X_{3}\mid X_{2}}\left(F_{X_{1}\mid X_{2}}^{-1}\left(u\mid F_{X}^{-1}(w)\right),F_{X_{3}\mid X_{2}}^{-1}\left(v\mid F_{X}^{-1}(w)\right)\mid F_{X}^{-1}(w)\right)}{f_{X_{1}\mid X_{2}}\left(F_{X_{1}\mid X_{2}}^{-1}\left(u\mid F_{X}^{-1}(w)\right)\mid F_{X}^{-1}(w)\right)f_{X_{3}\mid X_{2}}\left(F_{X_{3}\mid X_{2}}^{-1}\left(v\mid F_{X}^{-1}(w)\right)\mid F_{X}^{-1}(w)\right)}$
-and can be calculated approximately for the classical GARCH(1,1) model building on our analysis in Section 3.3.
-In Figure 4 we show contour plots for the cases where $w=0.5$, i.e.
-the intermediate observation is at the median of the marginal distribution (zero in this case), and $w=0.75$, i.e.
-the intermediate observation is at a quartile of the marginal distribution.
-While the two pictures are not identical, they are relatively similar to each other in terms of the shapes and heights of the contour lines.
-To examine this further we attempt to compute an approximate value for the absolute distance between the conditional copula densities and the independence copula density using the measure $D_{2}(w)=\int_{0}^{1}\int_{0}^{1}|c_{2}(u,v;w)-1|\mathrm{d}u\mathrm{d}v$.
-We obtain the estimates $D_{2}(0.5)=0.10$ and $D_{2}(0.75)=0.09$, which can be compared with the equivalent measure $D_{1}=0.16$ for the copula density $c_{1}$ in Figure 2.
+    In a symmetric GARCH(1,1)-type process $C_{k}$ and $C_{(k)}$ are jointly symmetric about the origin for $k\geqslant 2$.
+The computational tools we have developed to approximate $c_{1}$ in Section 2.4 can be extended to approximate $c_{2}$, and we use them to explore the simplifying assumption.
+Let $c_{2}(u,v;w):=c_{2}(u,v\mid F_{X}^{-1}(w))$ for $w\in(0,1)$.
+Contour plots of $c_{1}(u,v;w)$ for the cases where $w=0.5$ and $w=0.75$ are shown in Figure 3.
+The two pictures are very similar to each other in terms of the shapes and heights of the contour lines, and are also similar in pattern to the plot of $c_{1}$ in the lower-right panel of Figure 1.
+We compute an approximate value for the absolute distance between the conditional copula densities and the independence copula density using the measure $D_{2}(w)=\int_{0}^{1}\int_{0}^{1}|c_{2}(u,v;w)-1|\mathrm{d}u\mathrm{d}v$.
+We obtain the estimates $D_{2}(0.5)=0.10$ and $D_{2}(0.75)=0.09$, which can be compared with the equivalent measure $D_{1}=0.16$ for the copula density $c_{1}$ in the lower-right panel of Figure 1.
+Figure 3: Contour plots of $c_{2}(u,v;w)$ for the classical GARCH(1,1) process with Gaussian innovations. Left: $w=0.5$. Right: $w=0.75$.
 As might be expected, the dependence weakens from lag 1 to lag 2.
-Moreover, the value of the conditioning variable does not seem to have a particularly large effect on the conditional copula at lag 2.
-This might be taken as weak evidence that the simplifying assumption is approximately tenable; obviously we should look at values of $w$ further into the tail, but our approximation methods then become more unreliable.
-Figure 4: Contour plots of $c_{2}(u,v;w)$ for the classical GARCH(1,1) process with Gaussian innovations. In the left-hand figure $w=0.5$ and in the right-hand figure $w=0.75$.
-4 Copula densities and v-transforms
+Moreover, the value of $w$ does not seem to have a particularly large effect on the conditional copula at lag 2.
+This might be taken as weak evidence in favour of the simplifying assumption.
+It would be desirable to look at values of $w$ further into the tail and consider lags $k>2$, but our approximation methods then become less accurate leading to less confident conclusions.
+3 Using v-transforms to approximate GARCH pair copulas
 3.1 V-transforms and their inverses
-The function $\mathcal{V}(u)=|2u-1|$ which appeared in Proposition 3 is a symmetric v-shaped function and the simplest example of a v-transform.
-These are special cases of uniform-measure-preserving transformations (Porubský et al.,, 1988) and are extensively analysed in McNeil, (2021) and Bladt and McNeil, 2022b .
-In this section we recap the essential elements of the theory and examine its relevance for understanding and mimicking the behaviour of the copula of processes of GARCH(1,1) type.
+It has long been observed that serial dependence in financial asset-price return series, and in the models like GARCH that describe them, shows up as serial correlation of absolute or squared values.
+Loaiza-Maya et al., (2018) considered the componentwise application of symmetric v-shaped transformations $T$ (such as the absolute value function $T(x)=|x|$) to a pair of random variables $(Y,Z)^{\top}$ and derived an expression for the relationship of the copula $C_{T}$ of $(T(Y),T(Z))^{\top}$ to the copula $C$ of $(Y,Z)^{\top}$.
+They showed that the resulting copula $C_{T}$ involves the marginal distributions $F_{Y}$ and $F_{Z}$ except in the special case of symmetric margins.
+McNeil, (2021) observed that the effect on the copula of such transformations, as well as more general asymmetric transformations, can be isolated by introducing v-shaped transformations of uniform random variables that are uniform-distribution-preserving (Porubský et al.,, 1988).
+Importantly, for the purposes of this paper, these so-called v-transforms can be inverted stochastically in a way that also preserves uniform distributions.
+We use this inversion technique to create new copulas that show behaviour similar to GARCH pair copulas.
+In this section we recap the essential elements of the theory.
 Definition 3.
@@
 $\mathcal{V}(u)=\begin{cases}(1-u)-(1-\delta)\Psi\left(\frac{u}{\delta}\right)&u\leqslant\delta,\\ u-\delta\Psi^{-1}\left(\frac{1-u}{1-\delta}\right)&u>\delta,\end{cases}$ (11)
-for a parameter $10<\delta<$ 1, known as the fulcrum, and a continuous and strictly increasing distribution function $\Psi$ on $[0,1]$, known as the generator.
+for a parameter $0<\delta<1$, known as the fulcrum, and a continuous and strictly increasing distribution function $\Psi$ on $[0,1]$, known as the generator.
 Mappings of this form are always v-shaped with $\mathcal{V}(0)=\mathcal{V}(1)=1$ and $\mathcal{V}(\delta)=0$.
-They preserve uniformity in the sense that if $U\sim U(0,1)$, then $\mathcal{V}(U)\sim U(0,1)$.
-If the generator $\Psi$ is taken to be standard uniform, we obtain the linear v-transform, which satisfies $\mathcal{V}(u;\delta)=(\delta-u)/\delta$ for $u\leqslant\delta$ and $\mathcal{V}(u;\delta)=(u-\delta)/(1-\delta)$ for $u>\delta$; this includes the special case $\mathcal{V}(u;0.5)=|2u-1|$, which is the symmetric v-transform.
+They preserve uniformity in the sense that if $U\sim\mathcal{U}(0,1)$ is standard uniform, then $\mathcal{V}(U)\sim\mathcal{U}(0,1)$.
+If the generator is $\Psi(x)=x$ (the uniform df), we obtain the linear v-transform, which satisfies $\mathcal{V}(u;\delta)=(\delta-u)/\delta$ for $u\leqslant\delta$ and $\mathcal{V}(u;\delta)=(u-\delta)/(1-\delta)$ for $u>\delta$; this includes the special case $\mathcal{V}(u;0.5)=|2u-1|$, which is the symmetric v-transform.
 It is possible to introduce further parameters.
@@
 We also introduce the function $\Delta(y)=-1/\mathcal{V}^{\prime}(\mathcal{V}^{-1}(y))$ where $\mathcal{V}^{\prime}$ is the derivative of $\mathcal{V}$.
-If $U\sim U(0,1)$ then McNeil, (2021) shows that $\Delta(y)=\mathbb{P}(U\leqslant\delta\mid\mathcal{V}(U)=y)$ for $y\in(0,1]$; for the linear v-transform $\Delta(y)=\delta$, but otherwise $\Delta(y)$ depends explicitly on $y$.
+If $U\sim\mathcal{U}(0,1)$ then McNeil, (2021) shows that $\mathbb{P}(U\leqslant\delta\mid\mathcal{V}(U)=y)=\Delta(y)$ for $y\in(0,1]$; for the linear v-transform $\Delta(y)=\delta$, but otherwise $\Delta(y)$ depends explicitly on $y$.
 We can stochastically invert a v-transform in such a way that we also preserve uniformity of random variables.
-For $y\in(0,1]$, the stochastic inverse $\mathcal{V}^{\leftarrow}(y)$ of the v-transform $\mathcal{V}$ is defined to take the value $\mathcal{V}^{-1}(y)$ (from the left branch of the transform) with probability $\Delta(y)$ and the value $\mathcal{V}^{-1}(y)+y$ (from the right branch of the transform) with probability $1-\Delta(y)$; for $y=0$ we set $\mathcal{V}^{\leftarrow}(y)=\delta$.
-If $U\sim U(0,1)$ then it can also be shown that $\mathcal{V}^{\leftarrow}(U)\sim U(0,1)$.
-Note, however, that while $\mathcal{V}\circ\mathcal{V}^{\leftarrow}(y)=y$ holds for all $y\in[0,1]$, the equation $\mathcal{V}^{\leftarrow}\circ\mathcal{V}(u)=u$ is only true with probability $\Delta(u)$, if $u<\delta$, and with probability $1-\Delta(u)$, if $u>\delta$.
+For $y\in(0,1]$, the stochastic inverse $\mathcal{V}^{\leftarrow}(y,W)$ of the v-transform $\mathcal{V}$ is defined by taking a uniform random variable $W\sim\mathcal{U}(0,1)$ and setting $\mathcal{V}^{\leftarrow}(y,W)=\mathcal{V}^{-1}(y)$ if $W\leqslant\Delta(y)$ and $\mathcal{V}^{\leftarrow}(y,W)=\mathcal{V}^{-1}(y)+y$ otherwise; we always set $\mathcal{V}^{\leftarrow}(0,W)=\mathcal{V}^{-1}(0)=\delta$.
+If $U$ and $W$ are iid $\mathcal{U}(0,1)$ variables, then it can also be shown that $\mathcal{V}^{\leftarrow}(U,W)\sim\mathcal{U}(0,1)$.
+Note, however, that while $\mathcal{V}\left(\mathcal{V}^{\leftarrow}(y,W)\right)=y$ holds for all $y\in[0,1]$, the equation $\mathcal{V}^{\leftarrow}\left(\mathcal{V}(u),W\right)=u$ is only true with probability $\Delta(u)$, if $u<\delta$, and with probability $1-\Delta(u)$, if $u>\delta$.
 3.2 V-transformation and inverse v-transformation of copulas
 The following result, which is an adaptation of a result in McNeil, (2021), is key to constructing new copulas using v-transforms and their inverses.
+Part (i) is related to the result of Loaiza-Maya et al., (2018) but is expressed in a manner that involves only copulas and is free of marginal distributions.
 Theorem 1.
-Let $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ be v-transforms with associated partial inverses $\mathcal{V}_{1}^{-1}$ and $\mathcal{V}_{2}^{-1}$ and $\Delta$-functions $\Delta_{1}$ and $\Delta_{2}$.
-Let $(U,V)^{\top}$ be distributed according to a copula $C_{U,V}$ with density $c_{U,V}$.
+Let $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ be v-transforms with associated partial inverses $\mathcal{V}_{1}^{-1}$ and $\mathcal{V}_{2}^{-1}$, $\Delta$-functions $\Delta_{1}$ and $\Delta_{2}$ and stochastic inverses $\mathcal{V}_{1}^{\leftarrow}$ and $\mathcal{V}_{2}^{\leftarrow}$.
+Let $(U,V)^{\top}$ be distributed according to a copula $C$ with density $c$ and let $W_{1}$ and $W_{2}$ be iid $\mathcal{U}(0,1)$ variables, independent of $(U,V)^{\top}$.
   1. (i)
-    $(\mathcal{V}_{1}(U),\mathcal{V}_{2}(V))^{\top}$ is distributed according to a copula $C_{\mathcal{V}_{1}(U),\mathcal{V}_{2}(V)}$ with density
-    $c_{\mathcal{V}_{1}(U),\mathcal{V}_{2}(V)}(u,v)=\;$ $c_{U,V}(\mathcal{V}_{1}^{-1}(u),\mathcal{V}_{2}^{-1}(v))\;\Delta_{1}(u)\,\Delta_{2}(v)\;+$
-    $c_{U,V}(\mathcal{V}_{1}^{-1}(u)+u,\mathcal{V}_{2}^{-1}(v))\;\big{(}1-\Delta_{1}(u)\big{)}\,\Delta_{2}(v)\;+$ (14)
-    $c_{U,V}(\mathcal{V}_{1}^{-1}(u),\mathcal{V}_{2}^{-1}(v)+v)\;\Delta_{1}(u)\,\big{(}1-\Delta_{2}(v)\big{)}\;+$
-    $c_{U,V}(\mathcal{V}_{1}^{-1}(u)+u,\mathcal{V}_{2}^{-1}(v)+v)\;\big{(}1-\Delta_{1}(u)\big{)}\,\big{(}1-\Delta_{2}(v)\big{)}.$
+    $(\mathcal{V}_{1}(U),\mathcal{V}_{2}(V))^{\top}$ is distributed according to a copula $C^{*}$ with density
+    $c^{*}(u,v)=c(\mathcal{V}_{1}^{-1}(u),\mathcal{V}_{2}^{-1}(v))\;\Delta_{1}(u)\,\Delta_{2}(v)\;+c(\mathcal{V}_{1}^{-1}(u)+u,\mathcal{V}_{2}^{-1}(v))\;\big(1-\Delta_{1}(u)\big)\,\Delta_{2}(v)\;+$
+    $c(\mathcal{V}_{1}^{-1}(u),\mathcal{V}_{2}^{-1}(v)+v)\;\Delta_{1}(u)\,\big(1-\Delta_{2}(v)\big)+c(\mathcal{V}_{1}^{-1}(u)+u,\mathcal{V}_{2}^{-1}(v)+v)\;\big(1-\Delta_{1}(u)\big)\,\big(1-\Delta_{2}(v)\big).$ (13)
   2. (ii)
-    $(\mathcal{V}_{1}^{\leftarrow}(U),\mathcal{V}_{2}^{\leftarrow}(V))^{\top}$ is distributed according to a copula $C_{\mathcal{V}^{\leftarrow}_{1}(U),\mathcal{V}^{\leftarrow}_{2}(V)}$ with density
-    $c_{\mathcal{V}_{1}^{\leftarrow}(U),\mathcal{V}_{2}^{\leftarrow}(V)}(u,v)=c_{U,V}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v)).$ (15)
+    $(\mathcal{V}_{1}^{\leftarrow}(U,W_{1}),\mathcal{V}_{2}^{\leftarrow}(V,W_{2}))^{\top}$ is distributed according to a copula $C^{\dagger}$ with density
+    $c^{\dagger}(u,v)=c(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v)).$ (14)
 Proof.
 This may be proved by a simple adaptation of the proof of Theorem 3 in McNeil, (2021) which deals with the case of two (or more) identical v-transforms. ∎
-We refer to the copula density $c_{\mathcal{V}_{1}(U),\mathcal{V}_{2}(V)}$ in part (i) as a v-transformed version of $c_{U,V}$ while we refer to the copula density $c_{\mathcal{V}_{1}^{\leftarrow}(U),\mathcal{V}_{2}^{\leftarrow}(V)}$ in part (ii) as an inverse-v-transformed version of $c_{U,V}$.
-Inverse v-transformation of a copula density using (15) followed by v-transformation using (14) returns the original copula density $c_{U,V}$ (since for any v-transform $\mathcal{V}\circ\mathcal{V}^{\leftarrow}(y)=y$ for $y\in[0,1]$).
-However v-transformation of a copula density using (14) followed by inverse v-transformation using (15) does not, in general, return the original density $c$, except in special cases, such as the following.
+Part (i) refers to what we call (componentwise) v-transformation of $(U,V)^{\top}$, or its distribution $C$, while part (ii) refers to what we call (componentwise independent) inverse v-transformation of $(U,V)^{\top}$, or $C$.
+Inverse v-transformation of $(U,V)^{\top}$ followed by v-transformation returns the original random vector $(U,V)^{\top}$ (since $\mathcal{V}(\mathcal{V}^{\leftarrow}(y,W))=y$ for $y\in[0,1]$ and any v-transform).
+However v-transformation of $(U,V)^{\top}$ followed by inverse v-transformation returns a different random vector with, in general, a different distribution.
+An exception is the following special case.
 Corollary 2.
-If $C_{U,V}$ is a jointly symmetric copula and $\mathcal{V}_{1}(u)=\mathcal{V}_{2}(u)=|2u-1|$ for $u\in[0,1]$, then the density of the copula $C_{\mathcal{V}_{1}(U),\mathcal{V}_{2}(V)}$ satisfies
-$c_{\mathcal{V}_{1}(U),\mathcal{V}_{2}(V)}(u,v)$ $=$ $c_{U,V}\left(\frac{1-u}{2},\frac{1-v}{2}\right)=c_{U,V}\left(\frac{1-u}{2},\frac{1+v}{2}\right)$ (16)
-$=$ $c_{U,V}\left(\frac{1+u}{2},\frac{1-v}{2}\right)=c_{U,V}\left(\frac{1+u}{2},\frac{1+v}{2}\right)$
-and
-$c_{\mathcal{V}_{1}^{\leftarrow}\circ\mathcal{V}_{1}(U),\mathcal{V}_{2}^{\leftarrow}\circ\mathcal{V}_{2}(V)}(u,v)=c_{\mathcal{V}_{1}(U),\mathcal{V}_{2}(V)}(|2u-1|,|2v-1|)=c_{U,V}(u,v).$ (17)
+Under the conditions of Theorem 1, if $C$ is jointly symmetric and $\mathcal{V}_{1}(u)=\mathcal{V}_{2}(u)=|2u-1|$ for $u\in[0,1]$, then $(\mathcal{V}_{1}^{\leftarrow}(\mathcal{V}_{1}(U),W_{1}),\mathcal{V}_{2}^{\leftarrow}(\mathcal{V}_{2}(V),W_{2}))^{\top}\stackrel{{\scriptstyle\text{\tiny d}}}{{=}}(U,V)^{\top}$.
 Proof.
-By part (i) of Theorem 1 and equation (14), the copula density satisfies
-$c_{\mathcal{V}_{1}(U),\mathcal{V}_{2}(V)}(u,v)$ $=\frac{1}{4}\left(c_{U,V}\left(\frac{1-u}{2},\frac{1-v}{2}\right)+c_{U,V}\left(\frac{1-u}{2},\frac{1+v}{2}\right)\right.$
-$\left.+c_{U,V}\left(\frac{1+u}{2},\frac{1-v}{2}\right)+c_{U,V}\left(\frac{1+u}{2},\frac{1+v}{2}\right)\right)$
-and (16) follows easily from the joint symmetry property of $C_{U,V}$.
-The identity (17) is then easily verified by considering the values of $|2u-1|$ and $|2v-1|$ in the four quadrants of the unit square defined by the lines $u=0.5$ and $v=0.5$ and using each of the four equations in (16) in turn.
-∎
-Remark 2.
-We can easily derive a corresponding result for the simpler case where the copula $C_{U,V}$ has h-symmetry and $V$ is v-transformed with the symmetric v-transform.
-The copula density of the random vector $(U,|2V-1|))^{\top}$ can be shown to satisfy
-$c_{U,|2V-1|}(u,v)=c_{U,V}\left(u,\frac{1-v}{2}\right)=c_{U,V}\left(u,\frac{1+v}{2}\right)$ (18)
-and it is easily verified that $c_{U,|2V-1|}(u,|2v-1|)=c_{U,V}(u,v)$.
-Remark 3.
-To see that the identity (17) does not always hold when $\mathcal{V}_{1}(u)=\mathcal{V}_{2}(u)=|2u-1|$, it suffices to start with any copula $C_{U,V}$ that is not jointly symmetric, since application of the inverse v-transform in the final step always results in a jointly symmetric copula.
-For example, when $C_{U,V}=C^{\text{Ga}}_{\rho}$, the Gaussian copula with parameter $\rho>0$, v-transformation followed by inverse v-transformation results in the jointly symmetric mixture $0.5C^{\text{Ga}}_{\rho}+0.5C^{\text{Ga}}_{-\rho}$.
-Corollary 1 and Remark 2 allow us to state the following simple result, which is relevant to our study of the symmetries in the densities of copulas of GARCH(1,1) type.
-Proposition 6.
-Let $C$ be an absolutely continuous copula with density $c$.
-  1. (i)
-    $C$ is h-symmetric if and only if $c(u,v)=c^{*}(u,|2v-1|)$ for some copula density $c^{*}$.
-  2. (ii)
-    $C$ is jointly symmetric if and only if $c(u,v)=c^{*}(|2u-1|,|2v-1|)$ for some copula density $c^{*}$.
-Proof.
-It is trivial to verify that copulas with the structures $c^{*}(u,|2v-1|)$ and $c^{*}(|2u-1|,|2v-1|)$ are h-symmetric and jointly symmetric, respectively.
-To show the reverse implication in part (i), we choose $c^{*}$ to be the copula density of $(U,|2V-1|)^{\top}$ where $(U,V)^{\top}$ is a random vector with copula density $c$.
-To show the reverse implication in part (ii), we choose $c^{*}$ to be the copula density of $(|2U-1|,|2V-1|)^{\top}$ where $(U,V)^{\top}$ is a random vector with copula density $c$.
-∎
+By part (ii) of Corollary 1 the density $c$ of the copula $C$ of $(U,V)^{\top}$ can be written as $c(u,v)=c^{*}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v))$ where $c^{*}$ is the density of the copula of $(\mathcal{V}_{1}(U),\mathcal{V}_{2}(V))^{\top}$.
+By part (ii) of Theorem 1 $c^{*}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v))$ is also the density of the inverse v-transformation of $(\mathcal{V}_{1}(U),\mathcal{V}_{2}(V))^{\top}$ with the stochastic inverses $\mathcal{V}_{1}^{\leftarrow}$ and $\mathcal{V}_{2}^{\leftarrow}$.
 3.3 Implicit v-transforms in general bivariate densities
-In this section we use part (i) of Theorem 1 to find a condition for our generic random vector $(Y,Z)^{\top}$ on $\mathbb{R}^{2}$ which allows us to write its copula density $c$ in the general form $c(u,v)=c^{*}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v))$ for v-transforms $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ and a copula density $c^{*}$.
-This result requires the concept of a v-shaped transformation on $\mathbb{R}$, which should be distinguished from a v-transform.
+In this section we use part (i) of Theorem 1 to find a condition that allows us to write the copula density $c$ of a random vector $(Y,Z)^{\top}$ on $\mathbb{R}^{2}$ in the form $c(u,v)=c^{*}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v))$ for v-transforms $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ and a copula density $c^{*}$.
+This result requires the concept of a v-shaped transformation on $\mathbb{R}$, which generalizes the idea found in Loaiza-Maya et al., (2018) to also allow non-symmetric transformations.
 Definition 4.
-A function $T$ is a v-shaped function on $\mathbb{R}$ if it can be written in the form
-$T(x)=\begin{cases}T_{L}(\mu_{T}-x)&\quad x\leqslant\mu_{T}\\ T_{R}(x-\mu_{T})&\quad x>\mu_{T}\end{cases}$
-where $T_{L}$ and $T_{R}$ are strictly increasing, continuous and differentiable functions on $\mathbb{R}^{+}=[0,\infty)$ such that $T_{L}(0)=T_{R}(0)$ and $\mu_{T}\in\mathbb{R}$.
+A function $T$ is a v-shaped function on $\mathbb{R}$ if it can be written in the form $T(x)=T_{L}(\mu_{T}-x)$ for $x\leqslant\mu_{T}$ and $T(X)=T_{R}(x-\mu_{T})$ for $x>\mu_{T}$ where $T_{L}$ and $T_{R}$ are strictly increasing, continuous and differentiable functions on $\mathbb{R}^{+}=[0,\infty)$ such that $T_{L}(0)=T_{R}(0)$ and $\mu_{T}\in\mathbb{R}$.
 Theorem 2.
@@
 If the density ratio $f_{Y,Z}(y,z)/(f_{Y}(y)f_{Z}(z))$ can be written in the form
-$\frac{f_{Y,Z}(y,z)}{f_{Y}(y)f_{Z}(z)}=g\left(T_{1}(y),T_{2}(z)\right)$
-for two v-shaped transformations $T_{1}$ and $T_{2}$ on $\mathbb{R}$ and some function $g:\mathbb{R}\times\mathbb{R}\to\mathbb{R}^{+}$, then we may write $c(u,v)=c^{*}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v))$ on $[0,1]^{2}$ where $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ are v-transforms given by $\mathcal{V}_{1}(u)=F_{T_{1}(Y)}\circ T_{1}\circ F_{Y}^{-1}(u)$ and $\mathcal{V}_{2}(v)=F_{T_{2}(Z)}\circ T_{2}\circ F_{Z}^{-1}(v)$ and where $c^{*}$ is the density of the copula of $(T_{1}(Y),T_{2}(Z))^{\top}$.
+$\frac{f_{Y,Z}(y,z)}{f_{Y}(y)f_{Z}(z)}=g\left(T_{1}(y),T_{2}(z)\right)$ (15)
+for two v-shaped transformations $T_{1}$ and $T_{2}$ on $\mathbb{R}$ and some function $g:\mathbb{R}\times\mathbb{R}\to\mathbb{R}^{+}$, then we may write $c(u,v)=c^{*}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v))$ on $0,1]^{2}$ where $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ are v-transforms given by $\mathcal{V}_{1}(u)=F_{T_{1}(Y)}\circ T_{1}\circ F_{Y}^{-1}(u)$ and $\mathcal{V}_{2}(v)=F_{T_{2}(Z)}\circ T_{2}\circ F_{Z}^{-1}(v)$ and where $c^{*}$ is the density of the copula of $(T_{1}(Y),T_{2}(Z))^{\top}$.
 Proof.
@@
 ∎
-We give two examples to illustrate the use of Theorem 2.
-Example 1 (Symmetric GARCH-type process).
-Theorem 2 gives an alternative proof of the fact that the copula density $c_{1}$ in a symmetric GARCH-type process can be written in the form $c^{*}(|2u-1|,|2v-1|)$ where $c^{*}$ is the density of the copula of $(|X_{1}|,|X_{2}|)^{\top}$.
-The joint density in (6) may be written as
-$f_{X_{1},X_{2}}(x,y)=\int_{0}^{\infty}\frac{1}{w\sigma\left(|x|,s\right)}f_{\epsilon}\left(\frac{|y|}{\sigma\left(|x|,s\right)}\right)f_{\epsilon}\left(\frac{|x|}{s}\right)f_{\sigma}(s)\mathrm{d}s$
-in terms of the absolute value function, which is a v-shaped transformation on $\mathbb{R}$; moreover the marginal distribution satisfies $f_{X}(x)=f_{X}(|x|)$.
+Symmetric models of GARCH(1,1)-type provide examples of models where the density ratio for $f_{X_{1},X_{2}}$ (and also for $f_{X_{1},X_{k+1}\mid\bm{X}_{[2:k]}}$) may be factorized in the form (15) by taking $T_{1}(x)=T_{2}(X)=|x|$.
+Unfortunately, when innovations distributions are asymmetric, or when leverage is introduced into a model that is not of ARCH(1) type effect, perfect factorization of the form (15) does not seem to be possible.
+However, it is possible in the following case.
 Example 1 (ARCH(1) process with symmetric innovations and leverage).
@@
 To derive $\mathcal{V}_{1}$ we must find the composition of functions $\mathcal{V}_{1}(u)=F_{\sigma}\circ\sigma\circ F_{X}^{-1}(u)$.
-By using the fact that the distribution function of $\sigma_{t}$ is
-$F_{\sigma}(s)=F_{X}\left(\sqrt{\frac{s^{2}-\alpha_{0}}{\alpha_{1}}}\right)-F_{X}\left(-\sqrt{\frac{s^{2}-\alpha_{0}}{\alpha_{1}+\gamma_{1}}}\right),$
-we obtain after some calculation that
+By using the fact that the distribution function of $\sigma_{t}$ is $F_{\sigma}(s)=F_{X}\left(\sqrt{(s^{2}-\alpha_{0})/\alpha_{1}}\right)-F_{X}\left(-\sqrt{(s^{2}-\alpha_{0})(\alpha_{1}+\gamma_{1})}\right)$ we obtain after some calculation that
 $\mathcal{V}_{1}(u)=\begin{cases}F_{X}\left(-\left(1+\frac{\gamma_{1}}{\alpha_{1}}\right)^{\frac{1}{2}}F_{X}^{-1}(u)\right)-u,&u\leqslant 0.5,\\ u-F_{X}\left(-\left(1+\frac{\gamma_{1}}{\alpha_{1}}\right)^{-\frac{1}{2}}F_{X}^{-1}(u)\right),&u>0.5,\end{cases}$
 which is an asymmetric function for $\gamma_{1}\neq 0$.
-For the classical ARCH(1) process with asymmetric innovation density in the upper right panel of Figure 3 it does not seem to be generally possible to write the copula density in the form $c_{1}(u,v)=c_{1}^{*}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v))$.
-Although the density ratio can be written in the form
-$\frac{f_{X_{1},X_{2}}(x,y)}{f_{X}(x)f_{X}(y)}=\frac{f_{\epsilon}\left(\frac{y}{\sigma(x)}\right)}{\sigma(x)f_{X}(y)}=h(\sigma(x),y),$
-it does not appear to be possible to find a second v-shaped transformation $T_{2}$ such that $h(\sigma(x),y)=g(\sigma(x),T_{2}(y))$.
-This conclusion extends to the more general models of first-order GARCH type with asymmetric innovation distributions.
-Despite the negative conclusion, copula densities of the form $c(u,v)=c^{*}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v))$ are still promising candidates for approximating or mimicking the implicit copulas of GARCH(1,1)-type models and this is the subject of the next section.
-5 GARCH-mimicking copulas
-5.1 Mimicking jointly symmetric GARCH copulas
-There are a number of ways of finding jointly symmetric copulas that might be capable of mimicking the jointly symmetric copula of a symmetric GARCH process. We could consider:
-  1. 1.
+3.4 Approaches to approximating GARCH copulas
+Despite the limited number of GARCH(1,1)-type models for which we can prove that pair copulas must take the exact form $c(u,v)=c^{*}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v))$, this construction remains of interest as a flexible means of approximating the cross-like behaviour of the implied GARCH pair copulas.
+Our approach to this problem is to use simulation studies to first identify copula densities of the form $c^{*}(|2u-1|,|2v-1|)$ to approximate the jointly symmetric copulas of symmetric GARCH processes and then to consider how asymmetric v-transforms $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ can accommodate the asymmetries introduced by leverage and asymmetric innovations.
+There are two ways of identifying appropriate copula densities $c^{*}$ for building models with densities of the form $c^{*}(|2u-1|,|2v-1|)$: one is the explicit method using part (ii) of Theorem 1 where we simply choose a candidate for the copula $C^{*}$ based on the insights we have collected so far; the other is the implicit method where we consider constructions that are known to yield a jointly symmetric copula and then infer the form of $c^{*}$.
+Explicit choices of $C^{*}$ might be the Gumbel, Joe or 180-degree-rotated (survival) Clayton copulas.
+These all have upper tail dependence and may be capable of approximating the copulas of the absolute value processes of symmetric GARCH(1,1)-type models.
+Recall that the densities of such copulas can be thought of as blown-up versions of the upper-right quadrants of the copula densities in Figure 1.
+We impose one technical condition on copulas $C^{*}$ which facilitates maximum likelihood inference for models with densities of the form $c^{*}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v))$: we require that $c^{*}(0,0)<\infty$ which ensures that the log-likelihood is properly defined for observations at the fulcrum values.
+As examples of the implicit construction we could consider:
+  1. (1)
     the copula of a spherically symmetric distribution;
-  2. 2.
+  2. (2)
     an equal (50/50) mixture of a radially symmetric copula and its 90-degree rotation;
-  3. 3.
-    an equal (25/25/25/25) mixture of any copula and its rotations through 90, 180 and 270 degrees;
-  4. 4.
-    a copula constructed from another copula $C^{*}$ by inverse v-transformation with the symmetric v-transform, i.e. a copula with density of the form $c^{*}(|2u-1|,|2v-1|)$.
-In view of Proposition 6, the densities of the copulas arising from the first 3 approaches can also be written in the form $c(u,v)=c^{*}(|2u-1|,|2v-1|)$ for some implicit copula $C^{*}$ with density $c^{*}$, so the 4th approach should be considered as the overarching approach.
-Proposition 3, part (ii), and the lower-left panel of Figure 3 give clues about the kind of copula density $c^{*}$ we should be targeting.
-In particular, we note that there should be strong dependence in the upper right tail and some asymmetry.
+  3. (3)
+    an equal (25/25/25/25) mixture of any copula and its rotations through 90, 180 and 270 degrees.
 An example of (1) is the copula of a spherical t distribution, i.e.
@@
 As the degree-of-freedom parameter tends to zero, this copula becomes increasingly cross-shaped and converges to an equal mixture of the comonotonicity and countermonotonicity copulas (McNeil et al.,, 2022).
-It is of interest to identify the copula corresponding to v-transformation of $C^{t}_{\nu}$, in other words, the copula of $(|Y|,|Z|)^{\top}$.
-We call this copula the absolute spherical t copula (ast) and denote it $C^{|t|}_{\nu}$.
-By Corollary 1 and (16) its density must satisfy
+We denote the copula of $(|Y|,|Z|)^{\top}$ by $C^{|t|}_{\nu}$ and call it the absolute spherical t copula (ast).
+By (4) its density must satisfy
 $c^{|t|}_{\nu}(u,v)=c^{t}_{\nu}\left(\frac{1-u}{2},\frac{1-v}{2}\right)=c^{t}_{\nu}\left(\frac{1-u}{2},\frac{1+v}{2}\right)=c^{t}_{\nu}\left(\frac{1+u}{2},\frac{1-v}{2}\right)=c^{t}_{\nu}\left(\frac{1+u}{2},\frac{1+v}{2}\right).$
-Integrating any one of these equalities and observing the boundary conditions for a copula gives us an expression for $C^{|t|}_{\nu}$. For example, integrating the final equality gives
-$C^{|t|}_{\nu}=4C_{\nu}^{t}\left(\frac{1+u}{2},\frac{1+v}{2}\right)-u-v-1.$ (20)
 This one-parameter copula, which appears to be new, interpolates between independence (as $\nu\to\infty$) and comonotonicity (as $\nu\to 0$).
-For small degrees of freedom, the density $c^{|t|}_{\nu}$ does bear some resemblance to the lower-left panel of Figure 3, but cannot reproduce the non-exchangeability.
-An example of (2) would be an equal mixture of the t copula $C^{t}_{\nu,\rho}$ and the t copula $C^{t}_{\nu,-\rho}$ for some $\rho\geqslant 0$; obviously this contains the spherical t copula $C^{t}_{\nu}$ as the special case where $\rho=0$.
-Examples of (3) could be constructed using the Gumbel, Clayton or Joe copulas, to take advantage of the strong tail dependence of these copulas in either the upper-right or lower-left corners; this includes the equally-weighted special case of the so-called convex Gumbel copula (Junker and May,, 2005) used by Loaiza-Maya et al., (2018) for financial return data.
-Mixtures of types (2) and (3) all imply v-transformed copulas $C^{*}$, whose forms may be easily derived using the method we have used for the spherical t copula.
-If we follow the explicit construction (4), we need to choose a copula $C^{*}$ that has stronger dependence in the upper tail than the lower tail.
-This suggests considering the Gumbel, Joe or 180-degree-rotated (survival) Clayton copulas.
+It has upper tail dependence like the Gumbel, Joe and survival Clayton copulas.
+More information is collected in Section S4 in the Supplementary Material.
+An example of (2) is an equal mixture of the t copulas $C^{t}_{\nu,\rho}$ and $C^{t}_{\nu,-\rho}$ for some $\rho\geqslant 0$, which is the jointly symmetric special case of a copula considered by Loaiza-Maya et al., (2018); the t mixture reduces to the spherical t copula $C^{t}_{\nu}$ when $\rho=0$.
+Construction (3), which is the bivariate version of a higher-dimensional family used in Oh and Patton, (2016), could be based on the Gumbel, Clayton or Joe copulas, to take advantage of the strong tail dependence of these copulas in either the upper-right or lower-left corners.
+The Gumbel case of (3) is the jointly symmetric special case of the convex Gumbel copula proposed by Junker and May, (2005) and used by Loaiza-Maya et al., (2018) in their analysis.
 So far none of these constructions address the issue of non-exchangeability which was clearly evident in certain examples in Figure 1.
-One possible construction for non-exchangeable copulas is the one proposed by Khoudraji, (1995) and later generalized by Liebscher, (2008).
-In this construction two copulas $C_{A}$ and $C_{B}$ are combined to give a new copula
-$C_{K}(u,v)=C_{A}(u^{1-a_{1}},v^{1-a_{2}})C_{B}(u^{a_{1}},v^{a_{2}})$ (21)
-by choosing constants $a_{i}\in[0,1]$ for $i=1,2$ and the resulting copula $C_{K}$ is non-exchangeable when $a_{1}\neq a_{2}$.
-In our use of this construction we set $C_{B}$ to be the independence copula.
+A possible construction for non-exchangeable copulas was proposed by Khoudraji, (1995) and we use a version of this construction in which we generalize an exchangeable copula $C_{\theta}$ with parameter $\theta$ by setting
+$C_{\theta,a_{1},a_{2}}(u,v)=C_{\theta}(u^{1-a_{1}},v^{1-a_{2}})u^{a_{1}}v^{a_{2}}$ (17)
+for $a_{1}\in[0,1]$ and $a_{2}\in[0,1]$.
+The resulting copula is non-exchangeable when $a_{1}\neq a_{2}$.
 We apply the Khoudraji extension to copulas before either mixing them with their rotations or applying inverse v-transformation.
-We note, however, that the class of copulas in (21) does have one problematic feature for maximum-likelihood inference; even when $C_{B}$ is the independence copula, in order to evaluate the copula density of $c_{K}$ of $C_{K}$, we need to be able to evaluate, not only the copula density $c_{A}$, but also the copula $C_{A}$ and its partial derivatives.
-This is a particular challenge for copulas related to the t copula [^3] For example, existing R software for evaluating the distribution function of the t copula imposes some limitations on $\nu$.
-The algorithms in the R package copula (Hofert et al.,, 2023), which are based on methods by Gentz and Bretz (Gentz and Bretz,, 2009), are limited to integer $\nu$ and this is insufficiently fine-grained for applications of $C^{|t|}_{\nu}$.
-The C++ library underlying rvinecopulib (Nagler and Vatter,, 2022) imposes the restriction that $\nu\geqslant 2$..
-5.2 Simulation study of jointly symmetric case
-We explore the ideas of the previous section via a Monte Carlo simulation study.
-We generate data from the classical symmetric ARCH(1) and GARCH(1,1) processes using the parameter values and innovation distributions summarized in Table 1.
-We then fit copulas with the property of joint symmetry to the generated realizations of the pairs $(X_{t},X_{t+1})^{\top}$.
-In this and all subsequent simulations, the analyses are based on simulated samples of 50000 points.
-Copulas are estimated by the method of pseudo maximum likelihood (PML) (Genest and Rivest,, 1993); in other words, marginal distributions are estimated empirically, the data are mapped to the unit square using the empirical distribution function and copula parameters are then determined by maximum likelihood.
+3.5 Simulation study of the approximation of $C_{1}$
+To get some insight into the approximations that work best for the pair copulas implied by GARCH-type models, we carry out a simulation study in which we focus on estimation of the first-order copula $C_{1}$.
+Samples of 50000 values are generated from ARCH(1)-type and GARCH(1,1)-type processes using the parameter values and innovation distributions summarized in Table 1.
+The copula $C_{1}$ is estimated by fitting a first-order D-vine to the data using the two-stage method developed by Chen and Fan, (2006) in which the marginal distribution of the process is estimated by the rescaled empirical distribution function (EDF) to produce pseudo-copula data.
+The explicit and implicit constructions of Section 3.4 are tested and fitted models are compared according to their AIC values.
+Note that, the first-order copulas $C_{1}$ are the only pair copulas for which we can produce pseudo-copula data using only the rescaled EDF; we address approximations for higher-order copulas $(C_{k})_{k>1}$ indirectly in Section 4.2 by fitting higher-order simplified D-vines to simulated GARCH data.
+Under the assumption that the model is first-order Markov with known copula, Chen and Fan, (2006) show consistency and asymptotic normality of the resulting parameter estimates subject to technical conditions which include that the process should be $\beta$-mixing at a geometric rate.
+All the GARCH(1,1)-type processes we consider in our simulation satisfy this mixing condition (Francq and Zakoïan,, 2006).
+However, the first-order Markov assumption is only true for ARCH(1)-type processes and our models contain approximating copulas rather than correctly specified copulas.
+Nagler et al., (2022) supply some very general results which can be applied to semiparametric inference for simplified D-vines of higher order when the pair copulas are misspecified.
+Their results suggest that consistency and asymptotic normality of the procedure hold in the misspecified model, albeit convergence is to the parameter values that solve the model score equations, which coincide with the true values in a correctly specified model.
+3.5.1 Jointly symmetric models
+Model ARCH(1) GARCH(1,1)
+Copula (form of density) $p$ | $(\epsilon_{t})$ Gauss t4 t2.5 Gauss t4 t2.5
+$c^{t}_{\nu}(u,v)=c^{|t|}_{\nu}(|2u-1|,|2v-1|)$ 1 528 0 126 103 7 95
+$\tfrac{1}{2}\left(c^{t}_{\rho,\nu}(u,v)+c^{t}_{-\rho,\nu}(u,v)\right)$ 2 530 2 128 101 9 97
+$\tfrac{1}{4}\left(c_{\theta}^{\text{Cl}}(u,v)+c_{\theta}^{\text{Cl},90}(u,v)+c_{\theta}^{\text{Cl},180}(u,v)+c_{\theta}^{\text{Cl},270}(u,v)\right)$ 1 1676 426 189 491 291 145
+$\tfrac{1}{4}\left(c_{\theta}^{\text{Gu}}(u,v)+c_{\theta}^{\text{Gu},90}(u,v)+c_{\theta}^{\text{Gu},180}(u,v)+c_{\theta}^{\text{Gu},270}(u,v)\right)$ 1 913 150 182 228 120 156
+$\tfrac{1}{4}\left(c_{\theta}^{\text{J}}(u,v)+c_{\theta}^{\text{J},90}(u,v)+c_{\theta}^{\text{J},180}(u,v)+c_{\theta}^{\text{J},270}(u,v)\right)$ 1 1982 593 276 666 431 232
+$c_{\theta}^{\text{Cl},180}(|2u-1|,|2v-1|)$ 1 529 146 292 174 105 216
+$c_{\theta}^{\text{Gu}}(|2u-1|,|2v-1|)$ 1 1167 444 379 422 307 290
+$c_{\theta}^{\text{J}}(|2u-1|,|2v-1|)$ 1 650 50 145 144 65 131
+$\tfrac{1}{4}\left(c_{\theta,a_{1},a_{2}}^{\text{Cl}}(u,v)+c_{\theta,a_{1},a_{2}}^{\text{Cl},90}(u,v)+c_{\theta,a_{1},a_{2}}^{\text{Cl},180}(u,v)+c_{\theta,a_{1},a_{2}}^{\text{Cl},270}(u,v)\right)$ 3 1170 430 102 385 288 70
+$\tfrac{1}{4}\left(c_{\theta,a_{1},a_{2}}^{\text{Gu}}(u,v)+c_{\theta,a_{1},a_{2}}^{\text{Gu},90}(u,v)+c_{\theta,a_{1},a_{2}}^{\text{Gu},180}(u,v)+c_{\theta,a_{1},a_{2}}^{\text{Gu},270}(u,v)\right)$ 3 854 154 66 199 124 64
+$\tfrac{1}{4}\left(c_{\theta,a_{1},a_{2}}^{\text{J}}(u,v)+c_{\theta,a_{1},a_{2}}^{\text{J},90}(u,v)+c_{\theta,a_{1},a_{2}}^{\text{J},180}(u,v)+c_{\theta,a_{1},a_{2}}^{\text{J},270}(u,v)\right)$ 3 1680 597 178 632 435 151
+$c_{\theta,a_{1},a_{2}}^{\text{Cl},180}(|2u-1|,|2v-1|)$ 3 46 147 156 89 98 119
+$c_{\theta,a_{1},a_{2}}^{\text{Gu}}(|2u-1|,|2v-1|)$ 3 598 448 221 330 303 159
+$c_{\theta,a_{1},a_{2}}^{\text{J}}(|2u-1|,|2v-1|)$ 3 93 54 11 47 63 18
+$c_{\nu,a_{1},a_{2}}^{|t|}(|2u-1|,|2v-1|)$ 3 0 3 0 0 0 0
+Table 2: Comparison of approximating copulas for the first-order copula of symmetric ARCH(1) and GARCH(1,1) processes.
+Estimates are obtained by fitting first-order D-vines using semiparametric method of Chen and Fan, (2006).
+Each column shows the differences between AIC and the minimum value of AIC for that column; thus a 0 indicates the model with lowest AIC.
+The superscripts Cl, Gu and J refer to Clayton, Gumbel and Joe copulas while 90, 180 and 270 denote rotations.
+Densities $c_{\theta,a_{1},a_{2}}$ correspond to the Khoudraji construction in (17).
+The results in Table 2 relate to classical ARCH(1) and GARCH(1,1) models with symmetric innovations.
+The innovation distributions considered are Gaussian, Student t4 and Student t2.5, giving six different datasets in total.
+The spherical t copula $C^{t}_{\nu}$ works quite well in all cases (despite its exchangeability) and is the favoured model for the case of ARCH(1) with t4 innovations.
+It is notable that the spherical t copula generally outperforms the mixture of two t copulas; in the latter the estimate of $\rho$ is very close to zero suggesting the model is poorly specified for this application.
+For the other datasets the favoured copula has density of the form $c^{*}(|2u-1|,|2v-1|)$ where $C^{*}$ is the Khoudraji extension (17) of the ast copula; the second-best copula is the model where $C^{*}$ is the Khoudraji extension of either the Joe copula (four cases) or the survival Clayton copula (one case).
+Models constructed using explicit inverse v-transformation are more effective than mixture copulas constructed according to approaches (2) and (3) of Section 3.4.
+While the Gumbel copula gives the best mixtures, the Joe and survival Clayton copulas give better inverse-v-transformed models than the Gumbel copula.
+3.5.2 Models with asymmetries
+When models contain asymmetric innovations, and potentially also leverage, we consider approximating GARCH pair copulas using the construction $c^{*}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v))$ with asymmetric v-transforms $\mathcal{V}(u;\delta,\kappa)$ taken from the family (12), which includes the linear special case $\mathcal{V}(u;\delta)$.
+Section S5.2 of the Supplementary Material contains extensive results for ARCH(1) and GARCH(1,1) models with skewed t innovations, with and without leverage effects.
+The most important observation in these analyses is that, when two different asymmetric v-transforms $\mathcal{V}_{1}(u)$ and $\mathcal{V}_{2}(v)$ are used, the Khoudraji extension of the copula $C^{*}$ adds very little.
+This is a helpful insight from a computational point of view because evaluating the density of the Khoudraji extension requires, not only the density $c^{*}$ of the parent copula, but also the function $C^{*}$ itself, leading to much slower evaluation (particularly in the ast case).
+The best models are of the two-parameter form $c^{*}(\mathcal{V}(u;\delta_{1},\kappa_{1}),\mathcal{V}(v;\delta_{2},\kappa_{2}))$ with $c^{*}$ taken to be $c^{|t|}_{\nu}$ or $c^{\text{J}}_{\theta}$.
+However models of the form $c^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))$ with two asymmetric linear v-transforms already offer a large improvement on jointly symmetric models and offer the advantages of greater parsimony and greater tractability, as we discuss in Section 4.1.
+Moreover, the analyses of Section 4.2 will show that, in the presence of asymmetry, models with two asymmetric linear v-transforms tend to give better estimates of $C_{1}$ than mixtures of t copulas or convex Gumbel copulas.
+4 Simplified D-vines for stochastic volatility
+4.1 Proposal for a simplified D-vine structure
+In this section we use insights from Sections 2 and 3 to propose a class of simplified D-vines that behave like GARCH(1,1) processes and offer potential alternative models for financial time series.
+A guiding principle in constructing our models is parsimony.
+GARCH(1,1) models have relatively few parameters and an effective model that incorporates both leverage and skewed innovations can be created with six parameters ($\alpha_{0}$, $\alpha_{1}$, $\beta_{1}$, $\gamma_{1}$, $\xi$, $\lambda$).
+The first four control the volatility equation and the final two describe the shape and skew of the innovation distribution; a seventh parameter $\mu$ could be added for a non-zero mean.
+With a handful of parameters, the GARCH(1,1) model is able to model the persistence of volatility observed in financial data in a way that low-order ARCH($p$) models are not.
+To achieve a similar effect in D-vine models, we construct sequences of pair copulas $(C_{k})_{k\in\mathbb{N}}$ using three main ideas.
+  1. 1.
+    We choose pair copula densities of the form $c_{k}(u,v)=c^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2});\theta_{k})$, based on two linear v-transforms and a copula family $C^{*}(u,v;\theta)$ satisfying $c^{*}(0,0;\theta)<\infty$ which can model the full range of positive dependence between independence and comonotonicity for different values of the parameter $\theta$ in a parameter space $\Theta$; examples are the ast, Joe and survival Clayton copulas.
+This construction offers flexibility at the cost of only three parameters: $\theta_{k}$ controls the strength of dependence while $\delta_{1}$ and $\delta_{2}$ accommodate possible asymmetries.
+  2. 2.
+    We model the parameters $(\theta_{k})_{k\in\mathbb{N}}$ as a parametric function of the lag $k$ so that $\theta_{k}=g(k,\bm{\vartheta})$ where $\bm{\vartheta}$ is a low-dimensional parameter vector.
+More details of this parameterisation are given below.
+  3. 3.
+    We assume that the asymmetry parameters $\delta_{1}$ and $\delta_{2}$ are constant for all lags $k$.
+This strong assumption is based on the observation that these parameters are equal to 0.5 at all lags in jointly symmetric models.
+It is also partly influenced by the GARCH(1,1) model where a single parameter $\gamma_{1}$ controls leverage and a further parameter $\lambda$ controls asymmetry of innovations.
+There is another practical reason for choosing linear v-transforms. After making the simplifying assumption, the copula density implied by the D-vine in equation (1) takes the form
+$c_{U_{1},\ldots,U_{n}}(u_{1},\ldots,u_{n})=\prod_{k=1}^{n-1}\prod_{j=k+1}^{n}c_{k}(F_{U_{1}|\bm{U}_{[2:k]}}(u_{j-k}\mid\bm{u}_{[j-k+1:j-1]}),F_{U_{k+1}|\bm{U}_{[2:k]}}(u_{j}\mid\bm{u}_{[j-k+1:j-1]})).$ (18)
+It is well known that the the evaluation of the conditional distribution functions $F_{U_{1}|\bm{U}_{[2:k]}}$ and $F_{U_{k+1}|\bm{U}_{[2:k]}}$ in (18) is straightforward if we have a method of computing the so-called $h$-functions of the copulas $C_{k}$, which are the partial derivatives $h_{k,1}(u,v)=\frac{\partial}{\partial u}C_{k}(u,v)$ and $h_{k,2}(u,v)=\frac{\partial}{\partial v}C_{k}(u,v)$; see, for example, Aas et al., (2009) or Smith et al., (2010).
+This is easy in the linear case but, in the non-linear case, cumbersome numerical integration seems in general to be necessary.
+More details of the formulas for the linear case are found in Section S6.1 of the Supplementary Material while in Section S6.2 we show that, in the special case of two identical linear v-transforms, D-vines with inverse-v-transformed copulas also belong to the family of vt-D-vines proposed in Bladt and McNeil, 2022b .
+The relationship $\theta_{k}=g(k,\bm{\vartheta})$ between the parameters $\theta_{k}$ and the lag $k$ is determined implicitly using a method developed in Bladt and McNeil, 2022a and Bladt et al., (2025).
+Let $\tau^{*}:\Theta\to[0,1)$ be the invertible function that maps the parameter value of the copula $C^{*}(u,v;\theta)$ into the Kendall’s tau measure of association and write $(\tau^{*})^{-1}$ for its inverse.
+Let $(\omega_{k}(\bm{\vartheta}))_{k\in\mathbb{N}}$ denote the partial autocorrelation (pacf) of a classical Gaussian ARMA($p$,$q$) process, where $\bm{\vartheta}$ denotes the vector of $p+q$ parameters and where we assume that the parameters are constrained to ensure $\omega_{k}(\bm{\vartheta})\in[0,1)$ for all $k\in\mathbb{N}$.
+We determine the sequence of parameters $(\theta_{k})_{k\in\mathbb{N}}$ by setting $\theta_{k}=(\tau^{*})^{-1}\left(2\pi^{-1}\arcsin(\omega_{k}(\bm{\vartheta}))\right)$ for all $k$.
+This means that we equate the Kendall’s tau values of the copula sequence $(C^{*}(u,v;\theta_{k}))_{k\in\mathbb{N}}$ with the Kendall’s tau values of the sequence of pair copulas that characterise the Gaussian ARMA($p$,$q$) model.
+In practice we favour the ARMA(1,1) model based on the fact that the autocorrelation function of the squared values of the classical GARCH(1,1) process decays like an ARMA(1,1) process, as noted in Section 2.4.3.
+We also consider AR($p$) models to faciliate comparison with D-vine($p$) models in the literature; these lead to the estimation of $p$ parameters $\theta_{1},\ldots,\theta_{p}$.
+4.2 Examples
+To illustrate the potential of the D-vine models proposed in Section 4.1, we provide four examples, three using simulated data from the GARCH(1,1)-type processes documented in Table 1 and one using real daily return data for the USD/AUD exchange rate, as previously analysed in Loaiza-Maya et al., (2018).
+The simulated datasets consist of 10000 values from the GARCH(1,1) model with (i) t4 innovations, (ii) skewed t4 innovations and (iii) skewed t4 innovations and leverage.
+The exchange-rate returns consist of 3669 values from 2 January 2001 to 7 August 2015.
+Models are fitted using the generalization of the two-stage method of Chen and Fan, (2006) to higher-order D-vines.
+For the simulated data, the marginal distribution function is estimated by the empirical distribution function while, for the real data, we use the same adaptive kernel density estimator as Loaiza-Maya et al., (2018) to facilitate comparison of results.
+The maximum-likelihood fit of the parametric copula model to the pseudo-copula observations can be accomplished in a single step, rather than by using the common stepwise procedure where the sequence of $C_{1},\ldots,C_{p}$ defining a D-vine($p$) model is estimated in $p$ successive steps.
+To all datasets we fit D-vine(1) and D-vine(5) models based on t copulas (T1, T5) and convex Gumbel mixtures (B1, B5), as used in Loaiza-Maya et al., (2018).
+For the new models of Section 4.1, we use inverse-v-transformed Joe, survival Clayton and ast copulas and parameterizations based on the partial acfs of Gaussian AR(1), AR(5) and ARMA(1,1) processes.
+The copula sequence in the ARMA(1,1) case is truncated at 40 for the real data and 20 for the simulated data, since persistence of volatility is weaker for the latter.
+For the empirical data only, we add D-vine(1) and D-vine(5) models based on mixtures of two t copulas (A1, A5) to facilitate further comparison with models in Loaiza-Maya et al., (2018).
+For the simulated data, we add a D-vine(2) model with two t copulas, as suggested by Zhao et al., (2022) for GARCH approximation.
+Since the different models have very different numbers of parameters, we compare models using both AIC and BIC, where the latter makes greater adjustments for overfitting.
+For all four datasets the best model, as selected by both the AIC and BIC information criteria, is ast-ARMA(1,1) models and we show estimated parameter values and standard errors in Table 3.
+It may be noted that the estimates of $\delta_{1}$ and $\delta_{2}$ are both close to 0.5 for the symmetric GARCH model, but more offset in the other cases.
+The estimates of the AR and MA parameters ($\phi$ and $\psi$) are larger in absolute value for the real data than for the GARCH simulations, implying a much slower decay of the partial acf and greater persistence in volatility.
+We also give the parameters $\nu_{1}$ and $\nu_{2}$ of the ast copulas at the first two lags as well as the corresponding values of Kendall’s tau, $\tau_{1}$ and $\tau_{2}$.
+Dataset $\phi$ $\psi$ $\delta_{1}$ $\delta_{2}$ $\nu_{1}$ $\nu_{2}$ $\tau_{1}$ $\tau_{2}$
+GARCH(1,1), t4 innovations 0.780 -0.614 0.510 0.515 2.96 5.07 0.132 0.079
+(0.012) (0.016) (0.017) (0.016)
+GARCH(1,1), skew t4 innovations 0.786 -0.614 0.493 0.427 2.82 4.85 0.139 0.082
+(0.012) (0.017) (0.016) (0.016)
+GARCH(1,1), skew t4 + leverage 0.790 -0.607 0.644 0.434 2.60 4.56 0.150 0.088
+(0.013) (0.018) (0.015) (0.014)
+USD/AUD daily exchange-rate returns 0.982 -0.934 0.528 0.446 5.82 6.59 0.069 0.061
+(0.0029) (0.0073) (0.024) (0.023)
+Table 3: Parameter estimates and (standard errors) for D-vines based on ast copula and partial acf of ARMA(1,1) model fitted to four datasets.
+Models have four primary parameters, $\phi$ and $\psi$ from the ARMA(1,1) partial acf and $\delta_{1}$ and $\delta_{2}$ from the linear v-transforms; $\nu_{1}$ and $\nu_{2}$ are implied parameters of the ast copulas at the first two lags and $\tau_{1}$ and $\tau_{2}$ are the corresponding values of Kendall’s tau.
+The full model comparisons for the USD/AUD data are found in Table 4 and we give more commentary on these results since they are broadly representative of all four examples; corresponding tables for the three simulated datasets are given in Section S7 of the Supplementary Material.
+Comparing D-vine(1) models, the copulas based on inverse-v-transformation of the Joe and ast copulas lead to higher log-likelihoods and lower AIC and BIC values than the convex Gumbel mixture model B1.
+This suggests the fundamental shape of these copulas is slightly better than the mixture copulas for modelling $C_{1}$.
+Comparing D-vine(5) models, B5 gives the highest log-likelihood value but is a very expensive model to fit in terms of the number of parameters and the time required to maximize the likelihood [^3] Loaiza-Maya et al., (2018) fit this model using MCMC in the Bayesian paradigm and we use their posterior mean estimates as starting values in maximum likelihood estimation..
+The D-vine(5) models based on inverse-v-transformed copulas have less than a third of the parameters, are quicker to fit and lead to lower values for the information criteria.
+However, our favoured D-vine(40) models based on the partial acf of an ARMA(1,1) model lead to the largest log-likelihood values and the lowest AIC and BIC values at the cost of only 4 parameters.
+Loaiza-Maya et al., (2018) show that their favoured models (B5 and T5) lead to one-step-ahead forecasts of conditional quantiles (value-at-risk) that are more accurate than the ones produced by classical symmetric GARCH models and are not rejected by Christoffersen’s test of conditional coverage at the 99% level (Christoffersen,, 1998).
+This is also true of our favoured models (Joe-AR(5), ast-AR(5), Clayton180-ARMA(1,1) and ast-ARMA(1,1)); results are documented in Section S8 of the Supplementary Material.
+Model D-vine order No. pars Loglik AIC BIC
+T1 1 2 36.72 -69.43 -57.01
+A1 1 5 36.71 -63.42 -32.38
+B1 1 5 38.70 -67.40 -36.36
+Joe-AR(1) 1 3 41.17 -76.33 -57.71
+Clayton180-AR(1) 1 3 38.26 -70.51 -51.89
+ast-AR(1) 1 3 41.13 -76.27 -57.64
+T5 5 10 153.04 -286.08 -224.01
+A5 5 25 155.63 -261.26 -106.07
+B5 5 25 163.09 -276.17 -120.98
+Joe-AR(5) 5 7 154.13 -294.26 -250.81
+Clayton180-AR(5) 5 7 151.32 -288.63 -245.18
+ast-AR(5) 5 7 155.72 -297.45 -253.99
+Joe-ARMA(1,1) 40 4 247.21 -486.42 -461.59
+Clayton180-ARMA(1,1) 40 4 255.29 -502.59 -477.76
+ast-ARMA(1,1) 40 4 261.91 -515.83 -491.00
+Table 4: Comparison of D-vine models in Loaiza-Maya et al., (2018) and the D-vine models of Section 4 fitted to return data for the USD/AUD exchange rate for the period 2 January 2001 to 7 August 2015; the marginal distribution is estimated nonparametrically using adaptive kernel density estimation.
+T1, A1 and B1 are D-vine(1) models based, respectively, on t copulas, mixed t copulas and mixed convex Gumbel copulas as used in Loaiza-Maya et al., (2018), while T5, A5 and B5 are D-vine(5) extensions.
+The models following the construction in Section 4.1 are described by the name of the copula $C^{*}$ and the name of the AR or ARMA model underlying the parameterisation $\theta_{k}=g(\bm{\vartheta})$; the copula sequence in the ARMA(1,1) case is truncated at order 40.
+5 Conclusion
+This paper has asked three questions.
+First, what are the bivariate copulas that underlie the serial dependencies and conditional dependencies of pairs of variables $(X_{t},X_{t+k})$ in first-order GARCH(1,1) models?
+Second, since the true pair copulas and conditional copulas of GARCH-type processes are inaccessible for practical use, can they be approximated by simpler constructions?
+Third, can these ideas help us to construct simplified D-vines that behave like GARCH-type models, but offer more flexible models for real data showing stochastic volatility?
+With regard to the first question we find a number of symmetries in the dependence structures of the models with symmetric innovation distributions, even when leverage effects are introduced, but these symmetries disappear in the presence of asymmetric innovations, which are often necessary to model real-world data in practice.
+What can always be said is that the underlying bivariate copulas are cross-shaped, in the sense that there is a tendency for points to accumulate in all fours corners of the unit square, and that this tendency increases at shorter lags as the size of the ARCH effect increases and persists at longer lags as the aggregate size of the ARCH and GARCH effects increases.
+Motivated by the recognition that v-transforms are implicitly embedded in the copulas of symmetric GARCH models, we propose a construction to approximate GARCH pair copulas based on v-transforms and positive-dependence copulas with strong dependence in the upper joint tail.
+Moreover, we identify a new copula - the absolute spherical t copula - which offers a promising basis for the construction.
+With regard to the final question, we believe that the simplified D-vine models proposed in Section 4.1 offer effective descriptions of GARCH-like behaviour and have potential for modelling volatile financial data.
+Their advantages come not only from the particular choice of pair copulas, but also from a method of parameterization, using the partial acf of Gaussian ARMA processes, which gives models with a similar degree of parsimony to GARCH processes.
+In combination with the flexibility to choose an appropriate marginal distribution, these models could be extremely effective when data depart from the power-lawed marginal behaviour imposed by a GARCH process, as discussed in Bladt and McNeil, 2022b .
+However, the proposed models also need more extensive empirical testing.
+It is important to conduct experiments with a wider variety of financial return data, such as stock-price, index, commodity-price and volatility-index returns.
+This is the subject of future work.
+Another issue requiring further research is the construction of D-vines to model ARMA-GARCH behaviour.
+Real-world financial return series sometimes show serial correlation in the raw data as well as in the squared or absolute value series and this is usually modelled by combining ARMA and GARCH effects.
+In Section S9 of the Supplementary Material we consider an AR-ARCH model and suggest that mixtures of standard positive-dependence copulas and the pair copulas of this paper could be successful in addressing combined ARMA and GARCH effects.
+Supplementary Material
+Appendix S1 Correlation and rank correlation for h-symmetric distributions
+Nelsen, (2006) notes that jointly symmetric distributions have zero correlation (when finite second moments permit the measure to be defined).
+Perhaps surprisingly, h-symmetry on its own is a sufficient condition for an absence of both correlation and rank correlation; the same is true of v-symmetry.
+Proposition S1.
+If a random vector $(Y,Z)^{\top}$ has finite second moments and is either h-symmetric or v-symmetric then the Pearson correlation $\rho(Y,Z)$ is zero.
+Proof.
+We give the proof in the case of h-symmetry, the method for v-symmetry being identical.
+If $(Y,Z)^{\top}$ is h-symmetric about $b$ then $\mathbb{E}(Z)=b$ and we can write $\operatorname{cov}(Y,Z)=\mathbb{E}(YZ)-\mathbb{E}(Y)\mathbb{E}(Z)=\mathbb{E}(Y(Z-b))+\mathbb{E}(Yb)-\mathbb{E}(Y)b=\mathbb{E}(Y(Z-b))$.
+However, h-symmetry implies $\mathbb{E}(Y(Z-b))=\mathbb{E}(Y(b-Z))=-\mathbb{E}(Y(Z-b))$ implying that $\mathbb{E}(Y(Z-b))=0$ from which the result follows.
+∎
+Proposition S2.
+If a random vector $(Y,Z)^{\top}$ has continuous marginal distributions and is either h-symmetric or v-symmetric, then the Spearman rank correlation $\rho_{S}(Y,Z)$ and the Kendall rank correlation $\tau(Y,Z)$ are both zero.
+Proof.
+Again it suffices to give the proof in the case of h-symmetry.
+Let the unique copula of $(Y,Z)^{\top}$ be denoted $C$.
+For Spearman’s rank correlation we need to calculate $\rho_{S}(Y,Z)=12\int_{0}^{1}\int_{0}^{1}(C(u,v)-uv)\mathrm{d}u\mathrm{d}v$.
+From the identity C(u,v)=uC(u,1v)C(u,v)=u-C(u,1-v) we infer that
+$\int_{0}^{1}\int_{0}^{1}C(u,v)\mathrm{d}u\mathrm{d}v$ $=$ $\int_{0}^{1}\int_{0}^{1}u\mathrm{d}u\mathrm{d}v-\int_{0}^{1}\int_{0}^{1}C(u,1-v)\mathrm{d}u\mathrm{d}v=\frac{1}{2}-\int_{0}^{1}\int_{0}^{1}C(u,y)\mathrm{d}u\mathrm{d}y$
+where we have made the change of variable y=1vy=1-v. We must have $\int_{0}^{1}\int_{0}^{1}C(u,v)\mathrm{d}u\mathrm{d}v=\tfrac{1}{4}$ and then simple calculations give $\rho_{S}(Y,Z)=0$.
+In the case of Kendall’s rank correlation we need to compute $\tau(Y,Z)=4\mathbb{E}(C(U,V))-1$ where $(U,V)^{\top}$ is a pair of variables with distribution function $C$.
+The identity C(u,v)=uC(u,1v)C(u,v)=u-C(u,1-v) implies that $(U,V)^{\top}\stackrel{{\scriptstyle\text{\tiny d}}}{{=}}(U,1-V)^{\top}$; indeed $(U,V-0.5)^{\top}\stackrel{{\scriptstyle\text{\tiny d}}}{{=}}(U,0.5-V)^{\top}$ since $(U,V)^{\top}$ is h-symmetric about 0.5.
+Hence we infer that
+$\mathbb{E}(C(U,V))$ $=$ $\mathbb{E}(U)-\mathbb{E}(C(U,1-V))=\frac{1}{2}-\mathbb{E}(C(U,V))$
+which implies that $\mathbb{E}(C(U,V))=\tfrac{1}{4}$ and $\tau(Y,Z)=0$. ∎
+Appendix S2 More illustrations of joint densities and copulas
+Figure S1: Contour plots of $f_{X_{1},X_{2}}(x_{1},x_{2})$ for symmetric GARCH(1,1)-type processes; see Table 1 for all parameter values.
+Top left, top right, bottom left: classical ARCH(1) processes with, respectively, Gaussian, Student t4 and Student t2.5 innovations.
+Bottom right: classical GARCH(1,1) process with Gaussian innovations.
+Figure S1 shows contour plots of $f_{X_{1},X_{2}}(x_{1},x_{2})$ for a number of symmetric GARCH(1,1)-type processes as described in Sections 2.4.12.4.3; these pictures are counterparts to the copula pictures in Figure 1 and show the joint symmetry of the distributions.
+It is notable that the contours in the top right picture for an ARCH(1) model with Student t4 innovations look almost circular, showing the distribution is very close to spherical; the corresponding copula density plot in Figure 1 suggests the copula is almost exchangeable.
+In the left panel of Figure S2 we show a contour plot of the absolute value copula (the copula of $(|X_{1}|,|X_{2}|)^{\top}$) for the ARCH(1) model with Gaussian innovations described in Section 2.4.1.
+This can be thought of as a blown-up version of the top right quadrant of the copula density for $(X_{1},X_{2})^{\top}$ and clearly shows non-exchangeability and stronger dependence in the upper tail.
+Note that this picture also illustrates the density of the volatility copula of ARCH(1), that is the copula of $(\sigma_{1},\sigma_{2})^{\top}$; this follows because $\sigma_{t}$ is a strictly increasing function of $|X_{t-1}|$ for all $t$ and copulas are invariant under strictly increasing transformations of variables.
+For the same reason, it is also the copula density for the variables $(X_{1}^{2},X_{2}^{2})^{\top}$.
+Figure S2: Left: copula density for copula of the absolute values for classical ARCH(1) copula with Gaussian innovations; see Section 2.4.1 for parameters.
+Right: AR(1)-ARCH(1) model with Gaussian innovations and parameters $\phi=0.3$, $\alpha_{0}=0.4$ and $\alpha_{1}=0.6$.
+Appendix S3 General proof of Proposition 3
+We complete the proof of Proposition 3.
+The case $k=2$ has been proved so let $k\geqslant 3$.
+Suppose we write $s_{2}(x,s)=\sigma(x,s)$, $s_{3}(z_{1},x,s)=\sigma(z_{1},s_{2}(x,s))$ and then for $j=4,\ldots,k+1$ we define recursive functions $s_{j}(z_{j-2},\ldots,z_{1},x,s)=\sigma(z_{j-2},s_{j-1}(z_{j-3},\ldots,z_{1},x,s))$.
+This permits us to write the joint density as
+$f_{X_{1},\ldots,X_{k+1}}(x,z_{1},\ldots,z_{k-1},y)=\int_{0}^{\infty}f_{X_{k+1}\mid\sigma_{k+1}}(y\mid s_{k+1}(z_{k-1},\ldots,z_{1},x,s))\\ \left(\prod_{j=3}^{k}f_{X_{j}\mid\sigma_{j}}(z_{j-1}\mid s_{j}(z_{j-2},\ldots,z_{1},x,s))\right)f_{X_{2}\mid\sigma_{2}}(z_{1}\mid s_{2}(x,s))f_{X_{1}\mid\sigma_{1}}(x\mid s)f_{\sigma}(s)\mathrm{d}s.$
+This satisfies $f_{X_{1},\ldots,X_{k+1}}(x,\ldots,y)=f_{X_{1},\ldots,X_{k+1}}(x,\ldots,-y)$ when the innovation distribution is symmetric.
+Moreover, in the symmetric model, we have $f_{X_{1},\ldots,X_{k+1}}(x,\ldots,y)=f_{X_{1},\ldots,X_{k+1}}(-x,\ldots,y)$ since all of the functions $s_{2},\ldots,s_{k+1}$ are even functions of $x$.
+These properties also hold for the conditional density
+$f_{X_{1},X_{k+1}\mid\bm{X}_{[2:k]}}(x,y\mid z_{1},\ldots,z_{k-1})=f_{X_{1},\ldots,X_{k+1}}(x,z_{1},\ldots,z_{k-1},y)/f_{X_{2},\ldots,X_{k}}(z_{1},\ldots,z_{k-1}).$
+To infer that the assertions regarding symmetry of the conditional copula hold we note that h-symmetry of $f_{X_{1},X_{k+1}\mid\bm{X}_{[2:k]}}$ implies that $F_{X_{k+1}\mid\bm{X}_{[2:k]}}$ is symmetric and joint symmetry of $f_{X_{1},X_{k+1}\mid\bm{X}_{[2:k]}}$ also implies that $F_{X_{1}\mid\bm{X}_{[2:k]}}$ is symmetric; the result then follows easily from equation (10).
+Since the joint density of $(X_{1},X_{k+1})^{\top}$ can be obtained as
+$f_{X_{1},X_{k+1}}(x,y)=\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}f_{X_{1},\ldots,X_{k+1}}(x,z_{1},\ldots,z_{k-1},y)\mathrm{d}z_{1}\cdots\mathrm{d}z_{k-1}$
+it follows by similar logic and application of Lemma 1 that the copula $C_{(k)}$ of the unconditional distribution of $(X_{1},X_{k+1})^{\top}$ shares the same symmetry properties as the conditional copula $C_{k}$.
+Appendix S4 The absolute spherical t copula
+The absolute spherical t copula has density given by
+$c^{|t|}_{\nu}(u,v)=c^{t}_{\nu}\left(\frac{1-u}{2},\frac{1-v}{2}\right)=c^{t}_{\nu}\left(\frac{1-u}{2},\frac{1+v}{2}\right)=c^{t}_{\nu}\left(\frac{1+u}{2},\frac{1-v}{2}\right)=c^{t}_{\nu}\left(\frac{1+u}{2},\frac{1+v}{2}\right).$
+It has an asymptote at $(1,1)$ but takes a finite value at $(0,0)$ given by
+$c^{|t|}_{\nu}(0,0)=\frac{\Gamma\left(\frac{\nu+2}{\nu}\right)\Gamma\left(\frac{\nu}{2}\right)}{\Gamma\left(\frac{\nu+1}{2}\right)^{2}}.$
+This behaviour facilitates the maximum likelihood estimation of fulcrum parameters in models that incorporate v-transforms.
+Integrating any of the four equalities for $c^{|t|}_{\nu}$ and observing the boundary conditions for a copula gives us an expression for $C^{|t|}_{\nu}$.
+For example, integrating the final equality gives
+$C^{|t|}_{\nu}(u,v)=4C_{\nu}^{t}\left(\frac{1+u}{2},\frac{1+v}{2}\right)-u-v-1.$ (S1)
+The h-functions for the copula may be calculated to be
+$h^{|t|}_{\nu,1}(u,v)$ $=\frac{\partial}{\partial u}C^{|t|}_{\nu}(u,v)=2t_{\nu+1}\left(\sqrt{\frac{\nu+1}{\nu+t_{\nu}^{-1}(\frac{1+u}{2})}}t_{\nu}^{-1}\left(\frac{1+v}{2}\right)\right)-1$ (S2)
+$h^{|t|}_{\nu,2}(u,v)$ $=\frac{\partial}{\partial v}C^{|t|}_{\nu}(u,v)=2t_{\nu+1}\left(\sqrt{\frac{\nu+1}{\nu+t_{\nu}^{-1}(\frac{1+v}{2})}}t_{\nu}^{-1}\left(\frac{1+u}{2}\right)\right)-1$
+where $t_{\nu}$ denotes the distribution function of a Student t distribution with $\nu$ degrees of freedom.
+For calculation purposes, the h-functions are easy to evaluate for all $\nu>0$, since the function $t_{\nu}$ is typically available in statistical software.
+This is not however the case for the copula $C^{|t|}_{\nu}$ itself, since available software for evaluating the t copula $C^{t}_{\nu,\rho}$ is currently restricted.
+The R package copula is restricted to integer $\nu$ while the package rvinecopulib is confined to $\nu\geqslant 2$.
+A one-dimensional numerical integral seems unavoidable for evaluating the copula for general $\nu$ and we solve this problem by integrating the h-functions in (S2).
+From expression (S1) we can infer that the absolute t copula interpolates between independence (as $\nu\to\infty$) and comonotonicity (as $\nu\to 0$).
+For the latter case this follows because the spherical t copula $C_{\nu}^{t}$ converges to $0.5M(u,v)+0.5W(u,v)$, an equal mixture of the comonotonicity and countermonotonicity copulas, as $\nu\to 0$ (McNeil et al.,, 2022).
+In view of the limiting behaviour of the copula, the Kendall’s tau value $\tau(\nu)$ of $C^{|t|}_{\nu}$ must range in the interval $(0,1)$ as $\nu$ ranges in the interval $(0,\infty)$.
+We can find values by numerical integration using the formula
+$\tau(\nu)=4\int_{0}^{1}\int_{0}^{1}C^{|t|}_{\nu}(u,v)c^{|t|}_{\nu}(u,v)\mathrm{d}u\mathrm{d}v-1=64\int_{0}^{0.5}\int_{0}^{0.5}C_{\nu}^{t}(u,v)c_{\nu}^{t}(u,v)\mathrm{d}u,\mathrm{d}v-1.$ (S3)
+For example, for $\nu\in\{4,2,1,0.5\}$ the corresponding values are $\{0.099,0.189,0.333,0.515\}$.
+The absolute spherical t copula has upper tail dependence (inherited from the t copula).
+However we can easily show that it is asymptotically independent in the lower tail.
+Appendix S5 Simulation study of the approximation of $C_{1}$
+In the simulation study, data are generated from ARCH(1) and GARCH(1,1)-type processes using the parameter values summarized in Table 1 and different innovation distributions, as recorded in the legend to the table.
+Analyses are based on simulated samples of 50000 points and the first-order copula $C_{1}$ is estimated by fitting a first-order D-vine to the data using the two-stage method developed by Chen and Fan, (2006) in which the marginal distribution of the process is estimated by the rescaled empirical distribution function.
 Models are compared according to their AIC values.
-The first copula fitted in Table 2 is $C^{t}_{\nu}$, the copula of the spherical t distribution, i.e.
-the t copula with the correlation parameter fixed at $\rho=0$; the only free parameter is the degree-of-freedom parameter $\nu$.
-We also label this copula iv-ast, to emphasize that it can be viewed as the inverse v-transformed version of the absolute spherical t copula $C^{|t|}_{\nu}$.
-The second model is the generalisation of $C^{t}_{\nu}$ consisting of an equal mixture of the t copula $C^{t}_{\nu,\rho}$ and its 90-degree rotation $C^{t}_{\nu,-\rho}$ for $\rho\geqslant 0$.
-According to AIC, the $C_{\nu}^{t}$ is the best model in the table in two of six cases, which are the cases where the innovations are Student t4.
-The mixture of two t copulas is an inferior model to $C_{\nu}^{t}$ in 5 of 6 cases, the only exception being GARCH with normal innovations.
-The maximum likelihood method for the t mixture returns an estimate $\rho\approx 0$ in those 5 cases, which accounts for the fact that the difference in AIC values is approximately 2.
-The next panel of 3 models contains equal mixtures of the Clayton, Gumbel and Joe copulas and their 90, 180 and 270 degree rotations.
-The Gumbel case is the equally-weighted special case of the convex Gumbel copula of Junker and May, (2005), as also used by Loaiza-Maya et al., (2018).
-The next 3 copulas in the first panel of Table 2 are inverse-v-transformed versions of the upper tail-dependent Gumbel, Joe and Clayton survival copulas (the latter described as Clayton-180), i.e.
-copulas with densities of the form $c^{*}(|2u-1|,|2v-1|)$ where $c^{*}$ is the density of either the Clayton survival copula, the Gumbel copula or the Joe copulas.
-Comparing all of the 1-parameter models, the $C_{\nu}^{t}$ (iv-ast) copula is always the best model; second place is usually occupied by the iv-Joe copula, except for ARCH with Gaussian innovations where the iv-Clayton180 copula is favoured.
-Interestingly, while the iv-Gumbel copula is the worst of the inverse-v-transformed copulas, the mixture of Gumbels is often the best of the mixtures.
-The remaining models in Table 2 are non-exchangeable.
-The fourth panel of the table contains mixture copulas, except that the Khoudraji extension in (21) is applied to each copula before the mixture of the 4 different rotations is constructed.
-Thus these models introduce two further parameters $a_{1}$ and $a_{2}$.
-The final panel contains inverse-v-transformed versions of the Khoudraji extensions of the Clayton survival, Gumbel and Joe copulas.
-It would be appealing to add the Khoudraji construction based on the iv-ast copula, but this model is computationally more demanding and not currently a viable candidate model for practical applications.
-The best of the 3-parameter non-exchangeable models are always in the inverse-v-transformed class.
-The mixture copulas are only really competitive for t2.5 distributed innovations.
-The best-fitting copulas for Gaussian innovations (iv-K-Clayton-190 for ARCH and iv-K-Joe for GARCH) give estimates of $a_{2}$ which are very close to zero, while the best-fitting copula for t2.5 innovations (iv-K-Joe) give estimates of $a_{1}$ which are very close to zero.
-This makes sense in relation to Figure 2 since the non-exchangeability in the copulas seems to be caused by a “twist” in different directions.
-It is notable that the best model of all is iv-ast in the two cases where the innovations are t4.
-In the case of ARCH, this is in accord with the near-exchangeability evident in Figures 1 and 2.
-The estimated degree of freedom parameters are 2.56 for ARCH and 2.67 for GARCH.
+The results of the jointly symmetric case are documented in Section 3.5.1.
+Further experiments are documented here.
+S5.1 Simulation study of h-symmetric case with leverage
 Model ARCH(1) GARCH(1,1)
-Copula $p$ | $(\epsilon_{t})$ Gauss t4 t2.5 Gauss t4 t2.5
-$C^{t}_{\nu}$ (iv-ast) 1 482 0 115 56 0 77
-mix-t 2 484 2 117 54 2 79
-mix-Clayton 1 1630 426 178 444 284 127
-mix-Gumbel 1 867 150 171 181 113 138
-mix-Joe 1 1936 593 265 619 424 214
-iv-Clayton-180 1 483 146 281 127 98 198
-iv-Gumbel 1 1121 444 368 375 300 272
-iv-Joe 1 604 50 134 97 58 113
-mix-K-Clayton 3 1124 430 91 338 281 52
-mix-K-Gumbel 3 728 154 55 152 117 46
-mix-K-Joe 3 1634 597 167 585 428 133
-iv-K-Clayton-180 3 0 147 145 42 91 101
-iv-K-Gumbel 3 552 448 210 283 296 141
-iv-K-Joe 3 47 54 0 0 56 0
-Table 2: Comparison of copula fits when jointly symmetric copulas are fitted to simulated data from symmetric ARCH(1) and GARCH(1,1) processes by the PML method.
-Each column shows the differences between AIC and the minimum value of AIC for that column; thus a 0 indicates the model with lowest AIC.
-5.3 Simulation study of h-symmetric case with leverage
-Model ARCH(1) GARCH(1,1)
-Copula $p$ | $(\epsilon_{t})$ Gauss t4 t2.5 Gauss t4 t2.5
-iv-ast 1 980 224 222 683 262 173
-iv-Clayton-180 1 1009 358 363 806 384 299
-iv-Joe 1 1114 267 237 727 319 200
-iva-ast 2 258 0 131 23 0 68
-iva-Clayton-180 2 364 152 280 209 139 202
-iva-Joe 2 374 48 152 54 60 100
-iv-K-Clayton-180 3 655 361 239 715 386 238
-iv-K-Joe 3 672 271 113 619 323 132
-iva-K-Clayton-180 4 22 153 131 151 136 112
-iva-K-Joe 4 0 52 0 0 61 0
-Table 3: Comparison of AIC values when various h-symmetric copulas are fitted to simulated data from ARCH(1) and GARCH(1,1) processes with symmetric innovations and leverage by the PML method.
+Copula (form of density) $p$ | $(\epsilon_{t})$ Gauss t4 t2.5 Gauss t4 t2.5
+$c^{|t|}_{\nu}(|2u-1|,|2v-1|)$ 1 1068 224 223 720 264 187
+$c_{\theta}^{\text{J}}(|2u-1|,|2v-1|)$ 1 1202 267 238 764 321 214
+$c_{\theta}^{\text{Cl},180}(|2u-1|,|2v-1|)$ 1 1097 358 364 843 386 313
+$c^{|t|}_{\nu}(\mathcal{V}(u;0.5,\kappa),|2v-1|)$ 2 346 0 132 60 2 82
+$c_{\theta}^{\text{J}}(\mathcal{V}(u;0.5,\kappa),|2v-1|)$ 2 462 48 153 91 62 114
+$c_{\theta}^{\text{Cl},180}(\mathcal{V}(u;0.5,\kappa),|2v-1|)$ 2 452 152 281 246 141 216
+$c_{\nu,a_{1},a_{2}}^{|t|}(|2u-1|,|2v-1|)$ 3 678 227 113 614 267 131
+$c_{\theta,a_{1},a_{2}}^{\text{J}}(|2u-1|,|2v-1|)$ 3 760 271 114 656 325 146
+$c_{\theta,a_{1},a_{2}}^{\text{Cl},180}(|2u-1|,|2v-1|)$ 3 743 361 240 752 388 252
+$c_{\nu,a_{1},a_{2}}^{|t|}(\mathcal{V}(u;0.5,\kappa),|2v-1|)$ 4 0 2 0 0 0 0
+$c_{\theta,a_{1},a_{2}}^{\text{J}}(\mathcal{V}(u;0.5,\kappa),|2v-1|)$ 4 88 52 1 37 63 14
+$c_{\theta,a_{1},a_{2}}^{\text{Cl},180}(\mathcal{V}(u;0.5,\kappa),|2v-1|)$ 4 110 153 132 188 138 126
+Table S1: Comparison of AIC values when various h-symmetric copulas are fitted to simulated data from ARCH(1) and GARCH(1,1) processes with symmetric innovations and leverage by the PML method.
 Each column shows the differences between AIC and the minimum value of AIC for that column; thus a 0 indicates the model with lowest AIC.
@@
 The parameterisation of the ARCH(1) model with leverage is the same as in Figure 2 while the parameters for the GARCH(1,1) are given in Table 1.
-In Table 3 we include a selection of the copulas from Table 2, omitting the mixture copulas, since the inverse-v-transformed copulas typically do better, and also omitting the iv-Gumbel copula, since iv-Clayton-180 and iv-Joe generally do better.
-We add copulas with the prefix ‘iva’ (referring to asymmetric inverse v-transform) which are models of the form $c(u,v)=c^{*}(\mathcal{V}_{1}(u),|2v-1|)$ where $\mathcal{V}_{1}$ is a member of the family $\mathcal{V}(u;\delta,\kappa)$ in (13) with $\delta=0.5$, so that only one extra unknown parameter $\kappa$ is introduced.
-The first panel contains jointly symmetric and exchangeable copulas; the second panel contains h-symmetric copulas based on an exchangeable copula for $c^{*}$ but incorporating an asymmetric v-transform for the first variable; the third panel contains jointly symmetric non-exchangeable copulas using the Khoudraji construction; the final panel contains h-symmetric copulas based on a non-exchangeable $c^{*}$ and incorporating an asymmetric v-transform for the first variable.
-For the two models with t4 innovations the best fitting copula is the absolute spherical t copula incorporating an asymmetric v-transform.
-For the other models the best fitting copula is a Khoudraji-transformed version of the Joe copula incorporating an asymmetric v-transform.
-Results are very similar for the ARCH(1) and GARCH(1,1) process.
-5.4 Simulation study of case of asymmetric innovations
+In Table S1 we include a selection of the copulas from Table 2, omitting the mixture copulas, since the inverse-v-transformed copulas typically do better, and also omitting the inverse-v-transformed Gumbel copula, since the other inverse-v-transformed copulas do better.
+We add copulas of the form $c(u,v)=c^{*}(\mathcal{V}_{1}(u),|2v-1|)$ where $\mathcal{V}_{1}$ is a member of the family $\mathcal{V}(u;\delta,\kappa)$ in (12) with $\delta=0.5$, so that only one extra unknown parameter $\kappa$ is introduced.
+The first panel contains jointly symmetric and exchangeable copulas; the second panel contains h-symmetric copulas based on an exchangeable copula for $c^{*}$ but incorporating an asymmetric v-transform $\mathcal{V}(u;0.5,\kappa)$ for the first variable; the third panel contains jointly symmetric non-exchangeable copulas using the Khoudraji construction (17); the final panel contains h-symmetric copulas based on a non-exchangeable $c^{*}$ and incorporating an asymmetric v-transform for the first variable.
+The best results in all cases are obtained from copulas developed from the absolute spherical t copula $C^{|t|}_{\nu}$ that are not jointly symmetric and that incorporate the non-symmetric v-transform, as theory suggests.
+In 5 of 6 cases the inverse v-transformation of the Khoudraji non-exchangeable extension of $C^{|t|}_{\nu}$ is selected while in one case (ARCH(1) with t4 innovations) the inverse v-transformation of $C^{|t|}_{\nu}$ is selected.
+Models based on the Joe copula also do well.
+S5.2 Simulation study of case of asymmetric innovations
 Model ARCH(1) GARCH(1,1)
-Copula $p$ leverage leverage
-iv-ast 1 168 679 49 436
-iv-Clayton-180 1 376 881 208 673
-iv-Joe 1 220 728 92 487
-ivl-ivl-ast 3 116 201 23 92
-ivl-ivl-Clayton-180 3 347 558 189 403
-ivl-ivl-Joe 3 164 231 64 140
-ivl-ivl-K-Clayton-180 5 336 561 191 403
-ivl-ivl-K-Joe 5 168 235 68 144
-iv2-iv2-ast 5 0 0 0 0
-iv2-iv2-Clayton-180 5 228 293 159 277
-iv2-iv2-Joe 5 47 24 45 50
-iv2-iv2-K-Clayton-180 7 221 297 160 279
-iv2-iv2-K-Joe 7 50 28 49 54
-Table 4: Comparison of AIC values when various copulas are fitted to simulated data from ARCH(1) and GARCH(1,1) processes with asymmetric innovations by the PML method.
+Copula (form of density) $p$ | leverage absent present absent present
+$c^{|t|}_{\nu}(|2u-1|,|2v-1|)$ 1 169 679 49 436
+$c_{\theta}^{\text{J}}(|2u-1|,|2v-1|)$ 1 221 728 92 487
+$c_{\theta}^{\text{Cl},180}(|2u-1|,|2v-1|)$ 1 377 881 208 673
+$c^{|t|}_{\nu}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))$ 3 117 201 23 92
+$c^{\text{J}}_{\nu}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))$ 3 165 231 64 140
+$c^{\text{Cl},180}_{\nu}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))$ 3 348 558 189 403
+$c^{|t|}_{\nu,a_{1},a_{2}}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))$ 5 116 205 25 94
+$c^{\text{J}}_{\nu,a_{1},a_{2}}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))$ 5 169 235 68 144
+$c^{\text{Cl},180}_{\nu,a_{1},a_{2}}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))$ 5 337 561 191 403
+$c^{|t|}_{\nu}(\mathcal{V}(u;\delta_{1},\kappa_{1}),\mathcal{V}(v;\delta_{2},\kappa_{2}))$ 5 1 0 0 0
+$c^{\text{J}}_{\nu}(\mathcal{V}(u;\delta_{1},\kappa_{1}),\mathcal{V}(v;\delta_{2},\kappa_{2}))$ 5 48 24 45 50
+$c^{\text{Cl},180}_{\nu}(\mathcal{V}(u;\delta_{1},\kappa_{1}),\mathcal{V}(v;\delta_{2},\kappa_{2}))$ 5 229 293 159 277
+$c^{|t|}_{\nu,a_{1},a_{2}}(\mathcal{V}(u;\delta_{1},\kappa_{1}),\mathcal{V}(v;\delta_{2},\kappa_{2}))$ 7 0 4 1 4
+$c^{\text{J}}_{\nu,a_{1},a_{2}}(\mathcal{V}(u;\delta_{1},\kappa_{1}),\mathcal{V}(v;\delta_{2},\kappa_{2}))$ 7 51 28 49 54
+$c^{\text{Cl},180}_{\nu,a_{1},a_{2}}(\mathcal{V}(u;\delta_{1},\kappa_{1}),\mathcal{V}(v;\delta_{2},\kappa_{2}))$ 7 222 297 160 279
+Table S2: Comparison of AIC values when various copulas are fitted to simulated data from ARCH(1) and GARCH(1,1) processes with asymmetric innovations by the PML method.
 Each column shows the differences between AIC and the minimum value of AIC for that column; thus a 0 indicates the model with lowest AIC.
@@
 We have already observed that asymmetric innovations remove all of the previous symmetries we have discussed and so our approach in this section is pure experimentation without the theoretical insights we exploited in the previous experiments.
-The first set of 3 models in Table 4 consists of jointly symmetric and exchangeable copulas, included as a benchmark.
-The iv-ast copula is the best of these but substantial improvements can be made by introducing asymmetry.
-The second set of 3 models is based on copula densities of the form $c(u,v)=c^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))$ where $\mathcal{V}(\cdot;\delta)$ is the linear v-transform with fulcrum at $\delta$ (asymmetric when $\delta\neq 0.5$) and $c^{*}$ is either the ast, Clayton survival or Joe copula [^4] It should be noted that full maximum likelihood estimation with respect to all parameters including the parameters of the v-transform presents no problems for these models since the ast, Clayton survival and Joe copula densities do not have asymptotes at $(0,0)$.
-If this were not the case we can run into a threshold estimation issue, as discussed in Bladt and McNeil, 2022b ..
-These obviously lower the AIC values considerably at the expense of 2 extra parameters.
-The third set of 2 models applies the Khoudraji transformation to $c^{*}$ in the case of Clayton and Joe, but this only offers an improvement in one case (Clayton-180 and ARCH(1) process).
+The first set of 3 models in Table S2 consists of jointly symmetric and exchangeable copulas, included as a benchmark.
+The spherical t copula is the best of these but substantial improvements can be made by introducing asymmetry.
+The second set of 3 models is based on copula densities of the form $c(u,v)=c^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))$ where $\mathcal{V}(\cdot;\delta)$ is the linear v-transform with fulcrum at $\delta$ (asymmetric when $\delta\neq 0.5$) and $c^{*}$ is either the ast, Clayton survival or Joe copula.
+It should be noted that full maximum likelihood estimation with respect to all parameters including the parameters of the v-transform presents no problems for these models since the ast, Clayton survival and Joe copula densities do not have asymptotes at $(0,0)$.
+If this were not the case we can run into a threshold estimation issue, as discussed in Bladt and McNeil, 2022b .
+Adding the parameters $\delta_{1}$ and $\delta_{2}$ lowers the AIC values considerably.
+The third set of 3 models applies the Khoudraji transformation to $c^{*}$ before application of the inverse linear v-transform, but this barely offers any improvement on the previous set of models; the AIC is only really lowered for the Clayton survival copula in the ARCH(1) model without leverage, and survival Clayton is the worst of the three candidates in this case.
 The fourth set of models is based on copula densities $c(u,v)=c^{*}(\mathcal{V}(u;\delta_{1},\kappa_{1}),\mathcal{V}(v;\delta_{2},\kappa_{2}))$ where $\mathcal{V}(\cdot;\delta,\kappa)$ is the 2-parameter v-transform in (12); such v-transforms are both non-linear and asymmetric.
-These models give the best results of all, with the model where $c^{*}$ is the ast copula taking first place for all 4 datasets, followed by the model based on Joe.
-Addition of the Khoudraji transformation in the final set of models again offers no discernible improvement to the models based on Joe.
-It seems that allowing asymmetric v-transform largely removes the need to consider the Khoudraji transformation, even when the v-transforms are linear.
-6 Mimicking GARCH(1,1) processes with simplified d-vines
-6.1 Practical issues in the construction of simplified d-vines
-In the final sections of this paper we provide some guidance to the construction of simplified d-vine copula models to mimic the behaviour of GARCH-type processes and offer potential alternative models for financial time series.
-The previous sections have given some insight into the kinds of pair copulas we should consider.
-We now turn to some of the issues that arise when we want to embed these in simplified vine structures.
-Equation (1) in the Introduction implies that the density $c_{(n)}$ of the copula of the variables $(X_{1},\ldots,X_{n})^{\top}$ in a strictly stationary time series model, such as GARCH, can be written in the form
-$c_{(n)}(u_{1},\ldots,u_{n})=\\ \prod_{k=1}^{n-1}\prod_{j=k+1}^{n}c_{k}(B_{k-1}(u_{j-k};\bm{u}_{[j-k+1:j-1]}),R_{k-1}(u_{j};\bm{u}_{[j-1:j-k+1]})\mid F_{X}^{-1}(u_{j-k+1}),\ldots,F_{X}^{-1}(u_{j-1}))$ (22)
-where, for $k\in\mathbb{N}$, we define the functions $R_{k}:(0,1)\times(0,1)^{k}\to(0,1)$ and $B_{k}:(0,1)\times(0,1)^{k}$
-$R_{k}(x;\bm{u})$ $=$ $F_{X_{k+1}|\bm{X}_{[1:k]}}(F_{X}^{-1}(x)\mid F_{X}^{-1}(u_{k}),\ldots,F_{X}^{-1}(u_{1}))$
-$B_{k}(x;\bm{u})$ $=$ $F_{X_{1}|\bm{X}_{[2:k+1]}}(F_{X}^{-1}(x)\mid F_{X}^{-1}(u_{1}),\ldots,F_{X}^{-1}(u_{k}))$
-which we refer to as forward and backward Rosenblatt functions, respectively.
-The notation is based on the one used in Bladt and McNeil, 2022a with minor changes.
-Note that, in (22), we adopt the convention that $R_{0}(x;\cdot)=B_{0}(x;\cdot)=x$ for all $x\in(0,1)$.
-In practice we want to mimic the copula (22) with a simplified d-vine, i.e. a model where the copula has joint density of the simpler form
-$c_{(n)}(\bm{u})=\prod_{k=1}^{n-1}\prod_{j=k+1}^{n}c_{k}\Big{(}B_{k-1}(u_{j-k};\bm{u}_{[j-k+1:j-1]}),R_{k-1}(u_{j};\bm{u}_{[j-1:j-k+1]})\Big{)}.$ (23)
-In this case the Rosenblatt functions are determined in a recursive, interlacing fashion by $R_{1}(x;u)=h_{1,1}(u,x)$, $B_{1}(x;v)=h_{1,2}(x,v)$ and
-[…]
-where $h_{k,1}(u,v)=\frac{\partial}{\partial u}C_{k}(u,v)$ and $h_{k,2}(u,v)=\frac{\partial}{\partial v}C_{k}(u,v)$ are the so-called $h$-functions of the copula $C_{k}$ with density $c_{k}$.
-It seems unlikely that we can perfectly mimic a GARCH-type process with a simplified d-vine, since the true copula (22) cannot be shown to conform to the simplifying assumption.
-However, the weak dependence of $c_{2}(u,v;w)$ on the conditioning variable investigated in Section 3.5 gives some optimism that we could at least approximate a GARCH-type model with a simplified d-vine.
-Other than ARCH(1)-type processes, which are Markov processes of order 1, GARCH(1,1)-type processes have infinite memory in the sense that $X_{t}$ depends on all previous values $X_{t-k}$, for $k\in\mathbb{N}$, albeit the serial dependence on past observations dies away relatively quickly; for example, the autocorrelation function of the squared values of the classical GARCH(1,1) process decays like an ARMA(1,1) process.
-To mimic a GARCH(1,1)-type process, it seems natural to consider d-vine processes with infinite-order partial copula dependence as in Bladt and McNeil, 2022a .
-These are models which are built from a sequence of pair copulas $(C_{k})_{k\in\mathbb{N}}$ which is parsimoniously parameterized using the Kendall partial autocorrelation function (kpacf) or, in other words, the sequence of Kendall rank correlations of $(C_{k})_{k\in\mathbb{N}}$.
-GARCH(1,1) models are very parsimoniously parameterized.
-For example, a model that incorporates both leverage and skewed Student t innovations would have only 6 parameters ($\alpha_{0}$, $\alpha_{1}$, $\beta_{1}$, $\gamma_{1}$, $\nu$, $\lambda$).
-In our d-vine analogues we also need to control the number of parameters and the method of parameterization via the kpacf developed in Bladt and McNeil, 2022a can be adapted to our context by applying it to an inverse-v-transformed version of the copula series $(C^{*}_{k})_{k\in\mathbb{N}}$.
-To be more precise, each pair copula $C_{k}$ used in the sequence defining the d-vine would be constructed to have density $c_{k}(u,v)=c_{k}^{*}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v))$ where the v-transforms $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ would be two different, probably asymmetric functions with a fixed form for all $k$; in essence they would take care of the effects caused by leverage and asymmetric innovations in a similar manner to $\gamma_{1}$ and $\lambda$ in the aforementioned GARCH(1,1) model.
-The copula sequence $(C^{*}_{k})_{k\in\mathbb{N}}$ would contain Joe, Clayton survival or ast copulas with Kendall correlations $\tau_{k}$ that coincided with the terms of the kpacf of a classical Gaussian ARMA(1,1) process, or some other low-order linear process.
-The combination of ARMA(1,1) and two linear v-transforms would give precisely 4 parameters to describe the d-vine copula process, although two or more parameters would also be required to model the marginal distribution of $(X_{t})_{t\in\mathbb{Z}}$.
-We have listed the ast copula $C^{|t|}_{\nu}$ as a candidate pair copula for the sequence $(C^{*}_{k})_{k\in\mathbb{N}}$, but it is not clear whether there is a simple parametric formula to link Kendall’s tau to the degree of freedom parameter $\nu$.
-Simple calculations from (20) show that
-$\tau(\nu)=64\int_{0}^{0.5}\int_{0}^{0.5}C_{\nu}^{t}(u,v)c_{\nu}^{t}(u,v)\mathrm{d}u,\mathrm{d}v-1$ (25)
-where $C_{\nu}^{t}$ and $c_{\nu}^{t}$ are the spherical t copula and its density, and that $\tau(\nu)$ ranges in the interval $(0,1)$ as $\nu$ ranges in the interval $(0,\infty)$.
-An analytical expression for $\tau(\nu)$ would make the ast copula a very attractive choice for these models but, in the absence of such an expression, the Joe copula may be the easier option.
-Finally we consider the choice of v-transforms for $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$.
-To estimate a d-vine model by maximum likelihood we have to be able to evaluate a likelihood based on (23) which means evaluating the Rosenblatt forward and backward functions using the recursive expressions (24).
-This in turn entails evaluating multiple $h$-functions.
-Moreover, to generate realizations from a copula with joint density (23), we have to be able to invert the Rosenblatt forward functions which entails the evaluation of inverse $h$-functions.
-To this end it is convenient to have closed form expressions for the copulas $C_{k}$ and their $h$-functions and inverse $h$-functions.
-For copula densities $c_{k}$ obtained from the inverse-v-transformation construction in (15) this is unfortunately not straightforward, except in the case where we have copula densities of the form $c(u,v)=c^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))$ constructed from two linear v-transforms $\mathcal{V}_{\delta_{1}}$ and $\mathcal{V}_{\delta_{2}}$ and a copula density $c^{*}$ which itself has a tractable joint distribution function and $h$-functions.
-We investigate this in the next section.
-6.2 Tractable processes using linear v-transforms
+These models give the best results of all, with the model where $c^{*}$ is the ast copula best for 3 out of 4 datasets datasets, followed by the model based on Joe.
+Addition of the Khoudraji transformation in the final set of models only offers an improvement for the ARCH(1) model without leverage, and this improvement is very minor.
+We conclude that allowing asymmetric v-transform largely removes the need to consider the Khoudraji transformation, even when the v-transforms are linear.
+This has positive consequences for practical work: evaluating the density of Khoudraji-extended copulas of type (17) is computationally challenging since the density involves the copula $C_{\theta}$ and its h-functions, as well as is density $c_{\theta}$.
+This is particularly burdensome when $C_{\theta}=C^{|t|}_{\nu}$, the ast copula, as numerical integration is required to evaluate $C^{|t|}_{\nu}$.
+Appendix S6 D-vine processes using pair copulas based on linear v-transforms
+S6.1 Calculating $h$-functions
 The following results shows that all the formulas required to work with linear v-transforms are easily obtained.
-Proposition 7.
+Proposition S3.
 The copula $C$ with density given by $c(u,v)=c^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))$ is
-$C(u,v)=\delta_{1}^{I_{\{u\leqslant\delta_{1}\}}}(\delta_{1}-1)^{I_{\{u>\delta_{1}\}}}\delta_{2}^{I_{\{v\leqslant\delta_{2}\}}}(\delta_{2}-1)^{I_{\{v>\delta_{2}\}}}C^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))+\delta_{1}v+\delta_{2}u-\delta_{1}\delta_{2}$ (26)
+$C(u,v)=\delta_{1}^{I_{\{u\leqslant\delta_{1}\}}}(\delta_{1}-1)^{I_{\{u>\delta_{1}\}}}\delta_{2}^{I_{\{v\leqslant\delta_{2}\}}}(\delta_{2}-1)^{I_{\{v>\delta_{2}\}}}C^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))+\delta_{1}v+\delta_{2}u-\delta_{1}\delta_{2}$ (S4)
 where $C^{*}$ is the copula with density $c^{*}$. Writing $h_{1}^{*}$ and $h_{2}^{*}$ for the partial derivatives of $C^{*}$, the corresponding partial derivatives $h_{1}$ and $h_{2}$ of $C$ are
@@
 and satisfy the equations
-$\mathcal{V}\left(h_{1}(u,v);\delta_{2}\right)$ $=h_{1}^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))$ (27)
+$\mathcal{V}\left(h_{1}(u,v);\delta_{2}\right)$ $=h_{1}^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))$ (S5)
 $\mathcal{V}\left(h_{2}(u,v);\delta_{1}\right)$ $=h_{2}^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2})).$
@@
 Proof.
-First note that $\frac{\partial^{2}}{\partial u\partial v}C(u,v)=c^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))$.
-We just need to check that the expression (26) defines a copula, which entails checking that $C(0,v)=0$ and $C(1,v)=v$ for any $v$; by symmetry it would follow that $C(u,0)=0$ and $C(u,1)=u$ for any $u$.
-Now
-$C(0,v)=\delta_{1}\delta_{2}^{I_{\{v\leqslant\delta_{2}\}}}(\delta_{2}-1)^{I_{\{v>\delta_{2}\}}}\mathcal{V}(v;\delta_{2})+\delta_{1}v-\delta_{1}\delta_{2}$
-and this yields 0 for both $v\leqslant\delta_{2}$ and $v>\delta_{2}$; moreover
-$C(1,v)=(\delta_{1}-1)\delta_{2}^{I_{\{v\leqslant\delta_{2}\}}}(\delta_{2}-1)^{I_{\{v>\delta_{2}\}}}\mathcal{V}(v;\delta_{2})+\delta_{1}v+\delta_{2}(1-\delta_{1})$
-and this yields $v$ for both $v\leqslant\delta_{2}$ and $v>\delta_{2}$.
-The $h$-functions are obtained by differentiation of (26) and the identities (27) are easy to verify.
+The joint density $c$ of the copula $C$ defined in  (S4) can be easily shown to satisfy $c(u,v)=\frac{\partial^{2}}{\partial u\partial v}C(u,v)=c^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))$.
+We just need to check that it is a valid copula, which entails checking that $C(0,v)=0$ and $C(1,v)=v$ for any $v$, which is straightforward.
+By symmetry, it then follows that $C(u,0)=0$ and $C(u,1)=u$ for any $u$.
+The $h$-functions are obtained by differentiation of (S4) and the identities (S5) are easy to verify.
 Noting that $h_{1}(u,\delta_{2})=\delta_{2}$ and $h_{2}(u,\delta_{1})=\delta_{1}$, it is also straightforward to derive the inverses of the $h$-functions.
 ∎
-We end by considering the special case of the model with linear v-transforms in which the two v-transforms are identical.
-We find that the resulting d-vine copula also belongs to the class of vt-d-vines in Bladt and McNeil, 2022b ; these models have been used with real financial data and found to be tractable and effective.
-The statement of the result and proof use the notation $\mathcal{V}(\bm{u};\delta)$ for the element-by-element v-transformation of a vector $\bm{u}$ with the same linear v-transform.
+S6.2 Link to vt-D-vine models
+In the special case of a model with two identical linear v-transforms we now show that the resulting D-vine copula also belongs to the class of vt-D-vines in Bladt and McNeil, 2022b which has been applied successfully to real financial data.
+The use of two different linear v-transforms does however give much more flexibility.
+For this result we introduce sequences of conditional distribution functions (Rosenblatt functions) $B_{k}:(0,1)\times(0,1)^{k}\to(0,1)$ and $R_{k}:(0,1)\times(0,1)^{k}\to(0,1)$ defined by
+$\begin{aligned} B_{k-1}(x;\bm{u})&=\mathbb{P}(U_{1}\leqslant x\mid U_{2}=u_{1},\ldots,U_{k}=u_{k-1})=F_{U_{1}\mid\bm{U}_{[2:k]}}(x\mid\bm{u}_{[1:k-1]}),\\ R_{k-1}(x;\bm{u})&=\mathbb{P}(U_{k}\leqslant x\mid U_{k-1}=u_{1},\ldots,U_{1}=u_{k-1})=F_{U_{k}\mid\bm{U}_{[1:k-1]}}(x\mid\bm{u}_{[k-1:1]}),\end{aligned}\quad k\in\mathbb{N}.$ (S6)
+Note the reversal of the order of the conditioning values $\bm{u}_{[k-1:1]}=(u_{k-1},\ldots,u_{1})^{\top}$ in the second equation.
+Note also that we interpret $B_{0}(x;\cdot)=R_{0}(x;\cdot)=x$.
+The Rosenblatt functions satisfy $R_{1}(x;u)=h_{1,1}(u,x)$, $B_{1}(x;v)=h_{1,2}(x,v)$ and
+$\begin{aligned} R_{k}(x;\bm{u})&=h_{k,1}\left(B_{k-1}(u_{k};\bm{u}_{[k-1:1]}),R_{k-1}(x;\bm{u}_{[1:k-1]})\right),\\ B_{k}(x;\bm{u})&=h_{k,2}\left(B_{k-1}(x;\bm{u}_{[1,k-1]}),R_{k-1}(u_{k};\bm{u}_{[k-1:1]})\right),\end{aligned}\quad k\geqslant 2,$ (S7)
+where $h_{k,1}(u,v)=\frac{\partial}{\partial u}C_{k}(u,v)$ and $h_{k,2}(u,v)=\frac{\partial}{\partial v}C_{k}(u,v)$ are the $h$-functions of the copula $C_{k}$ with density $c_{k}$.
+This allows us to write the density in (18) as
+$c_{U_{1},\ldots,U_{n}}(u_{1},\ldots,u_{n})=\prod_{k=1}^{n-1}\prod_{j=k+1}^{n}c_{k}\Big(B_{k-1}(u_{j-k};\bm{u}_{[j-k+1:j-1]}),R_{k-1}(u_{j};\bm{u}_{[j-1:j-k+1]})\Big).$ (S8)
 Theorem 3.
-Suppose we construct a d-vine copula (23) using a sequence of copulas $(C_{k})_{k\in\mathbb{N}}$ whose densities are all of the form $c_{k}(u,v)=c_{k}^{*}(\mathcal{V}(u;\delta),\mathcal{V}(v;\delta))$ for a sequence of copulas $(C_{k}^{*})_{k\in\mathbb{N}}$ with densities $(c_{k}^{*})_{k\in\mathbb{N}}$ and a single linear v-transform $\mathcal{V}(\cdot,\delta)$.
-Then
-$c_{(n)}(\bm{u})=\prod_{k=1}^{n-1}\prod_{j=k+1}^{n}c_{k}^{*}\Big{(}B^{*}_{k-1}(\mathcal{V}(u_{j-k};\delta);\mathcal{V}(\bm{u}_{[j-k+1:j-1]};\delta)),R^{*}_{k-1}(\mathcal{V}(u_{j};\delta);\mathcal{V}(\bm{u}_{[j-1:j-k+1]};\delta))\Big{)}$ (28)
-where $(R_{k}^{*})_{k\in\mathbb{N}}$ and $(B_{k}^{*})_{k\in\mathbb{N}}$ are the forward and backward Rosenblatt functions corresponding to the copula sequence $(C_{k}^{*})_{k\in\mathbb{N}}$.
+Suppose we construct a d-vine copula (S8) using a sequence of copulas $(C_{k})_{k\in\mathbb{N}}$ whose densities are all of the form $c_{k}(u,v)=c_{k}^{*}(\mathcal{V}(u;\delta),\mathcal{V}(v;\delta))$ for a sequence of copulas $(C_{k}^{*})_{k\in\mathbb{N}}$ with densities $(c_{k}^{*})_{k\in\mathbb{N}}$ and a single linear v-transform $\mathcal{V}(\cdot,\delta)$.
+Then the joint density $c_{U_{1},\ldots,U_{n}}(u_{1},\ldots,u_{n})$ may be written as
+$\prod_{k=1}^{n-1}\prod_{j=k+1}^{n}c_{k}^{*}\Big(B^{*}_{k-1}(\mathcal{V}(u_{j-k};\delta);\mathcal{V}(\bm{u}_{[j-k+1:j-1]};\delta)),R^{*}_{k-1}(\mathcal{V}(u_{j};\delta);\mathcal{V}(\bm{u}_{[j-1:j-k+1]};\delta))\Big)$ (S9)
+where $(R_{k}^{*})_{k\in\mathbb{N}}$ and $(B_{k}^{*})_{k\in\mathbb{N}}$ are the forward and backward Rosenblatt functions for the copula sequence $(C_{k}^{*})_{k\in\mathbb{N}}$ and where we use the compact notation $\mathcal{V}(\bm{u};\delta)=(\mathcal{V}(u_{1};\delta),\ldots,\mathcal{V}(u_{d};\delta))^{\top}$.
 Proof.
 The result will follow if we can show that the equalities
-$\mathcal{V}\left(B_{k-1}(u;\bm{v});\delta\right)=B^{*}_{k-1}(\mathcal{V}(u;\delta);\mathcal{V}(\bm{v};\delta))\quad\text{and}\quad\mathcal{V}\left(R_{k-1}(u;\bm{v});\delta\right)=R^{*}_{k-1}(\mathcal{V}(u;\delta);\mathcal{V}(\bm{v};\delta))$ (29)
+$\mathcal{V}\left(B_{k-1}(u;\bm{v});\delta\right)=B^{*}_{k-1}(\mathcal{V}(u;\delta);\mathcal{V}(\bm{v};\delta))\quad\text{and}\quad\mathcal{V}\left(R_{k-1}(u;\bm{v});\delta\right)=R^{*}_{k-1}(\mathcal{V}(u;\delta);\mathcal{V}(\bm{v};\delta))$ (S10)
 hold for any $k\in\mathbb{N}$, $u\in(0,1)$ and $\bm{v}\in(0,1)^{k-1}$.
 We will argue by induction.
-The equalities (29) are obviously true for $k=1$ since, by our convention, $B_{0}(u;\cdot)=B_{0}^{*}(u;\cdot)=u$ and $R_{0}(u;\cdot)=R_{0}^{*}(u;\cdot)=u$ for all $u$.
-Assume the equalities (29) are true when $k=l\in\mathbb{N}$ and consider $k=l+1$.
-For the first equality we can show, by first using formula (27) in Proposition 7 and then applying the inductive assumption, that
+The equalities (S10) are obviously true for $k=1$ since, by our convention, $B_{0}(u;\cdot)=B_{0}^{*}(u;\cdot)=u$ and $R_{0}(u;\cdot)=R_{0}^{*}(u;\cdot)=u$ for all $u$.
+Assume the equalities (S10) are true when $k=l\in\mathbb{N}$ and consider $k=l+1$.
+For the first equality we can show, by first using formula (S5) in Proposition S3 and then applying the inductive assumption, that
 $\mathcal{V}\left(B_{l}(u;\bm{v});\delta\right)$ $=\mathcal{V}\left(h_{l,2}\left(B_{l-1}(u;\bm{v}_{[1,k-1]}),R_{l-1}(v_{k};\bm{v}_{[k-1:1]})\right);\delta\right)$
@@
 and the same argument can be applied to the second equality. ∎
-Equation (28) describes a vt-d-vine copula process.
+Equation (S9) describes a vt-D-vine copula process.
 In such a model the copula density can be evaluated by first applying the v-transform to the entire vector $(u_{1},\ldots,u_{n})$ and then passing the resulting values to the density of a d-vine copula based on the sequence $(C_{k}^{*})_{k\in\mathbb{N}}$.
-McNeil, (2021) shows that if we apply such a copula process to a series $(X_{t})_{t\in\mathbb{Z}}$ we effectively use the copula sequence $(C_{k}^{*})_{k\in\mathbb{N}}$ to model the serial dependence structure of an element-wise transformation of the time series $(T(X_{t}))_{t\in\mathbb{Z}}$.
-The case where the copulas are all Gaussian corresponds to the vt-ARMA model in McNeil, (2021), although this tends to be an inferior model to the ones based on non-Gaussian d-vines in Bladt and McNeil, 2022b .
-The implied canonical form of $T$ for a linear v-transform $\mathcal{V}(\cdot;\delta)$ and a continuous marginal distribution function $F_{X}$ is the v-shaped function on $\mathbb{R}$ given by
-$T(x)=\begin{cases}F_{X}^{-1}\left(F_{X}(x)+\mathcal{V}(F_{X}(x);\delta)\right)-\mu&x\leqslant\mu,\\ x-\mu&x>\mu,\end{cases}$ (30)
-where $\mu=F_{X}^{-1}(\delta)$ [^5] The function $T$ in (30) is not the unique transformation that gives a time series whose dependence structure may be described by the joint density (28).
-There is, in fact, an equivalence class of transformations that all lead to the same copula density when paired with $F_{X}$.
-For example, in the case where $F_{X}$ is symmetric and $\delta=0.5$ we could also take $T(x)=x^{2}$.
-We pin down the canonical form by fixing the gradient to be 1 for $x>\mu$.
-See McNeil, (2021) for more details..
-For example, in the case where $F_{X}$ is symmetric around the origin and $\delta=0.5$, $\mu=0$ and $T(x)=|x|$.
-In general $T$ can be thought of as a function that measures the magnitude of a return relative to $\mu$ in a way that can also account for asymmetric effects for returns that are greater or less than $\mu$.
-7 Conclusion
-This paper has asked two questions.
-First, what are the bivariate copulas that underlie the serial dependencies and serial conditional dependencies of pairs of variables $(X_{t},X_{t+k})$ in first-order GARCH(1,1) models.
-Second, can these copulas be mimicked by simpler constructions with a view to developing simplified d-vines that can behave in a similar fashion to GARCH-type models, but offer more flexible solutions for datasets where GARCH works less well.
-With regard to the first question we find interesting symmetries in the dependence structures of the models with symmetric innovation distributions, even when leverage effects are introduced, but these symmetries disappear in the presence of asymmetric innovations, which are often necessary in practice to model real-world datasets.
-What can generally always be said is that the underlying bivariate copulas are cross-shaped, in the sense that there is a tendency for points to accumulate in all fours corners of the unit square, and that this tendency increases at shorter lags as the size of the ARCH effect increases and persists at longer lags as the aggregate size of the ARCH and GARCH effects increases.
-Heavier-tailed innovation distributions lead to weaker dependence.
-However, the true copulas and conditional copulas of GARCH-type processes remain inaccessible for practical use, mainly because the stationary distribution of these models is not available in closed form.
-This leads us to consider the second question of whether GARCH copulas can be mimicked with simpler copulas.
-Motivated by the recognition that v-transforms are implicitly embedded in the copulas of symmetric GARCH models, we propose a construction based on v-transforms and positive-dependence copulas with strong dependence in the upper joint tail; these constructions are naturally cross-shaped.
-Of the well-known copulas, Joe’s copula in combination with asymmetric v-transforms seems the most promising for mimicking GARCH.
-However, we also identify a new copula - the absolute spherical t copula - which seems to offer an even more promising basis for the construction.
-Could these copulas be used to construct simplified d-vines to mimic GARCH processes?
-To this question we give a more qualified answer.
-While the Markovian first-order ARCH-type models can definitely be mimicked by a d-vine(1) models, as soon as a GARCH effect is introduced, it is not clear whether the simplifying assumption will restrict our ability to perfectly mimic GARCH behaviour; a first investigation of this question in Section 5 produces some weak evidence that the dependence of the conditional copulas on the conditioning variables may not be particularly strong, but more investigation is required.
-Moreover, the computational requirements of d-vines, as discussed in Section 6, mean that we may have difficulty in integrating the most promising GARCH-mimicking pair copulas (such as iv2-iv2-ast) into d-vine structures and we may have to compromise and use copulas based on linear v-transforms.
-Further examination of these issues is required.
-Could the GARCH-mimicking copulas be used to construct d-vines to model real financial return data?
-This is a different question to the question of mimicking GARCH processes.
-Bladt and McNeil, 2022b  have shown that the class of vt-d-vine processes with linear v-transforms is competitive for modelling many financial return series and, in particular, can give better results for data which depart from the power-lawed marginal behaviour modelled by a GARCH process.
-The d-vine processes constructed from GARCH-mimicking pair copulas, as described in this paper, can offer additional flexibility because of the use of two v-transforms in their construction, even if we do restrict both to be linear.
-There is a need for further empirical research to evaluate the potential of such models.
-Another issue requiring further research is the construction of d-vines to mimic ARMA-GARCH processes.
-Real-world financial return series sometimes show serial correlation in the raw data as well as in the squared or absolute value series and this is usually modelled by combining ARMA and GARCH effects.
-In this paper we have concentrated on pure GARCH-type models and given only a single example in Section 3.4 of an AR-ARCH model.
-The illustration in the lower-right panel of Figure 3 suggests mixtures of positive dependence copulas and GARCH-mimicking copulas could be successful in addressing combined ARMA and GARCH effects.
-As a pragmatic alternative, the d-vine models discussed in this paper could be applied to the residuals from an ARMA model instead.
+McNeil, (2021) shows that if we apply such a copula process to a series $(X_{t})_{t\in\mathbb{Z}}$ we effectively use the copula sequence $(C_{k}^{*})_{k\in\mathbb{N}}$ to model the serial dependence structure of the time series $(T(X_{t}))_{t\in\mathbb{Z}}$ where $T$ is a v-shaped transformation on $\mathbb{R}$.
+Note that the Theorem above is a little more general than is needed.
+The copula sequence in our application is derived from a single parametric copula family $C^{*}$ and we have $C_{k}^{*}(u,v)=C^{*}(u,v;\theta_{k})$ for a sequence of parameters $(\theta_{k})_{k\in\mathbb{N}}$.
+However, we could obviously choose the $C_{k}^{*}$ from different parametric families.
+Appendix S7 Detailed estimation results
+Table S3 contains the estimation results for various D-vine models fitted to 10000 simulated data from a GARCH(1,1) process with symmetric innovations from a t distribution with 4 degrees of freedom.
+The most effective models in terms of AIC and BIC are the ast-AR(5) and ast-ARMA(1,1) models.
+Model D-vine order No. pars Loglik AIC BIC
+T1 1 2 442.8 -881.6 -867.2
+B1 1 5 436.6 -863.2 -827.1
+Joe-AR(1) 1 3 440.6 -875.1 -853.5
+Clayton180-AR(1) 1 3 424.6 -843.2 -821.6
+ast-AR(1) 1 3 442.8 -879.5 -857.9
+T2 2 4 576.3 -1144.6 -1115.7
+B2 2 10 568.6 -1117.3 -1045.2
+Joe-AR(2) 2 4 564.4 -1120.7 -1091.9
+Clayton180-AR(2) 2 4 552.1 -1096.2 -1067.4
+ast-AR(2) 2 4 573.1 -1138.2 -1109.4
+T5 5 10 695.6 -1371.3 -1299.2
+B5 5 25 689.7 -1329.4 -1149.2
+Joe-AR(5) 5 7 681.9 -1349.8 -1299.3
+Clayton180-AR(5) 5 7 667.9 -1321.9 -1271.4
+ast-AR(5) 5 7 693.9 -1373.7 -1323.3
+Joe-ARMA(1,1) 20 4 682.0 -1356.0 -1327.2
+Clayton180-ARMA(1,1) 20 4 669.5 -1331.0 -1302.2
+ast-ARMA(1,1) 20 4 694.0 -1380.0 -1351.1
+Table S3: Comparison of AIC and BIC values when various D-vines are fitted to simulated data from a GARCH(1,1) process with symmetric t4 innovations.
+Table S4 contains corresponding results for 10000 data from a GARCH(1,1) process with skewed t4 innovations while Table S5 is for a model which has both skewed t4 innovations and leverage.
+Recall that parameter values may be found in Table 1.
+In both these analyses the preferred models are ast-AR(5) and ast-ARMA(1,1).
+The improvement of ast-ARMA(1,1) on ast-AR(5) is not quite so large as in the case of the empirical data in Table 4, since the persistence of volatility in the simulated GARCH processes is lower.
+Model D-vine order No. pars Loglik AIC BIC
+T1 1 2 503.0 -1002.1 -987.7
+B1 1 5 499.5 -989.0 -953.0
+Joe-AR(1) 1 3 500.8 -995.6 -974.0
+Clayton180-AR(1) 1 3 481.1 -956.2 -934.6
+ast-AR(1) 1 3 505.5 -1005.1 -983.4
+T2 2 4 649.2 -1290.5 -1261.7
+B2 2 10 650.7 -1281.4 -1209.3
+Joe-AR(2) 2 4 648.5 -1289.0 -1260.1
+Clayton180-AR(2) 2 4 626.3 -1244.6 -1215.7
+ast-AR(2) 2 4 656.4 -1304.9 -1276.0
+T5 5 10 747.5 -1475.0 -1402.9
+B5 5 25 756.4 -1462.8 -1282.6
+Joe-AR(5) 5 7 744.4 -1474.7 -1424.2
+Clayton180-AR(5) 5 7 723.0 -1431.9 -1381.4
+ast-AR(5) 5 7 758.4 -1502.8 -1452.4
+Joe-ARMA(1,1) 20 4 744.5 -1481.0 -1452.2
+Clayton180-ARMA(1,1) 20 4 726.5 -1445.0 -1416.1
+ast-ARMA(1,1) 20 4 759.1 -1510.3 -1481.4
+Table S4: Comparison of AIC and BIC values when various D-vines are fitted to simulated data from a GARCH(1,1) process with skew t4 innovations.
+Model D-vine order No. pars Loglik AIC BIC
+T1 1 2 569.2 -1134.4 -1120.0
+B1 1 5 588.6 -1167.3 -1131.2
+Joe-AR(1) 1 3 590.2 -1174.3 -1152.7
+Clayton180-AR(1) 1 3 568.5 -1130.9 -1109.3
+ast-AR(1) 1 3 597.6 -1189.1 -1167.5
+T2 2 4 723.8 -1439.5 -1410.7
+B2 2 10 760.5 -1501.0 -1428.8
+Joe-AR(2) 2 4 754.4 -1500.8 -1472.0
+Clayton180-AR(2) 2 4 725.6 -1443.3 -1414.4
+ast-AR(2) 2 4 763.6 -1519.1 -1490.3
+T5 5 10 813.8 -1607.6 -1535.5
+B5 5 25 888.2 -1726.5 -1546.3
+Joe-AR(5) 5 7 868.1 -1722.1 -1671.6
+Clayton180-AR(5) 5 7 828.8 -1643.6 -1593.1
+ast-AR(5) 5 7 879.6 -1745.1 -1694.6
+Joe-ARMA(1,1 20 4 866.0 -1724.0 -1695.1
+Clayton180-ARMA(1,1) 20 4 831.0 -1654.1 -1625.2
+ast-ARMA(1,1) 20 4 880.0 -1752.0 -1723.2
+Table S5: Comparison of AIC and BIC values when various D-vines are fitted to simulated data from a GARCH(1,1) process with skew t4 innovations and leverage.
+Appendix S8 Analysis of VaR exceedances
+After model estimation, the Rosenblatt functions $R_{k}$ in (S6) can be used to construct one-step-ahead conditional quantile or value-at-risk (VaR) forecasts for all our fitted D-vine copula models.
+We follow Loaiza-Maya et al., (2018) by using this as the basis of a method for evaluating the quality of the fitted models.
+Specifically, for $t=2,\ldots,n=3668$ we compare the realized values $u_{t}$ with VaR forecasts $\operatorname{VaR}_{t}(\alpha)$ derived according to the formula
+$\operatorname{VaR}_{t}(\alpha)=R_{k(t)}^{-1}(\alpha;u_{t-1},\ldots,u_{t-k(t)})$
+for $\alpha\in\{0.01,0.05,0.1,0.9,0.95,0.99\}$.
+The number of conditioning variable in the forecast is taken to be $k(t)=\min(t-1,p,12)$ where $p$ is the Markov order of the fitted model and 12 is a cap on the Markov order to ensure reasonably fast run times.
+Note that, we do not compare the realized raw data $x_{t}$ with estimates of the conditional quantiles of the composite model for $(X_{t})$, but rather we compare the data transformed to the copula scale $u_{t}=\widehat{F}_{X}(x_{t})$ with estimates of the conditional quantiles of the D-vine process $(U_{t})$; this means that we evaluate the quality of the copula model independently of the quality of the adaptive kernel density estimator $\widehat{F}_{X}$ of Loaiza-Maya et al., (2018) which is used for the marginal distribution $F_{X}$.
+The hit rates $\hat{\alpha}=\tfrac{1}{n-1}\sum_{t=2}^{n}I_{\{u_{t}<\operatorname{VaR}_{t}(\alpha)\}}$ for our favoured models (Joe-AR(5), ast-AR(5), Clayton180-ARMA(1,1) and ast-ARMA(1,1)) are shown in Table S6 and are very similar to the rates reported by Loaiza-Maya et al., (2018).
+In all cases the conditional coverage test of Christoffersen, (1998) leads to an insignificant test result with confidence level 99%.
+Quantile level $\alpha$ (percent)
+Model 1 5 10 90 95 99
+Joe-AR(5) 0.98 4.85 9.54 90.35 95.07 99.13
+ast-AR(5) 0.95 5.04 9.60 90.08 94.98 99.18
+ast-ARMA(1,1) 1.06 4.83 9.87 90.21 94.98 99.05
+Clayton180-ARMA(1,1) 1.04 4.88 10.33 90.19 94.77 99.15
+Table S6: Empirical hit rates (expressed as percentages) for 1-day-ahead conditional $\alpha$-quantile (VaR) estimates in analysis of USD/AUD exchange-rate returns.
+For details of models see Section 4.2 and Table 4.
+Appendix S9 Copula of an AR(1) model with ARCH(1) errors
+In this section we consider a time series model with stochastic volatility as well as a non-zero stochastic conditional mean term.
+Specifically we consider an AR(1) model with ARCH(1) errors which is a first-order Markov process $(X_{t})_{t\in\mathbb{Z}}$ satisfying equations
+$X_{t}=\phi X_{t-1}+\sigma(X_{t-1})\;\epsilon_{t}\quad\text{with}\quad\sigma(x)=\sqrt{\alpha_{0}+\alpha_{1}x^{2}},$ (S11)
+where the innovations $(\epsilon_{t})_{t\in\mathbb{Z}}$ have a symmetric distribution.
+This model was proposed and analysed by Borkovec and Klüppelberg, (2001) who showed that it is covariance stationary provided $\phi^{2}+\alpha_{1}<1$ and has a symmetric stationary distribution $F_{X}$.
+Note that it differs from the usual version of the AR(1)-ARCH(1) model, $X_{t}=\phi X_{t-1}+\sigma(X_{t-1}-\phi X_{t-2})\;\epsilon_{t}$, but this is not first-order Markov and thus less straightforward to treat with our methods.
+The formulas (8) require only simple modification to account for the non-zero conditional mean term. In particular, the copula becomes
+$c_{1}(u,v)=\frac{1}{\sigma\left(F_{X}^{-1}(u)\right)}f_{\epsilon}\left(\frac{F_{X}^{-1}(v)-\phi F_{X}^{-1}(u)}{\sigma\left(F_{X}^{-1}(u)\right)}\right)\frac{1}{f_{X}(F_{X}^{-1}(v))}\,.$
+It may be easily verified that the copula density satisfies c(u,v)=c(1u,1v)c(u,v)=c(1-u,1-v) and so the copula is radially symmetric.
+However, the copula is neither h-symmetric nor v-symmetric, and thus it is not jointly symmetric.
+The copula density is shown in the bottom right panel of Figure S2 in the case where $\phi=0.3$, $\alpha_{0}=0.4$ and $\alpha_{1}=0.6$.
+It is evident that more of the probability mass is now on the main diagonal of the copula.
+Behaviour of this kind could perhaps be modelled by a mixture of a Gauss copula (or Frank copula) and one of the inverse-v-transformed copulas examined earlier.